### v1

# Characterizing higher Auslander(-Gorenstein) Algebras

Abstract.
It is well known that for Auslander algebras, the category of all (finitely generated) projective modules is an abelian category and this property of abelianness characterizes Auslander algebras by Tachikawa theorem in 1974.
Let $n$ be a positive integer. In this paper, by using torsion theoretic methods, we show that $n$-Auslander algebras can be characterized by the abelianness of the category of modules with projective dimension less than $n$ and a certain additional property, extending the classical Auslander-Tachikawa theorem. By Auslander-Iyama correspondence a categorical characterization of the class of Artin algebras having $n$-cluster tilting modules is obtained.
Since higher Auslander algebras are a special case of higher Auslander-Gorenstein algebras, the results are given in the general setting as extending previous results of Kong. Moreover, as an application of some results, we give categorical descriptions for the semisimplicity and selfinjectivity of an Artin algebra.
Higher Auslander-Gorenstein Algebras are also studied from the viewpoint of cotorsion pairs and, as application, we show that they satisfy in two nice equivalences.
Key words and phrases:
Auslander algebra; Auslander-Gorenstein algebra; Cluster tilting module; Gorenstein projective module; Torsion theory; Cotorsion theory
2020 Mathematics Subject Classification:
16G10; 16E65
Introduction
An important problem in the representation theory of algebras is to study algebras of finite representation type; i.e. Artin algebras that have only a finite number of isomorphism classes of finitely generated indecomposable modules.
In 1971, Auslander [Au71] proved a remarkable classical result, called Auslander correspondence, about algebras of finite representation type which introduced a completely new insight to the representation theory of Artin algebras. He showed that there is a bijective correspondence between the set of Morita equivalence classes of Artin algebras $\Lambda$ of finite representation type and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an Auslander algebra; i.e. i.e. Artin algebras $\Gamma$ whose global dimensions (denoted by ${\rm{gldim}}(\Gamma)$) are at most two and whose dominant dimensions (denoted by ${\rm{domdim}}(\Gamma)$) are at least two. See [Au71, Chapter III, Section 4, Page 52] or [ARS, Theorem VI.5.7].
For an Artin algebra $\Gamma$, let ${\rm{Mod}}(\Gamma)$ be the category of all left $\Gamma$-modules and ${\rm{{mod}}}(\Gamma)$ be the category of all finitely generated left $\Gamma$-modules. If $\Gamma$ is an Auslander algebra, then the full subcategory $\mathrm{proj}(\Gamma)$ of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated projective $\Gamma$-modules is equivalent to a module category and so is itself an abelian category [Au71, §III, Section 4, Page 52]. This gave Tachikawa the motivation to study rings $\Gamma$ such that the full subcategory ${\mathrm{Proj}}(\Gamma)$ of ${\rm{Mod}}(\Gamma)$ consisting of all projective $\Gamma$-modules is abelian. He obtained some characterizations of such rings and proved that these algebras are precisely Auslander algebras [Ta74, Theorem 1]. For Artin algebras, by combining Auslander’s result and Tachikawa’s result, we have the following characterization of Auslander algebras.
Theorem 1 (Auslander-Tachikawa Theorem).
For an Artin algebra $\Gamma$, the following statements are equivalent:
  • $(a)$
    $\Gamma$ is an Auslander algebra;
  • $(b)$
    the full additive subcategory $\mathrm{proj}(\Gamma)$ of ${\rm{{mod}}}(\Gamma)$ is an abelian category;
  • $(c)$
    the full additive subcategory ${\mathrm{Proj}}(\Gamma)$ of ${\rm{Mod}}(\Gamma)$ is an abelian category.
Hence, for Artin algebras, by Auslander correspondence and Auslander-Tachikawa theorem, there are bijections between the following four sets:
  • (i)
    Morita equivalence classes of Artin algebras $\Lambda$ such that $\Lambda$ is of finite representation type;
  • (ii)
    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an Auslander algebra;
  • (iii)
    Morita equivalence classes of Artin algebras $\Delta$ such that $\mathrm{proj}(\Delta)$ is an abelian category;
  • (iv)
    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{Proj}}(\Sigma)$ is an abelian category.
Moreover, an Artin algebra $\Lambda$ is of finite representation type if and only if ${\rm{{mod}}}(\Lambda)\simeq\mathrm{proj}(\Gamma)$, where $\Gamma$ is an Auslander algebra; see also [Iy04, Theorem 3.3].
In 2007, Iyama in a series of important papers [Iy07a, Iy07b] gave a generalization of Auslander correspondence, called higher Auslander correspondence or Auslander-Iyama correspondence. He showed that for every positive integer $n$, there is a bijective correspondence between the set of equivalence classes of $n$-cluster tilting subcategories with additive generators (Definition 5.5) of Artin algebras and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that ${\rm{gldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$, called $n$-Auslander algebras [Iy07b, Theorem 0.2].
A natural question is whether there exists a characterization for higher Auslander algebras analogous to Auslander-Tachikawa Theorem. The first goal of this paper is to answer this question and we have the following characterization for higher Auslander algebras; see Theorem 5.1.
Theorem 2 (Higher Auslander-Tachikawa Theorem).
Let $n$ be a positive integer and $\Gamma$ be an Artin algebra. Then the following statements are equivalent:
  • $(a)$
    $\Gamma$ is an $n$-Auslander algebra;
  • $(b)$
    $\mathrm{proj}^{\leq n-1}(\Gamma)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$, and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$;
  • $(c)$
    ${\mathrm{Proj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$, and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n}}\Gamma}$.
Here, for an Artin algebra $\Gamma$, we denote by $\Gamma^{{\rm{op}}}$ the opposite algebra of $\Gamma$, by ${\mathrm{Inj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all injective $\Gamma$-modules, by $\mathrm{inj}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated injective $\Gamma$-modules; for every non-negative integer $m$, by ${\mathrm{Proj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of projective dimension at most $m$, and by $\mathrm{proj}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of projective dimension at most $m$. For $m\geq 1$, we write
[…]
and
[…]
we also write
${}^{\perp_{0}}\Gamma:=\{M\in{\rm{{mod}}}(\Gamma)\ |\ {\rm{Hom}}_{\Gamma}(M,\Gamma)=0\},\ {}^{\underline{\perp}_{0}}\Gamma:=\{M\in{\rm{Mod}}(\Gamma)\ |\ {\rm{Hom}}_{\Gamma}(M,\Gamma)=0\}.$
Note that only the abelianness property of the category of modules with projective dimension less than $n$ can not characterize $n$-Auslander algebras. For $n\geq 2$, the other conditions are necessary, although they hold automatically for $n=1$; see Remarks 5.2 and 5.3.
In 1993, Auslander and Solberg [AS93] established a Gorenstein analogue of Auslander correspondence by using the technique of relative homology developed in [AS93a, AS93b, AS93c]. They introduced $\tau$-selfinjective algebras, where $\tau$ is the Auslander-Reiten translation, as a replacement of representation-finite algebras. The corresponding analogue of Auslander algebras are characterized as Artin algebras whose injective dimensions are at most $2$ and whose dominant dimensions are at least $2$, that is, Artin algebras $\Gamma$ such that ${\rm{id}}({}_{\Gamma}\Gamma)\leq 2\leq\rm{domdim}(\Gamma)$. Kong named this class of algebras as quasi-Auslander algebras [Ko14]. In his work [Ko14] Kong also proved an extension of Auslander-Tachikawa’s theorem by characterizing quasi-Auslander algebras as Artin algebras over which the category of finitely generated Gorenstein projective modules is an abelian category.
Hence, for Artin algebras, by Auslander-Solberg correspondence ([AS93, Proposition 5.11]) and Kong’s theorem [Ko14, Theorem 2.1], there are bijections between the following four sets:
  • (i)
    Equivalence classes of suitable modules of Artin algebras;
  • (ii)
    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is a quasi-Auslander algebra;
  • (iii)
    Morita equivalence classes of Artin algebras $\Delta$ such that ${\mathrm{Gproj}}(\Delta)$ is an abelian category;
  • (iv)
    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{GProj}}(\Sigma)$ is an abelian category.
To obtain a higher analogue of Auslander-Solberg correspondence, Iyama and Solberg [IS18] introduced pre-cluster tilting subcategories in module categories of Artin algebras which are called $\tau_{n}$-selfinjective algebras as a higher version of $\tau$-selfjinjective algebras. They also gave the notion of $n$-minimal Auslander-Gorenstein algebras as a higher version of quasi-Auslander algebras, say Artin algebras $\Gamma$ with ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$. They showed that there exists a bijective correspondence between Morita equivalence classes of $n$-minimal Auslander-Gorenstein algebras and equivalence classes of finite $n$-precluster tilting subcategories of Artin algebras.
In view of the above Iyama-Solberg correspondence, we could establish a Gorenstein version of Theorem 2 which are also higher versions of Kong’s result [Ko14, Theorem 2.1].
Theorem 3 (Theorem 2.6).
Let $n$ be a positive integer and $\Gamma$ be an Artin algebra. Then the following statements are equivalent.
  • $(a)$
    $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra;
  • $(b)$
    ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$;
  • $(c)$
    ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$ and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$.
Here for every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$, by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
Hence, for Artin algebras, by Iyama-Solberg correspondence [IS18, Theorem 4.5] and our results (Theorem 2.6), there are bijections between the following sets:
  • (I)
    Equivalence classes of finite $n$-precluster tilting subcategories of Artin algebras;
  • (II)
    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra;
  • (III)
    Morita equivalence classes of Artin algebras $\Delta$ such that ${\mathrm{Gproj}}^{\leq n-1}(\Delta)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Delta)\cap\mathrm{inj}(\Delta)\subseteq\mathrm{proj}(\Delta)$, and ${{}^{\perp_{0}}\Delta}\subseteq{{}^{\perp_{n}}\Delta}$;
  • (IV)
    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{GProj}}^{\leq n-1}(\Sigma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Sigma)\cap{\mathrm{Inj}}(\Sigma)\subseteq{\mathrm{Proj}}(\Sigma)$, and ${{}^{\underline{\perp}_{0}}\Sigma}\subseteq{{}^{\underline{\perp}_{n}}\Sigma}$.
While the results of Section 2 mainly make use of torsion theoretic methods, in Section 4, we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS]. Especially, by Corollary 4.4, if $n$ is a positive integer and $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, then we have the following equivalences:
${{}^{\perp_{0}}\Gamma}\simeq\frac{\mathrm{inj}^{\leq 1}(\Gamma)}{\mathrm{inj}^{\leq 1}(\Gamma)\bigcap\mathrm{proj}^{\leq n}(\Gamma)}\simeq\frac{\mathrm{inj}^{\leq 1}(\Gamma)}{\mathrm{inj}(\Gamma)}.$
The layout of this paper is as follows.
The first section contains some preliminaries including elementary terminology, basic notions and facts about dominant dimension, torsion pairs, projectivization and injectivization, injective resolutions, and Gorenstein projective modules.
Since Theorem 2 is a special case of Theorem 3, we study higher Auslander-Gorenstein algebras and prove our main result Theorem 3 in the second section.
In the third section, as an application of our work we study $\tau_{n}$-selfinjective algebras and give some characterizations of them in the categorical sense.
The fourth section of this paper, Section 4, is devoted to study the relationship between higher Auslander-Gorenstein algebras and cotorsion pairs and in it we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples. See Corollary 4.4.
We end the paper in Section 5, where we restrict our attention to higher Auslander algebras and rewrite some results of previous sections for them. The end example gives a better perspective about some results of the paper.
1. Preliminaries
In this section, for the convenience of the reader, we collect some definitions and results that will be used throughout the paper.
1.1. Conventions
Throughout this article all Artin algebras are assumed to be finitely generated modules over a fixed commutative Artinian ring $R$ with Jacobson radical $J_{R}$.
Let $\Gamma$ be an Artin algebra defined over $R$. We denote by $\mathbf{D}$ the canonical duality ${\rm{Hom}}_{R}(-,\mathbf{E}$ $(R/J_{R})):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$, where $\mathbf{E}(R/J_{R})$ is the injective envelop of $R/J_{R}$ as $R$-module.
For a left $\Gamma$-module $X$, $\mathbf{E}(X)$ denotes the injective envelope of $X$; ${\rm{pd}}({}_{\Gamma}X)$ (resp. ${\rm{id}}({}_{\Gamma}X)$) is the projective (resp. injective) dimension of $X$; while ${\rm{Add}}(X)$ is the full subcategory of ${\rm{Mod}}(\Gamma)$ whose objects are direct summands of direct sums of copies of $X$ and ${\rm{add}}(X)$ is the full subcategory of ${\rm{{mod}}}(\Gamma)$ whose objects are direct summands of finite direct sums of copies of $X$.
Given a subcategory $\mathcal{C}$ of ${\rm{Mod}}(\Gamma)$, for each $i\geq 1$, denote $\mathrm{sub}^{i}(\mathcal{C})$ to be the subcategory consisting of all modules $X$ which admit copresentations
$0\to X\to C^{0}\to\cdots\to C^{i-1}$
with $C^{0},\dots,C^{i-1}\in\mathcal{C}$, while $\mathrm{fac}_{i}(\mathcal{C})$ is formed by all modules $Y$ which admit presentations
$C_{i-1}\to\cdots\to C_{0}\to Y\to 0$
with $C_{0},\dots,C_{i-1}\in\mathcal{C}.$ Also, by convention, $\mathrm{Sub}^{0}(\mathcal{C})=\mathrm{Fac}_{0}(\mathcal{C}):={\rm{Mod}}(\Gamma)$.
For a module $M\in{\rm{{mod}}}(\Gamma)$ and $i\geq 0$, write $\mathrm{sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{add}(M))$, $\mathrm{fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{add}(M))$, $\mathrm{Sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{Add}(M))$, $\mathrm{Fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{Add}(M))$. In particular, $\mathrm{sub}(M):=\mathrm{sub}^{1}(M)$, $\mathrm{fac}(M)$ $:=\mathrm{fac}_{1}(M)$, $\mathrm{Sub}(M):=\mathrm{Sub}^{1}(M)$, $\mathrm{Fac}(M):=\mathrm{Fac}_{1}(M)$.
1.2. Dominant Dimension
Let $n$ be a positive integer and $\Gamma$ be a ring. Recall that for a left $\Gamma$-module $M$ if in its minimal injective resolution
$0\longrightarrow{{}_{\Gamma}M}\longrightarrow I^{0}\longrightarrow I^{1}\longrightarrow\cdots,$
the first $n$ terms $I^{0},\dots,I^{n-1}$ are projective, then the dominant dimension of $M$ is at least $n$, denoted by $\mathrm{domdim}({{}_{\Gamma}M})\geq n$. Hence, the dominant dimension of $M$ is equal to the smallest $n$ such that $I^{n}$ is not projective, or it is infinite if no such $n$ exists.
The left dominant dimension of the ring $\Gamma$ is defined as the dominant dimension of the left regular module ${{}_{\Gamma}\Gamma}$. Note that the dominant dimension of a right module and the right dominant dimension of a ring are defined similarly. It is well-known that ${\rm{domdim}}({{}_{\Gamma}\Gamma})={\rm{domdim}}(\Gamma_{\Gamma})$ ; see [Mu68, Theorem 4]. So for the rest of the paper, we will denote both left and right dominant dimension of $\Gamma$, by ${\rm{domdim}}(\Gamma)$ and call it the dominant dimension of $\Gamma$.
The following simple observation will be useful in Section 2.
Lemma 1.1.
For a ring $\Gamma$, if ${\rm{domdim}}(\Gamma)\geq n+1,$ then ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$.
Proof.
Let
(1.1) $0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\longrightarrow\cdots$
be the minimal injective resolution of ${}_{\Gamma}\Gamma$, then $I^{j}$ is a projective $\Gamma$-module for every $0\leq j\leq n$.
To prove that ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$, let $M\in{{}^{\perp_{0}}\Gamma}$. In the minimal injective resolution (1.1), as $I^{j}$ is projective for every $0\leq j\leq n$, we have ${\rm{Hom}}_{\Gamma}(M,I^{j})=0$. Applying the functor ${\rm{Hom}}_{\Gamma}(M,-)$ to (1.1) shows that ${\rm{Ext}}^{i}_{\Gamma}(M,\Gamma)=0$, for every $1\leq i\leq n$. So $M\in{{}^{\perp_{n}}\Gamma}$.
The inclusion ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$ can be proves exactly in the same way. ∎
1.3. Torsion Pairs
Let $\mathcal{A}$ be an abelian category. If $X,Y\in\mathcal{A}$, then we denote the set of morphisms from $X$ to $Y$ in $\mathcal{A}$ by $\mathcal{A}(X,Y)$ or ${\rm{Hom}}_{\mathcal{A}}(X,Y)$. All subcategories considered are assumed to be full and closed under isomorphisms and finite direct sums.
A torsion pair (or torsion theory) in the abelian category $\mathcal{A}$ is a pair $(\mathcal{X},\mathcal{Y})$ of subcategories such that $\mathcal{A}(X,Y)=0$ for all $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$, and these two classes are maximal for this property [Di66, Sections 1 and 2], that is,
$\mathcal{X}={{}^{\perp_{0}}\mathcal{Y}}:=\{X\in\mathcal{A}\ |\ \mathcal{A}(X,Y)=0,\forall\ Y\in\mathcal{Y}\},$
$\mathcal{Y}={\mathcal{X}^{\perp_{0}}}:=\{Y\in\mathcal{A}\ |\ \mathcal{A}(X,Y)=0,\forall\ X\in\mathcal{X}\}.$
If $(\mathcal{X},\mathcal{Y})$ is a torsion pair in $\mathcal{A}$, $\mathcal{X}$ is called a torsion class and $\mathcal{Y}$ is called a torsion-free class.
A subcategory of $\mathcal{A}$ is the torsion class (resp. the torsion-free class) of some torsion pair if and only if it is closed under quotients, direct sums and extensions (resp. subobjects, direct products and extensions, respectively) [Di66, Theorem 2.3].
If $(\mathcal{X},\mathcal{Y})$ is a torsion pair in $\mathcal{A}$, to each object $A$ in $\mathcal{A}$ is associated a unique short exact sequence (called the canonical sequence for $A$)
$0\longrightarrow xA\longrightarrow A\longrightarrow yA\longrightarrow 0$
such that $xA\in\mathcal{X}$ and $yA\in\mathcal{Y}$. In fact, $xA$ is the maximal subobject of $A$ contained in $\mathcal{X}$ [Di66, Proposition 2.4]. Also, the canonical sequence is functorial, i.e. it defines the functors $x:\mathcal{A}\longrightarrow\mathcal{X}$ given by $A\mapsto xA$ and $y:\mathcal{A}\longrightarrow\mathcal{Y}$ given by $A\mapsto yA$, see [BR07, Section 1.1]. Note that the composition of the functor $x:\mathcal{A}\longrightarrow\mathcal{X}$ by the inclusion functor $\mathcal{X}\longrightarrow\mathcal{A}$ is an idempotent radical, see [St75, Proposition VI.2.3].
A torsion pair $(\mathcal{X},\mathcal{Y})$ in $\mathcal{A}$ is called hereditary if $\mathcal{X}$ is also hereditary, i.e. $\mathcal{X}$ is closed under subobjects (which is equivalent to $\mathcal{Y}$ being closed under taking injective envelopes [Di66, Theorem 2.9]).
Let now $\Gamma$ be an Artin algebra. We would like to mention that in this case, a subcategory of ${\rm{{mod}}}(\Gamma)$ is the torsion class (resp. the torsion-free class) of some torsion pair if and only if it is closed under quotients and extensions (resp. submodules and extensions, respectively). For more details about torsion theories in ${\rm{{mod}}}(\Gamma)$; see [AIR, Subsection 1.1] or [Iy04, Subsection 1.2].
1.4. Functorially Finite Subcategories
Let $\Gamma$ be an Artin algebra. A subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Gamma)$ is said to be covariantly finite in ${\rm{{mod}}}(\Gamma)$, if all representable covariant functors on ${\rm{{mod}}}(\Gamma)$ are finitely generated when considered as functors on $\mathcal{X}$ by restriction. In other words, for each $M\in{\rm{{mod}}}(\Gamma)$, there exist $X\in\mathcal{X}$ and a morphism $\varphi:M\longrightarrow X$ so that the sequence
${\rm{Hom}}_{\Gamma}(X,-)\mid_{\mathcal{X}}\longrightarrow{\rm{Hom}}_{\Gamma}(M,-)\mid_{\mathcal{X}}\longrightarrow 0$
is exact. Here, the notation $\mid_{\mathcal{X}}$ means restriction to $\mathcal{X}$.
Dually, $\mathcal{X}$ is said to be contravariantly finite in ${\rm{{mod}}}(\Gamma)$ if the representable contravariant functors on ${\rm{{mod}}}(\Gamma)$ stay finitely generated when restricted to $\mathcal{X}$.
A subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Gamma)$ is said to be functorially finite if it is both contravariantly and covariantly finite in ${\rm{{mod}}}(\Gamma)$.
Note that if the pair $(\mathcal{X},\mathcal{Y})$ is a torsion pair in ${\rm{{mod}}}(\Gamma)$, then an easy investigation shows that $\mathcal{X}$ is a contravariantly finite and $\mathcal{Y}$ is covariantly finite in ${\rm{{mod}}}(\Gamma)$ [Sm84, Page 519].
1.5. (Co)tiliting Modules
We recall some definitions and notations about (co)tilting modules and their relations to torsion theory; see [Ho82, As84, Sm84, AIR].
Definition 1.2 (Tilting and Cotilting Modules).
Let $\Gamma$ be an Artin algebra. A $\Gamma$-module $T$ is said to be a tilting module, if the following three properties are satisfied.
  • (i)
    The projective dimension of $T$ is at most one;
  • (ii)
    ${\rm{Ext}}_{\Gamma}^{1}(T,T)=0$;
  • (iii)
    There is an exact sequence $0\longrightarrow\Gamma\longrightarrow T^{\prime}\longrightarrow T^{\prime\prime}\longrightarrow 0$ with $T^{\prime}$ and $T^{\prime\prime}$ direct summands of direct sums of copies of $T$.
Dually, the $\Gamma$-module $T$ will be called a cotilting module if $\mathbf{D}(T)$ is a tilting $\Gamma^{\rm op}$-module.
Definition 1.3 (Ext-projectives and Ext-injectives).
Let $\Gamma$ be an Artin algebra and $\mathcal{X}$ be a subcategory of ${\rm{{mod}}}(\Gamma)$ closed under extensions. A module $X\in\mathcal{X}$ is said to be ${\rm{Ext}}$-projective (resp. ${\rm{Ext}}$-injective) if ${\rm{Ext}}^{1}_{\Gamma}(X,X^{\prime})=0$ (resp. ${\rm{Ext}}^{1}_{\Gamma}(X^{\prime},X)=0$) for all $X^{\prime}\in\mathcal{X}$.
We denote by $\mathcal{P}(\mathcal{X})$ the direct sum of one copy of each of the indecomposable ${\rm{Ext}}$-projective objects in $\mathcal{X}$ up to isomorphism.
Similarly, we denote by $\mathcal{I}(\mathcal{X})$ the direct sum of one copy of each of the indecomposable ${\rm{Ext}}$-injective objects in $\mathcal{X}$ up to isomorphism. Also, the annihilator ideal of $\mathcal{X}$ is denoted by $\mathrm{Ann}_{\Gamma}(\mathcal{X})$, i.e. $\mathrm{Ann}_{\Gamma}(\mathcal{X}):=\{\gamma\in\Gamma|\gamma X=0,\ \forall\ X\in\mathcal{X}\}$.
For the convenience of the reader, we recall the following relevant result on torsion pairs that we will use of it in next sections.
Theorem 1.4 ([Sm84, Theorem] and [AIR, Proposition 1.1]).
Let $\Gamma$ be an Artin algebra and the pair $(\mathcal{X},\mathcal{Y})$ be a torsion pair in ${\rm{{mod}}}(\Gamma)$. Then the following conditions are equivalent.
  • $(a)$
    $\mathcal{X}$ is functorially finite;
  • $(b)$
    $\mathcal{Y}$ is functorially finite;
  • $(c)$
    $\mathcal{X}=\mathrm{fac}(X)$ for some $X\in{\rm{{mod}}}(\Gamma)$;
  • $(d)$
    $\mathcal{Y}=\mathrm{sub}(Y)$ for some $Y\in{\rm{{mod}}}(\Gamma)$;
  • $(e)$
    $\mathcal{P}(\mathcal{X})$ is a tilting $(\Gamma/\mathrm{Ann}_{\Gamma}(\mathcal{X}))$-module;
  • $(f)$
    $\mathcal{I}(\mathcal{Y})$ is a cotilting $(\Gamma/\mathrm{Ann}_{\Gamma}(\mathcal{Y}))$-module;
  • $(g)$
    $\mathcal{X}=\mathrm{fac}(\mathcal{P}(\mathcal{X}))$;
  • $(h)$
    $\mathcal{Y}=\mathrm{sub}(\mathcal{I}(\mathcal{Y}))$.
1.6. Auslander-Reiten Theory
In this subsection we recall several results in Auslander-Reiten theory.
Let $\Gamma$ be an Artin algebra. The functor
()*$(-)^{*}:={\rm{Hom}}_{\Gamma}(-,\Gamma):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{{\rm{op}}})$
defined by
$M^{*}:={\rm{Hom}}_{\Gamma}(M,\Gamma)$
induces a duality ()*$(-)^{*}:{\rm{add}}(\Gamma)\longrightarrow{\rm{add}}(\Gamma^{{\rm{op}}})$, see e.g. [ARS, Proposition II.4.3].
For a $\Gamma$-module $X$, take a projective presentation $P_{1}\stackrel{{\scriptstyle g}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle f}}{{\longrightarrow}}X\longrightarrow 0.$ Define $\Gamma^{{\rm{op}}}$-module $\mathrm{Tr}X$ by an exact sequence $P_{0}^{*}\stackrel{{\scriptstyle g^{*}}}{{\longrightarrow}}P_{1}^{*}\longrightarrow\mathrm{Tr}X\longrightarrow 0.$ In fact $\mathrm{Tr}X$ is defined as the cokernel of $g^{*}$.
On the other hand, the canonical duality $\mathbf{D}:{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$ induces a duality $\mathbf{D}:\underline{{\rm{{mod}}}}(\Gamma)\longrightarrow\overline{{\rm{{mod}}}}(\Gamma^{\rm{\rm{op}}})$ and we have mutually quasi-inverse equivalences
$\tau:=\mathbf{D}\mathrm{Tr}:\underline{{\rm{{mod}}}}(\Gamma)\stackrel{{\scriptstyle\mathrm{Tr}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma^{{\rm{op}}})\stackrel{{\scriptstyle\mathbf{D}}}{{\longrightarrow}}\overline{{\rm{{mod}}}}(\Gamma)$
$\tau^{-}:=\mathrm{Tr}\mathbf{D}:\overline{{\rm{{mod}}}}(\Gamma)\stackrel{{\scriptstyle\mathbf{D}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma^{{\rm{op}}})\stackrel{{\scriptstyle\mathrm{Tr}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma)$
called the Auslander-Reiten translations. Especially, $\tau$ gives a bijection from isoclasses of indecomposable non-projective $\Gamma$-modules to isoclasses of indecomposable non-injective $\Gamma$-modules, see e.g. [ARS, Propositions IV.1.9 and IV.1.10].
1.7. Cotorsion Pairs
For a given class $\mathcal{C}$ of an abelian category $\mathcal{A}$, we let
${}^{\perp_{1}}\mathcal{C}:=\{A\in\mathcal{A}\mid{\rm{Ext}}^{1}_{\mathcal{A}}(A,C)=0,\ \forall\ C\in\mathcal{C}\},$
$\mathcal{C}^{\perp_{1}}:=\{A\in\mathcal{A}\mid{\rm{Ext}}^{1}_{\mathcal{A}}(C,A)=0,\ \forall\ C\in\mathcal{C}\}.$
Let now $\mathcal{X}$ and $\mathcal{Y}$ be full subcategories of the abelian category $\mathcal{A}$. Recall that a pair $(\mathcal{X},\mathcal{Y})$ is called a cotorsion pair (or cotorsion theory) if $\mathcal{X}={}^{\perp_{1}}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp_{1}}$, see for instance, [EJ11, Definition 7.1.2]. In this case, the class $\mathcal{X}$ is called a cotorsion class and the class $\mathcal{Y}$ is called a cotorsion-free class.
The cotorsion pair $(\mathcal{X},\mathcal{Y})$ is said to be hereditary, if ${\rm{Ext}}^{i}_{\mathcal{A}}(X,Y)=0$, for all $i\geq 1,X\in\mathcal{X}$, and $Y\in\mathcal{Y}$.
The cotorsion pair $(\mathcal{X},\mathcal{Y})$ is also called complete, if it satisfies the following two conditions:
  • $(i)$
    For any object $A$ of $\mathcal{A}$, there exists a short exact sequence $0\longrightarrow yA\longrightarrow xA\longrightarrow A\longrightarrow 0$, where $yA\in\mathcal{Y}$ and $xA\in\mathcal{X}$, i.e. $\mathcal{X}$ is a special precovering class;
  • $(ii)$
    For any object $A$ of $\mathcal{A}$, there exists a short exact sequence $0\longrightarrow A\longrightarrow\tilde{y}A\longrightarrow\tilde{x}A\longrightarrow 0$, where $\tilde{y}A\in\mathcal{Y}$ and $\tilde{x}A\in\mathcal{X}$, i.e. $\mathcal{Y}$ is a special preenveloping class.
Salce’s lemma (which goes back to [Sa79]) asserts that $(i)$ and $(ii)$ are equivalent in the case where $\mathcal{A}$ is a module category, see, e.g. [EJ11, Proposition 7.1.7]. The proof of this lemma shows that if the abelian category $\mathcal{A}$ has enough injectives, then $(i)\Rightarrow(ii)$ and if $\mathcal{A}$ has enough projectives, then $(ii)\Rightarrow(i)$.
In contrast to the case with torsion pairs, the constructions $x,\tilde{x}$ and $y,\tilde{y}$ are not functorial. But they define functors from $\mathcal{A}$ to $\frac{\mathcal{X}}{\mathcal{X}\cap\mathcal{Y}}$ and $\frac{\mathcal{Y}}{\mathcal{X}\cap\mathcal{Y}}$, respectively [BBOS, Lemma 2.11].
Note that there are slightly different definitions of cotorsion pairs which have appeared in the literature, see [BR07, KS03]. By [BR07, Remark V.3.2], there is no difference between the notion of a complete cotorsion pair that we recalled it in the above definition and the notion of a cotorsion pair in the sense of Beligiannis-Reiten [BR07, Definition V.3.1]. Also, cotorsion pairs in the sense of Krause-Solberg [KS03, Definition 2.1] are hereditary complete cotorsion pairs.
1.8. Projectivization and Injectivization
Let $\Lambda$ be an Artin algebra and $M\in{\rm{{mod}}}(\Lambda)$. Denote $\Gamma=\mathrm{End}_{\Lambda}(M)^{{\rm{op}}}$. Then the natural functor
$\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$
restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$; see, for instance, [ARS, Proposition II.2.1 (c)].
Lemma 1.5.
Let $P\in\mathrm{proj}(\Lambda)$ be a finitely generated projective module and $\Gamma=\mathrm{End}_{\Lambda}(P)^{{\rm{op}}}$.
  • (a)
    The functor
    $\mathrm{Hom}_{\Lambda}(P,-):\mathrm{fac}_{2}(P)\to{\rm{{mod}}}(\Gamma)$
    is an equivalence.
  • (b)
    The equivalence in (a) extends to another equivalence
    $\mathrm{Hom}_{\Lambda}(P,-):\mathrm{Fac}_{2}(P)\to{\rm{Mod}}(\Gamma).$
Proof.
The first statement is well known; see, for instance, [ARS, Proposition II.2.5].
For the second statement, since $P$ is finitely generated, $\mathrm{Hom}_{\Lambda}(P,-)$ commutes with direct sums, thus it establishes an equivalence between $\mathrm{Add}(P)$ and ${\mathrm{Proj}}(\Gamma)$. Then the proof proceeds as that of [ARS, Proposition II.2.5]. ∎
Dually, let $\Gamma$ be an Artin algebra and $N\in{\rm{{mod}}}(\Gamma)$. Denote $\Lambda=\mathrm{End}_{\Gamma}(N)^{{\rm{op}}}$. Then the natural functor
$\mathbf{D}\mathrm{Hom}_{\Gamma}(-,N):{\rm{{mod}}}(\Gamma)\to{\rm{{mod}}}(\Lambda)$
restricts to an equivalence $\mathrm{add}(N)\simeq\mathrm{inj}(\Lambda)$.
Lemma 1.6.
Let $\mathcal{Q}\in\mathrm{inj}(\Gamma)$ be a finitely generated injective module and $\Lambda=\mathrm{End}_{\Gamma}(\mathcal{Q})^{{\rm{op}}}$.
  • (a)
    The functor
    $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q}):\mathrm{sub}^{2}(\mathcal{Q})\simeq{\rm{{mod}}}(\Lambda)$
    is an equivalence.
  • (b)
    Let $I=\mathrm{Hom}_{\Gamma^{{\rm{op}}}}(\mathbf{D}\mathcal{Q},\Gamma)$, Then the two functors $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q})$ and $\mathrm{Hom}_{\Gamma}(I,-)$ from ${\rm{{mod}}}(\Gamma)$ to ${\rm{{mod}}}(\Lambda)$ are naturally isomorphic, which are furthermore isomorphic to $\mathbf{D}\mathcal{Q}\otimes_{\Gamma}-$.
  • (c)
    The equivalence in (a) extends to another equivalence
    $\mathrm{Hom}_{\Gamma}(I,-)\simeq\mathbf{D}\mathcal{Q}\otimes_{\Gamma}-:\mathrm{Sub}^{2}(\mathcal{Q})\to{\rm{Mod}}(\Lambda).$
  • (d)
    The two categories $\mathrm{sub}^{2}(\mathcal{Q})$ and $\mathrm{Sub}^{2}(\mathcal{Q})$ are abelian categories.
Proof.
(a) This is the first lemma of [Au71, Section 4].
(b) It is easy to see that $I$ is a finitely generated projective module, as $I\cong\nu^{-1}\mathcal{Q}$, where $\nu^{-1}$ is the inverse Nakayama functor.
The two functors $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q})$ and $\mathrm{Hom}_{\Gamma}(I,-)$ are exact, in particular, right exact, so by Watts’ theorem (see, for example, [Ro09, Corollary 5.47]), they are isomorphic to a tensor functor. It suffices to check that $\mathbf{D}\mathrm{Hom}_{\Gamma}(\Gamma,\mathcal{Q})\cong\mathbf{D}\mathcal{Q}\cong\mathrm{Hom}_{\Gamma}(I,\Gamma)$. In fact, since $\mathbf{D}\mathcal{Q}$ is finitely generated projective, its double $\Gamma$-dual $\mathrm{Hom}_{\Gamma}(\mathrm{Hom}_{\Gamma^{{\rm{op}}}}(\mathbf{D}\mathcal{Q},\Gamma),\Gamma)=\mathrm{Hom}_{\Gamma}(I,\Gamma)$ is isomorphic to itself.
(c). Since $I$ is finitely generated projective,
$\mathrm{Hom}_{\Gamma}(I,-)\simeq\mathbf{D}\mathcal{Q}\otimes_{\Gamma}-:\mathrm{Sub}^{2}(\mathcal{Q})\to{\rm{Mod}}(\Lambda)$
commutes with direct sums, so the equivalence $\mathrm{add}(\mathcal{Q})\simeq\mathrm{inj}(\Lambda)$ extends to $\mathrm{Add}(\mathcal{Q})\simeq{\mathrm{Inj}}(\Lambda)$. As in the proof of the first lemma of [Au71, Section 4], the latter equivalence extends to another equivalence
$\mathrm{Hom}_{\Gamma}(I,-)\simeq\mathbf{D}\mathcal{Q}\otimes_{\Gamma}-:\mathrm{Sub}^{2}(\mathcal{Q})\to{\rm{Mod}}(\Lambda).$
(d) follows from (a) and (c). ∎
1.9. Injective Resolutions
For the convenience of the reader, this subsection is devoted to state some important results about injective resolutions which we’ll make use of them in several proofs.
Recall that an Iwanaga-Gorenstein algebra is a two-sided Noetherian ring such that its injective dimensions both as a left and as a right $\Gamma$-module are finite, which are equal by Zaks [Za69, Lemma A] ( see also [EJ11, Proposition 9.1.8]).
Theorem 1.7 (Iwanaga [Iw79, Theorem 2]).
Let $\Gamma$ be an Iwanaga-Gorenstein algebra with the minimal injective resolution
$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow I^{1}\longrightarrow\cdots.$
Then $\bigoplus_{i\geq 0}I^{i}$ is an injective cogenerator.
Theorem 1.8 (Iwanaga [Iw80, Theorem 3 (1)]).
Let $\Gamma$ be an Artinian Iwanaga-Gorenstein algebra with the minimal injective resolution
$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow 0$
Then the projective dimension of every direct summand $I$ of $I^{n}$ is $n$, i.e. ${\rm{pd}}({}_{\Gamma}I)=n$.
Hence, for Gorenstein algebras of Gorenstein dimension $n+1$, we have the following proposition.
Proposition 1.9.
Let $\Gamma$ be an Artinian Iwanaga-Gorenstein algebra of injective dimension at most $n+1$ with the minimal injective resolution
$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\to 0$
Then $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq{\rm{add}}(\oplus_{i=0}^{n}I^{i}).$
Proof.
If $M\in\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)$, then $M=\bigoplus_{l=1}^{r}M_{l}$, where $M_{l}$ is an injective indecomposable $\Gamma$-module and of projective dimension at most $n$. By Theorem 1.7, each $M_{l}$ is a direct summand of $I^{j}$ with $10\leq j\leq n+$ 1. But ${\rm{pd}}(M_{l})\leq n$, so by Theorem 1.8, $M_{l}$ is not a direct summand of $I^{n+1}$. Thus $M_{l}$ is a direct summand of $I^{j}$ with $0\leq j\leq n$. This shows that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq{\rm{add}}(\oplus_{i=0}^{n}I^{i}).$
Theorem 1.10 (Miyachi [Mi00, Coro llary 1.3]).
Let
$0\longrightarrow Y_{n}\longrightarrow\cdots\longrightarrow Y_{1}\longrightarrow Y_{0}\longrightarrow X\longrightarrow 0$
be an exact sequence of $\Gamma$-modules. If
$0\longrightarrow Y_{i}\longrightarrow I^{-i,0}\longrightarrow I^{-i,1}\longrightarrow\cdots$
is an injective resolution of $Y_{i}$ ($0\leq i\leq n$), then $X$ has the injective resolution
$0\longrightarrow X\longrightarrow T\longrightarrow\bigoplus_{r+s=1}I^{r,s}\longrightarrow\bigoplus_{r+s=2}I^{r,s}\longrightarrow\cdots\longrightarrow\bigoplus_{r+s=n}I^{r,s}\longrightarrow\cdots$
where $T$ is a direct summand of $\bigoplus_{r+s=0}I^{r.s}$.
1.10. Infinitely Generated Modules
For later use, we record some facts about infinitely generated modules over an Artin algebra.
Facts 1.11.
Let $\Gamma$ is an Artin algebra.
  • (a)
    ([CE56, p. 17, Exercise 8]) Since $\Gamma$ is left Noetherian, arbitrary direct sums of injective (left) modules are injective modules, see also [Ch60, Proposition 4.1];
  • (b)
    ([KMT, Theorem 3.2]) Since $\Gamma$ is left Artinian, every injective module is a direct sum of indecomposable injective modules; see also Matlis [Ma58, Theorem 2.5];
  • (c)
    (Chase [Ch60, Theorem 3.3]) Since $\Gamma$ is left prefect and right coherent, direct products of projective modules are projective modules;
  • (d)
    (Müller [Mu70, Theorem 3]) Since $\Gamma$ is a semi-perfect ring, each projective module is a direct sum of finitely generated indecomposable projective modules.
1.11. Gorenstein Projective Modules
Let $\Gamma$ be an Artin algebra. Recall that a $\Gamma$-module $M$ is said to be Gorenstein projective if there exists an exact sequence of projective modules
${\bf P}=\cdots\longrightarrow P_{1}\longrightarrow P_{0}\longrightarrow P^{0}\longrightarrow P^{1}\longrightarrow\cdots$
such that $M\simeq\mathrm{Im}(P_{0}\longrightarrow P^{0})$ and such that ${\rm{Hom}}_{\Gamma}(-,Q)$ leaves the sequence ${\bf P}$ exact whenever $Q$ is a projective module.
The class of all Gorenstein projective modules is denoted by ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{Gproj}}(\Gamma)$ denotes the subcategory of finitely generated Gorenstein projective modules; see [EJ11, Definition 10.2.1].
Gorenstein injective modules are defined dually. Denote by ${\mathrm{GInj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all Gorenstein injective $\Gamma$-modules and by ${\mathrm{Ginj}}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated Gorenstein injective $\Gamma$-modules; see [EJ95, Definition 2.1] or [EJ11, Definition 10.1.1].
The Gorenstein projective dimension of a $\Gamma$-module $M$ is at most $n\ (n\in\mathbb{N}_{0})$ if and only if $M$ has a Gorenstein projective resolution of length $n$, written as ${\rm{Gpd}}({}_{\Gamma}M)\leq n$. By [Ho04a, Theorem 2.10], ${\rm{Gpd}}({}_{\Gamma}M)\leq n$ if and only if there exists an exact sequence
$0\longrightarrow P_{n}\longrightarrow\cdots\longrightarrow P_{1}\longrightarrow G\longrightarrow M\longrightarrow 0,$
where $P_{n},\dots,P_{1}$ are projective modules and $G$ is a Gorenstein projective modules. Similarly, one can define the Gorenstein injective dimension ${\rm{Gid}}({}_{\Gamma}M)$ of $M$ [Ho04a, Definition 2.8]. Note that if ${\rm{pd}}({}_{\Gamma}M)$ is finite, then ${\rm{Gpd}}({}_{\Gamma}M)={\rm{pd}}({}_{\Gamma}M)$ [Ho04a, Proposition 2.27]. For every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$ and by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
Like the classical case, the supremum of Gorenstein projective dimensions of all left $\Gamma$-modules is called left Gorenstein global dimension of $\Gamma$ and is denoted by ${\rm{Ggldim}}(\Gamma)$. By [BM10, Theorem 1.1], the left Gorenstein global dimension of $\Gamma$ is also the supremum of Gorenstein injective dimensions of all left $\Gamma$-modules. By [EJ11, Theorem 12.3.1], there is also no difference between left and right Gorenstein global dimensions of $\Gamma$ (as every Artin algebra is Noetherian).
Also, the Gorenstein global dimension of $\Gamma$ is the supremum of Gorenstein projective dimensions of all finitely generated $\Gamma$-modules; see [Ch10, Theorem 3.2.8]. Moreover, ${\rm{Ggldim}}(\Gamma)\leq n$ if and only if $\Gamma$ is an $n$-Iwanaga-Gorenstein algebra (i.e. ${\rm{id}}({}_{\Gamma}\Gamma)\leq n$ and ${\rm{id}}(\Gamma_{\Gamma})\leq n$). In fact, in this case, we have ${\rm{Ggldim}}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)={\rm{id}}(\Gamma_{\Gamma})$; see [EJ11, Corollary 12.3.2] or [Ch10, Corollary 3.2.6].
For any left $\Gamma$-module $M$ with finite projective dimension, ${\rm{Gpd}}({}_{\Gamma}M)\leq n$ if and only if ${\rm{Ext}}^{i}_{\Gamma}(M,\Gamma$ )=0)=0) = 0 for all $i>n$; see [Ho04a, Theorem 2.20]. Also, the Gorenstein injective dimension of $M$, ${\rm{Gid}}({}_{\Gamma}M)$, equals the usual injective dimension of $M$, i.e. if $M$ is a $\Gamma$-module with ${\rm{pd}}({}_{\Gamma}M)<\infty$, then ${\rm{Gid}}({}_{\Gamma}M)={\rm{id}}({}_{\Gamma}M)$; see [Ho04b, Theorem 2.1]. This result has a dual counterpart and if $M$ is a $\Gamma$-module with ${\rm{id}}({}_{\Gamma}M)<\infty$, then ${\rm{Gpd}}({}_{\Gamma}M)={\rm{pd}}({}_{\Gamma}M)$; see [Ho04b, Theorem 2.2].
2. Higher Auslander-Gorenstein Algebras
In this section, we will study higher Auslander-Gorenstein algebras introduced by Iyama and Solberg in [IS18] as a generalization of higher Auslander algebras and give a new characterization of them in terms of the abelianness property of the category of modules of Gorenstein projective dimension less than $n$. Of course, the results can be considered as a higher generalization of Auslander-Tachikawa theorem.
We recall some basic notions and facts about higher Auslander-Gorenstein algebras.
Let $n$ be a positive integer. Recall that an Artin algebra $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$ [IS18, Definition 1.1]. $\Gamma$ is also called a higher Auslander-Gorenstein algebra, if $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra for a certain $n$.
By [AR94, Corollary 5.5], these algebras are Iwanaga-Gorenstein algebras. It is also easily checked that any $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is either selfinjective or satisfies ${\rm{id}}({}_{\Gamma}\Gamma)=n+1={\rm{domdim}}(\Gamma)$; see [IS18, Proposition 4.1].
Notice that, as mentioned in Subsection 1.11, for Gorenstein algebras, there is no difference between Gorenstein global dimention and injective dimension of $\Gamma$, i.e. $\mathrm{Ggldim}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)$, and so $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if and only if ${\rm{Ggldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$.
The following observation is the starting point and the key fact of this paper.
Proposition 2.1.
Let $n$ be a positive integer and $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Then ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is an abelian category. Moreover, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$
Proof.
By Lemma 1.1, $\rm{domdim}(\Gamma)\geq n+1$ implies ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$.
Let
$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\to 0$
be the minimal injective resolution of ${}_{\Gamma}\Gamma$. By Proposition 1.9, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq{\rm{add}}(\oplus_{i=0}^{n}I^{i})$, while the latter is included in $\mathrm{proj}(\Gamma)$, as $\rm{domdim}(\Gamma)\geq n+1$. This shows that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$.
Let $\mathcal{Q}$ be the maximal injective summand of $\Gamma$ and $\Lambda={\rm{End}}_{\Gamma}(\mathcal{Q})^{{\rm{op}}}$. Since $\mathcal{Q}$ is an injective $\Gamma$-module, by Lemma 1.6, $\mathbf{D}{\rm{Hom}}_{\Gamma}(-,\mathcal{Q})$ establishes an equivalence between $\mathrm{sub}^{2}(\mathcal{Q})$ and ${\rm{{mod}}}(\Lambda)$ and so $\mathrm{sub}^{2}(\mathcal{Q})$ is an abelian category. To complete the proof we show that
${\mathrm{Gproj}}^{\leq n-1}(\Gamma)=\mathrm{sub}^{2}(\mathcal{Q}).$
Let $X\in\mathrm{sub}^{2}(\mathcal{Q})$. Thus $X$ fits into a short exact sequence
$0\longrightarrow X\longrightarrow E^{0}\stackrel{{\scriptstyle f}}{{\longrightarrow}}E^{1}$
such that $E^{0},E^{1}\in{{\rm{add}}(\mathcal{Q})}$. Since $\mathcal{Q}$ is also projective, $E^{0}$ and $E^{1}$ are also projective. Let $X^{\prime}$ be the cokernel of the morphism $E^{0}\longrightarrow E^{1}$. Since ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}X^{\prime})\leq n+1$, so by [Ho04a, Theorem 2.20] $,{\rm{Gpd}}({}_{\Gamma}X)\leq n-1$, i.e. $X\in{\mathrm{Gproj}}^{\leq n-1}(\Gamma)$. This shows that ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)\supseteq\mathrm{sub}^{2}(\mathcal{Q})$.
Let $X$ be a finitely generated Gorenstein projective $\Gamma$-module. By definition, $X$ is a submodule of a projective $\Gamma$-module $P$. Since ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq{\rm{domdim}}(\Gamma)$, $P$ has the minimal injective resolution
$0\longrightarrow P\longrightarrow J^{0}\longrightarrow J^{1}\longrightarrow\cdots\longrightarrow J^{n+1}\longrightarrow 0$
such that $J^{i}\in\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$, $i=0,\cdots,n$. Hence the injective envelope $\mathbf{E}(X)$ of $X$ is a direct summand of $J^{0}$ and itself is also a projective module. In the short exact sequence
$0\longrightarrow X\longrightarrow\mathbf{E}(X)\longrightarrow K\longrightarrow 0,$
we have ${\rm{Gpd}}({}_{\Gamma}K)\leq 1$ and by [Ho04a, Theorem 2.10], there exists an exact sequence
$0\longrightarrow P^{\prime}\longrightarrow G\longrightarrow K\longrightarrow 0$
such that $G$ is a finitely generated Gorenstein projective module and $P^{\prime}$ is a finitely generated projective $\Gamma$-module. Since ${\rm{domdim}}(\Gamma)\geq n+1\geq 2$, both $\mathbf{E}(G)$ and $\mathbf{E}(\mathbf{E}(P^{\prime})/P^{\prime})$ are projective, so by Theorem 1.10, it follows that $\mathbf{E}(K)$ is a projective $\Gamma$-module as well. Hence, $X$ fits into
(2.1) $0\to X\to\mathbf{E}(X)\to\mathbf{E}(K)$
with both $\mathbf{E}(X)$ and $\mathbf{E}(K)$ being projective-injective, hence $X$ falls into $\mathrm{sub}^{2}(\mathcal{Q})$.
Now assume that $X$ is a $\Gamma$-module such that ${\rm{Gpd}}({}_{\Gamma}X)\leq n-1$. By [Ho04a, Theorem 2.10], there exists an exact sequence
(2.2) $0\longrightarrow P_{n-1}\longrightarrow\cdots\longrightarrow P_{1}\longrightarrow G\longrightarrow X\longrightarrow 0$
where $P_{i}$ is a finitely generated projective module for every $11\leq i\leq n-$ 1 and $G$ is a finitely generated Gorenstein projective module. Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, ${\rm{Ggldim}}(\Gamma)\leq n+1\leq{\rm{domdim}}(\Gamma)$ and so every projective $\Gamma$-module $P_{i}$ has the minimal injective resolution
(2.3) $0\longrightarrow P_{i}\longrightarrow J^{-i,0}\longrightarrow J^{-i,1}\longrightarrow\cdots\longrightarrow J^{-i,n+1}\longrightarrow 0$
such that $J^{-i,k}\in\mathrm{proj}(\Gamma)\ \cap\ \mathrm{inj}(\Gamma)$, for every $k=0,\dots,n$. Applying Theorem 1.10 to the exact sequence (2.2) together with (2.1) and (2.3) then immediately gives the required result.
We have shown that ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)=\mathrm{sub}^{2}(\mathcal{Q})$. ∎
In fact, we can show that the properties obtained in Proposition 2.1 are also sufficient to characterize higher Auslander-Gorenstein algebras. We need the following simple lemmas which are of independent interest. They should be known to the expert. However, we could not find them in the literature, so we include a proof.
Lemma 2.2.
Let $\Gamma$ be an Artin algebra. If a full additive category $\mathcal{C}$ of ${\rm{{mod}}}(\Gamma)$ has kernels and contains $\Gamma$, then the kernels in $\mathcal{C}$ coincident with the kernels in ${\rm{{mod}}}(\Gamma)$ and the inclusion functor from $\mathcal{C}$ to ${\rm{{mod}}}(\Gamma)$ is left exact.
Proof.
Let $f:X\longrightarrow Y$ be a morphism in $\mathcal{C}$. Let $k^{\prime}:K^{\prime}\longrightarrow X$ be the kernel of $f$ in $\mathcal{C}$. We consider the following two complexes and the natural isomorphism between them.
[…]
Since $\Gamma\in\mathcal{C}$ the upper complex is an exact sequence of abelian groups, so the lower complex is also exact in the category of abelian groups. Thus $k^{\prime}$ is also the kernel of $f$ in ${\rm{{mod}}}(\Gamma)$. ∎
Lemma 2.3.
Let $\Gamma$ be an Artin algebra and $n\geq 1,0\leq k\leq n+1$. The following statements are equivalents:
  • $(a)$
    $\mathrm{Ggldim}(\Gamma)\leq n+1$;
  • $(b)$
    $\mathrm{sub}^{k}({\mathrm{Gproj}}^{\leq n+1-k}(\Gamma))\subseteq{\mathrm{Gproj}}^{\leq n+1-k}(\Gamma)$;
Specially, the above statements are equivalent to each of the following:
  • $(b^{\prime})$
    ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under taking submodules;
  • $(b^{\prime\prime})$
    ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is closed under taking kernels in ${\rm{{mod}}}(\Gamma)$.
Proof.
$(a)\Rightarrow(b)$: We prove by induction on $k$. For $k=0$, this is trivial. For $k=1$, Let $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and $Y$ be a submodule of $X$. Form the short exact sequence $0\to Y\to X\to Z\to 0$. As ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$, and combining with ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$ by [Ho04a, Theorem 2.20].
Suppose $(b)$ holds for $k$ ($1\leq k\leq n$). Then for $Y\in\mathrm{sub}^{k+1}({\mathrm{Gproj}}^{\leq n-k}(\Gamma))$, we have an exact sequence
$0\longrightarrow Y\longrightarrow X^{0}\stackrel{{\scriptstyle d^{0}}}{{\longrightarrow}}X^{1}\longrightarrow\cdots\longrightarrow X^{k}$
where $X^{0},\dots,X^{k}\in{\mathrm{Gproj}}^{\leq n-k}(\Gamma)$. Let $K^{1}:={\rm{Coker}}(d^{0})$. Hence
$K^{1}\in\mathrm{sub}^{k}({\mathrm{Gproj}}^{\leq n-k}(\Gamma))\subseteq\mathrm{sub}^{k}({\mathrm{Gproj}}^{\leq n-k+1}(\Gamma))\subseteq{\mathrm{Gproj}}^{\leq n-k+1}(\Gamma)$
where the last inclusion holds by inductive assumption. Combining with ${\rm{Gpd}}({}_{\Gamma}X^{0})\leq n-k$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n-k$ by[Ho04a, Theorem 2.20].
$(b)\Rightarrow(a)$: Let $X\in{\rm{{mod}}}(\Gamma)$ and $k\geq 1$. Given a presentation of $X$ by finitely generated projective modules:
$P_{k-1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{k-2}\longrightarrow\cdots\longrightarrow X\longrightarrow 0,$
since $P_{0},\dots,P_{k-1}\in\mathrm{proj}(\Gamma)\subseteq{\mathrm{Gproj}}^{\leq n+1-k}(\Gamma)$, we have $K:=\ker(\beta)\in\mathrm{sub}^{k}({\mathrm{Gproj}}^{\leq n+1-k}(\Gamma))\subseteq{\mathrm{Gproj}}^{\leq n+1-k}(\Gamma)$. So ${\rm{Gpd}}({}_{\Gamma}X)\leq n+1$. This shows that ${\rm{Ggldim}}(\Gamma)\leq n+1$. ∎
Proposition 2.4.
Let $n$ be a positive integer and $\Gamma$ be an Artin algebra. Then $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$, ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$.
Before giving the proof, we would like to mention that torsion pairs play an important rule in the proof of Tachikawa’s theorem ([Ta74, Theorem 1]), Iyama’s theorem ([Iy04, Theorem 3.3]), and Kong’s theorem ([Ko14, Theorem 2.1]). We also follow this procedure and prove the result in four steps.
Proof.
Step I. We show that ${\rm{Ggldim}}(\Gamma)\leq n+1$.
Since ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is abelian, it has kernels which, by Lemma 2.2, coincident with the kernels in ${\rm{{mod}}}(\Gamma)$. Hence, ${\rm{Ggldim}}(\Gamma)\leq n+1$ follows from Lemma 2.3.
Step II. We show that ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under taking injective envelopes, i.e. $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ $\Rightarrow\mathbf{E}(X)\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
The idea is to show that ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is the torsion-free class of some hereditary torsion pair, and so is closed under taking injective envelopes [Di66, Theorem 2.9].
Obviously the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under extensions. By Lemma 2.3 ($b^{\prime}$), it is also closed under taking submodules. Hence, the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under taking extensions and submodules and so is the torsion-free class of some torsion pair $(\mathcal{T},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ in ${\rm{{mod}}}(\Gamma)$ [Di66, Theorem 2.3].
We now show that ${{}^{\perp_{0}}\Gamma}={{}^{{\perp}_{0}}{\mathrm{Gproj}}^{\leq n}(\Gamma)}=\mathcal{T}$. Let $X\in\mathcal{T}$. Then ${\rm{Hom}}_{\Gamma}(X,M)=0$ for $M=\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, we get $X\in{{}^{\perp_{0}}\Gamma}$. This shows that ${{}^{\perp_{0}}\Gamma}\supseteq\mathcal{T}$.
On the other hand, suppose that $X\in{{}^{\perp_{0}}}\Gamma$. By ${}^{\perp_{0}}\Gamma\subseteq{}^{\perp_{n}}\Gamma$, ${\rm{Ext}}^{i}_{\Gamma}(X,\Gamma)=0$ for every $0\leq i\leq n$. Hence, if $Y$ is a projective $\Gamma$-module, then ${\rm{Ext}}^{i}_{\Gamma}(X,Y)=0$ for every $0\leq i\leq n$. Thus, by a dimension shifting argument, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0={\rm{Ext}}^{1}_{\Gamma}(X,Y)$ for each $Y\in{\mathrm{Proj}}^{\leq n-1}(\Gamma)$. We also deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in{\mathrm{Proj}}^{\leq n}(\Gamma)$.
Also, if $Y$ is a Gorenstein projective $\Gamma$-module, then $Y$ is a submodule of a projective $\Gamma$-module and so ${\rm{Hom}}_{\Gamma}(X,Y)=0$.
Let now $Y$ be a $\Gamma$-module with ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$. Then, by [Ho04a, Theorem 2.10], there exists an exact sequence
$0\longrightarrow K\longrightarrow G\longrightarrow Y\longrightarrow 0$
such that $G$ is a Gorenstein projective module and $K$ is a $\Gamma$-module of projective dimension less than $n$. By applying ${\rm{Hom}}_{\Gamma}(X,-)$ to this exact sequence, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. We have shown that ${{}^{\perp_{0}}\Gamma}=\mathcal{T}$.
We now prove that ${{}^{\perp_{0}}}\Gamma$ is closed under taking submodules. For every short exact sequence
$0\longrightarrow X^{\prime}\longrightarrow X\longrightarrow X^{\prime\prime}\longrightarrow 0$
in ${\rm{{mod}}}(\Gamma)$, if $X\in{{}^{\perp_{0}}}\Gamma$, then $X^{\prime\prime}\in{{}^{\perp_{0}}}\Gamma$. By the exact sequence
$0\longrightarrow{\rm{Hom}}_{\Gamma}(X^{\prime\prime},\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(X,\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(X^{\prime},\Gamma)\longrightarrow{\rm{Ext}}^{1}_{\Gamma}(X^{\prime\prime},\Gamma)$
and the fact that ${\rm{Ext}}^{1}_{\Gamma}(X^{\prime\prime},\Gamma)=0$ (as ${}^{\perp_{0}}\Gamma\subseteq{}^{\perp_{n}}\Gamma$), ${\rm{Hom}}_{\Gamma}(X^{\prime},\Gamma)=0$. Hence, ${{}^{\perp_{0}}}\Gamma$ is also closed under taking submodules and so the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a hereditary torsion pair in ${\rm{{mod}}}(\Gamma)$.
This finishes the proof of Step 2.
Step III. We show that ${\rm{domdim}}(\Gamma)\geq n+1$.
By Step II, for every $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, $\mathbf{E}(X)\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. On the other hand, as mentioned in Subsection 1.11, by [Ho04b, Theorem 2.2], the projective dimension of $\mathbf{E}(X)$ equals to its Gorenstein projective dimension and so $\mathbf{E}(X)\in\mathrm{proj}^{\leq n}(\Gamma)$.
Hence, by the assumption, $\mathbf{E}(X)$ is a projective $\Gamma$-module. Since $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, $I^{0}:=\mathbf{E}(\Gamma)\in\mathrm{proj}(\Gamma)$ and so there is an exact sequence
$0\longrightarrow\Gamma\longrightarrow I^{0}\longrightarrow K_{0}\longrightarrow 0$
with $I^{0}\in\mathrm{proj}(\Gamma)$ and $K_{0}\in\mathrm{proj}^{\leq 1}(\Gamma)$. Hence, $K_{0}\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and so $I^{1}:=\mathbf{E}(K_{0})\in\mathrm{proj}(\Gamma)$. Now by applying this argument for finite steps, we deduce that there is an exact sequence
$0\longrightarrow{}_{\Gamma}\Gamma\longrightarrow I^{0}\longrightarrow I^{1}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\longrightarrow 0$
with $I^{i}\in\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$ for each $0\leq i\leq n$ and hence ${\rm{domdim}}\ \Gamma\geq n+1$. ∎
Propositions 2.1 and 2.4 could be enhanced to infinitely generated modules, that is, we could show the following result. Since its proof is analogous to that of Propositions 2.1 and 2.4, we only give a sketch of proof.
Proposition 2.5.
Let $n$ be a positive integer and $\Gamma$ be an Artin algebra. Then $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if and only if ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$ and ${{}^{\underline{\perp}_{0}}}\Gamma\subseteq{{}^{\underline{\perp}_{n}}}\Gamma$.
Proof.
$\Rightarrow$”. Let $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Let $X$ be an arbitrary injective $\Gamma$-module. By Facts 1.11 $(b)$, $X$ is a direct sum of indecomposable injective $\Gamma$-modules, the latter being finitely generated over an Artin algebra. From this fact, we deduce that each injective $\Gamma$-module is a direct sum of finitely generated injective $\Gamma$-modules.
If $X\in{\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)$, then by the previous paragraph, each indecomposable direct summand of $X$ lies in $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)$, which by Proposition 2.1 also falls into $\mathrm{proj}(\Gamma)$. Therefore, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$.
By a similar argument used to prove Lemma 1.1, one can also easily show that ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n}}}\Gamma$.
Let now $\mathcal{Q}$ be the maximal injective summand of $\Gamma$. The statement that ${\mathrm{Proj}}^{\leq n-1}(\Gamma)=\mathrm{Sub}^{2}(\mathcal{Q})$ can be proved similarly as in the proof of Proposition 2.1. By Lemma 1.6 (d), ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is abelian.
$\Leftarrow$”. For this direction, the proof is nearly the same as the proof of the implication $\Leftarrow$” in Proposition 2.4 except that in Step II, one needs to observe that the subcategory ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is also closed under taking direct products and to use [Di66, Theorem 2.3] to conclude. ∎
Now by combining Propositions 2.1, 2.4, and 2.5 we have the following theorem that can be considered as a higher generalization of Auslander-Tachikawa theorem.
Theorem 2.6.
Let $n$ be a positive integer and $\Gamma$ be an Artin algebra. Then the following statements are equivalent:
  • $(a)$
    $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra;
  • $(b)$
    ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$;
  • $(c)$
    ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$ and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$;
Based on the proof of Theorem 2.6 (in fact, Propositions 2.1, 2.4, and 2.5) we have the following results.
Corollary 2.7.
Let $n$ be a positive integer, $\Gamma$ be an $n$-minimal Auslander algebra, and $\mathcal{Q}$ be the maximal injective summand of $\Gamma$. Then the pairs $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ and $({{}^{\underline{\perp}_{0}}\Gamma},{\mathrm{GProj}}^{\leq n}(\Gamma))$ are hereditary torsion pairs.
Proof.
In Step 2 of the proof of Proposition 2.4, we showed that the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a hereditary torsion pair. A similar argument shows that the pair $({{}^{\underline{\perp}_{0}}\Gamma},{\mathrm{GProj}}^{\leq n}(\Gamma))$ is also a hereditary torsion pair. ∎
Corollary 2.8.
Let $n$ be a positive integer and $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Then the category of all $\Gamma$-modules (resp. finitely generated $\Gamma$-modules) with dominant dimension at least two is precisely the subcategory ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ (resp. ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$). Also, for every non-negative integer $i$, the full subcategories ${\mathrm{GProj}}^{\leq i}(\Gamma)$ and ${\mathrm{Gproj}}^{\leq i}(\Gamma)$ are closed under taking injective envelopes.
Proof.
For the first statement, it suffices to observe that a $\Gamma$-module (resp. finitely generated $\Gamma$-module) with dominant dimension at least two is precisely a module in $\mathrm{Sub}^{2}(\mathcal{Q})$ (resp. $\mathrm{sub}^{2}(\mathcal{Q})$).
To prove the second statement, note that by the proof of Theorem 2.6, the full subcategory ${\mathrm{GProj}}^{\leq n}(\Gamma)$ of ${\rm{Mod}}(\Gamma)$ is closed under taking injective envelopes. Thus for every $0\leq i\leq n$, if $X\in{\mathrm{GProj}}^{\leq i}(\Gamma)\subseteq{\mathrm{GProj}}^{\leq n}(\Gamma)$, then $\mathbf{E}(X)\in{\mathrm{GProj}}^{\leq n}(\Gamma)$. But ${\mathrm{GProj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$, and so $\mathbf{E}(X)\in{\mathrm{Proj}}(\Gamma)\subseteq{\mathrm{GProj}}^{\leq i}(\Gamma)$. For each $i\geq n+1$, since ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\mathrm{GProj}}^{\leq i}(\Gamma)={\rm{Mod}}(\Gamma)$ and so clearly is closed under taking injective envelopes.
Similarly, for every non-negative integer $i$, ${\mathrm{Gproj}}^{\leq i}(\Gamma)$ is also closed under taking injective envelopes. ∎
Kong used of the notion of submodule categories and characterized Artin algebras whose Gorenstein projective modules form an abelian category; see [Ko14, Corollary 2.3]. In the following, we will prove a similar property for higher Auslander-Gorenstein algebras.
Corollary 2.9.
Let $n$ be a positive integer and $\Gamma$ be an Artin algebra. If $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, then ${\mathrm{Gproj}}^{\leq n}(\Gamma)=\rm{sub}(\Gamma)$ and ${\mathrm{GProj}}^{\leq n}(\Gamma)=\rm{Sub}(\Gamma)$.
Proof.
We prove the first equality, the proof of the second being similar. Let $\Gamma$ be an $n$-Auslander algebra. If $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, by Corollary 2.8, $\mathbf{E}(X)\in\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$, so $X\in\rm{sub}(\Gamma)$. On the other hand, if a $\Gamma$-module $X$ is a submodule of a projective $\Gamma$-module, then ${\rm{Hom}}_{\Gamma}(T,X)=0$ for every $T\in{{}^{\perp_{0}}\Gamma}$. But, by Corollary 2.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair and so $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. Therefore, ${\mathrm{Gproj}}^{\leq n}(\Gamma)=\rm{sub}(\Gamma)$. ∎
A combination of our results and some classical results about torsion pairs gives us a better perspective about higher Auslander-Gorenstein algebras, see Corollaries 2.10 and 2.11.
Corollary 2.10.
Let $n$ be a positive integer and $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Then the following equivalent statements hold:
  • $(a)$
    ${{}^{\perp_{0}}\Gamma}$ is functorially finite;
  • $(b)$
    ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is functorially finite;
  • $(c)$
    ${{}^{\perp_{0}}\Gamma}=\mathrm{fac}(X)$ for some $X\in{\rm{{mod}}}(\Gamma)$;
  • $(d)$
    ${\mathrm{Gproj}}^{\leq n}(\Gamma)=\mathrm{sub}(Y)$ for some $Y\in{\rm{{mod}}}(\Gamma)$;
  • $(e)$
    $\mathcal{P}({{}^{\perp_{0}}\Gamma})$ is a tilting $(\Gamma/\mathrm{Ann}_{\Gamma}({{}^{\perp_{0}}\Gamma}))$-module;
  • $(f)$
    $\mathcal{I}({\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a cotilting $\Gamma$-module;
  • $(g)$
    ${{}^{\perp_{0}}\Gamma}=\mathrm{fac}(\mathcal{P}({{}^{\perp_{0}}\Gamma}))$;
  • $(h)$
    ${\mathrm{Gproj}}^{\leq n}(\Gamma)=\mathrm{sub}(\mathcal{I}({\mathrm{Gproj}}^{\leq n}(\Gamma)))$.
Proof.
By Corollary 2.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair. By [AR94, Proposition 5.8], ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is a functorially finite subcategory of ${\rm{{mod}}}(\Gamma)$ and so by Theorem 1.4 that we recalled it in the Subsection 1.5, all the statements hold and also are equivalent with each other. Note that $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and so $\mathrm{Ann}_{\Gamma}({\mathrm{Gproj}}^{\leq n}(\Gamma))=0$. But, by [Sm84, Lemma 0.2], $\mathrm{Ann}_{\Gamma}({{}^{\perp_{0}}\Gamma})\neq 0$. ∎
Corollary 2.11.
Let $n$ be a positive integer, $\Gamma$ be an $n$-Auslander algebra, and $\tau$ and $\tau^{-}$ be the Auslander-Reiten translations. Then the following statements hold.
  • (i)
    A $\Gamma$-module $X$ is ${\rm{Ext}}$-projective in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ if and only if $X$ is a projective $\Gamma$-module;
  • (ii)
    An indecomposable $\Gamma$-module $X$ is ${\rm{Ext}}$-projective in ${{}^{\perp_{0}}\Gamma}$ if and only if $\tau(X)\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, i.e. ${\rm{Gpd}}_{\Gamma}(\tau X)\leq n$;
  • (iii)
    An indecomposable $\Gamma$-module $X$ is ${\rm{Ext}}$-injective in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ if and only if $\tau^{-}(X)\in{{}^{\perp_{0}}\Gamma}$.
Proof.
By Corollary 2.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair and so, by [Ho82, Lemmas 1,2, and 3], all the statements hold; see also [AS81, Corollaries 3.4 and 3.7]. ∎
Corollary 2.12.
Let $n$ be a positive integer, $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra, and $t$ be the idempotent radical corresponding to the torsion pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$. Moreover, suppose $X\in{{}^{\perp_{0}}\Gamma}$ and $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ are indecomposable $\Gamma$-modules.Then the following statements hold.
  • (i)
    If $X$ is not ${\rm{Ext}}$-projective in ${{}^{\perp_{0}}\Gamma}$, then $t(\tau X)$ is indecomposable and for the Auslander-Reiten sequence $0\longrightarrow\tau X\longrightarrow E\longrightarrow X\longrightarrow o$, the induced sequence $0\longrightarrow t(\tau X)\longrightarrow t(E)\longrightarrow X\longrightarrow o$ is the Auslander-Reiten sequence in ${{}^{\perp_{0}}\Gamma}$;
  • (ii)
    If $Y$ is not ${\rm{Ext}}$-injective in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$, then $\tau^{-}Y/t(\tau^{-}Y)$ is indecomposable and for the Auslander-Reiten sequence $0\longrightarrow Y\longrightarrow E\longrightarrow\tau^{-}Y\longrightarrow o$, the induced sequence $0\longrightarrow Y\longrightarrow E/t(E)\longrightarrow\tau^{-}Y/t(\tau^{-}Y)\longrightarrow o$ is the Auslander-Reiten sequence in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$;
Proof.
By [Ho82, Lemmas 2 and 3], all the statements hold; see also [AS81, Corollaries 3.4 and 3.7]. ∎
3. Selfinjective and $\tau_{n}$-selfinjective Algebras
As an application of our results in Section 2, in this section we study selfinjective and $\tau_{n}$-selfinjective algebras and characterize them in a categorical sense.
We first have the following result that gives a categorical description for when an $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is self-injective. Recall that an Artin algebra $\Gamma$ is said to be selfinjective if it is injective as a left $\Gamma$-module.
Proposition 3.1.
Let $n$ be a positive integer and an Artin algebra $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Then the following statements are equivalent:
  • $(a)$
    $\Gamma$ is selfinjective;
  • $(b)$
    ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category;
  • $(c)$
    ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$;
  • $(d)$
    ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is an abelian category;
  • $(e)$
    ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n+1}}\Gamma}$.
Proof.
$(a)\Leftrightarrow(b)$. Let ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ be an abelian category and $\Gamma$ be non-selfinjective. Then ${\rm{Ggldim}}(\Gamma)$ $=n+1$, so there exists a finitely generated $\Gamma$-module $X$ such that ${\rm{Gpd}}({}_{\Gamma}X)=n+1$. Let
$0\to K\stackrel{{\scriptstyle\kappa}}{{\to}}P_{0}\stackrel{{\scriptstyle\varepsilon}}{{\longrightarrow}}X\longrightarrow 0$
be a short exact sequence with $P_{0}\in\mathrm{proj}(\Gamma)$. Then $K\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$. If $\pi:P_{0}\longrightarrow Y$ is the cokernel of the natural monomorphism $\kappa$ in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$, then the sequence
$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}Y$
is an exact sequence in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ as ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category. Since $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, the sequence
$0\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,K)\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,P_{0})\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,Y)$
of abelian groups is also exact and so the sequence
$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}Y$
is also exact in ${\rm{{mod}}}(\Gamma)$. Thus the induce map $\overline{\pi}:X\simeq P_{0}/K\longrightarrow Y$ is a monomorphism. Hence, $X$ is a submodule of $Y$ and so there exists a short exact sequence
$0\longrightarrow X\longrightarrow Y\longrightarrow Z\longrightarrow 0$
in ${\rm{{mod}}}(\Gamma)$ such that ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$. Since ${\rm{Ggldim}}(\Gamma)=n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$ and so ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, which is a contradiction.
We have shown that if ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category, then $\Gamma$ is selfinjective. The other direction is trivial.
$(a)\Leftrightarrow(c)$. Assume that ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$. Suppose that $\Gamma$ has Gorenstein global dimension $n+1$. We first show that ${{}^{\perp_{n+1}}\Gamma}\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$. Let $X\in{}^{\perp_{n+1}}\Gamma$. Then ${\rm{Ext}}_{\Gamma}^{n+1}(X,\Gamma)=0$. Since ${\rm{Ggldim}}(\Gamma)=n+1$, ${\rm{Gpd}}({}_{\Gamma}X)\leq n$ and hence ${}^{\perp_{n}}\Gamma\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$. Then ${{}^{\perp_{0}}\Gamma}\subseteq{}^{\perp_{n+1}}\Gamma\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$. By Corollary 2.7, $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair in ${\rm{{mod}}}(\Gamma)$ and so ${{}^{\perp_{0}}\Gamma}=0$. Therefore, ${\mathrm{Gproj}}^{\leq n}(\Gamma)={\rm{{mod}}}(\Gamma)$, which contradicts the assumption ${\rm{Ggldim}}(\Gamma)=n+1$. So ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$ implies that $\Gamma$ is selfinjective.
The other direction is trivial.
$(a)\Leftrightarrow(d)$. The proof is similar to that of $(a)\Leftrightarrow(b)$.
$(a)\Leftrightarrow(e)$. The proof is similar to that of $(a)\Leftrightarrow(c)$. ∎
Now we can give a categorical description for the selfinjectivity of an Artin algebra.
Corollary 3.2.
Let $\Gamma$ be an Artin algebra. Then $\Gamma$ is selfinjective if and only if the categories ${\mathrm{Gproj}}(\Gamma)$ and ${\mathrm{Gproj}}^{\leq 1}(\Gamma)$ are abelian categories if and only if the categories ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{GProj}}^{\leq 1}$ $(\Gamma)$ are abelian categories.
Proof.
Let the categories ${\mathrm{Gproj}}(\Gamma)$ and ${\mathrm{Gproj}}^{\leq 1}(\Gamma)$ be abelian categories. By [Ko14, Theorem 2.1], $\Gamma$ is an Auslander algebra and hence by Proposition 3.1, $\Gamma$ is selfinjective. For the second equivalence, the proof is similar. ∎
We now turn our attention to $\tau_{n}$-selfinjective algebras. Let us recall their definition.
Definition 3.3 ($\tau_{n}$-selfinjective algebras).
Let $\Lambda$ be an Artin algebra and $\mathcal{X}$ be a full subcategory of ${\rm{{mod}}}(\Lambda)$. The subcategory $\mathcal{X}$ is cogenerating if for every object $M\in{\rm{{mod}}}(\Lambda)$ there exist an object $X\in\mathcal{X}$ and a monomorphism $M\longrightarrow X$. The concept of generating subcategories is defined dually. We also call $\mathcal{X}$ $n$-rigid if ${\rm{Ext}}^{i}_{\Lambda}(\mathcal{X},\mathcal{X})=0$ for each $0<i<n$.
Let now $\tau_{n}$ and $\tau_{n}^{-}$ be the $n$-Auslander-Reiten translations; see [Iy07a, Subsection 1.4] or [Iy11, Definition 1.1]. A generating-cogenerating $n$-rigid subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Lambda)$ is called $n$-precluster tilting subcategory if it is functorially finite, $\tau_{n}(\mathcal{X})\subseteq\mathcal{X}$, and $\tau_{n}^{-}(\mathcal{X})\subseteq\mathcal{X}$ [IS18, Definition 3.2]. Moreover, if $\mathcal{X}$ admits an additive generator $M$, we say that $\mathcal{X}$ is a finite $n$-precluster tilting subcategory and $M$ is an $n$-precluster tilting module. By following Iyama-Solberg [IS18, Definition 3.4], Artin algebras which have $n$-precluster tilting modules are also called $\tau_{n}$-selfinjective.
[…]
Now, by combining Iyama-Solberg correspondence and our results (Theorem 2.6), we can prove a result that gives us a new characterization of the class of Artin algebras having $n$-precluster tilting modules, i.e. $\tau_{n}$-selfinjective algebras. We need the following lemma which is of independent interest and follows from Lemmas 1.5 and 1.6.
Lemma 3.4.
Let $n$ be a positive integer and $\Lambda$ be an Artin algebra. Let $M$ be a $\Lambda$-module which is generating and cogenerating. Denote $\Gamma=\mathrm{End}_{\Lambda}(M)^{{\rm{op}}}.$ Let $\mathcal{Q}:=\mathrm{Hom}_{\Lambda}(M,\mathbf{D}\Lambda)\cong\mathbf{D}M$ and $I=\mathrm{Hom}_{\Gamma^{{\rm{op}}}}(\mathbf{D}\mathcal{Q},\Gamma)$. Then the following statements hold:
  • (a)
    The functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$, which further restricts to another equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$, so $\mathrm{add}(\mathcal{Q})$ $=\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$.
  • (b)
    The right $\Gamma$-module $M_{\Gamma}$ is projective; $I\cong\mathrm{Hom}_{\Gamma}(M,\Gamma)\cong\nu^{-1}\mathcal{Q}$, where $\nu^{-1}$ is the inverse Nakayama functor, so $\Lambda\cong\mathrm{End}_{\Gamma}(\mathcal{Q})^{{\rm{op}}}\cong\mathrm{End}_{\Gamma}(I)^{{\rm{op}}}.$
  • (c)
    The three functors $M\otimes_{\Gamma}-,\mathrm{Hom}_{\Gamma}(I,-)$ and $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q})$ from ${\rm{{mod}}}(\Gamma)$ to ${\rm{{mod}}}(\Lambda)$ are naturally isomorphic. So there exists an adjoint triple
    ${{\rm{{mod}}}(\Lambda)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{Hom}_{\Lambda}(M,-)}$$\scriptstyle{I\otimes_{\Lambda}-}$${{\rm{{mod}}}(\Gamma)\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{M\otimes_{\Gamma}-\simeq\mathrm{Hom}_{\Gamma}(I,-)}$
  • (d)
    The functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ is fully faithful with essential image $\mathrm{sub}^{2}(\mathcal{Q})$, so $\mathrm{sub}^{2}(\mathcal{Q})\simeq{\rm{{mod}}}(\Lambda)$ is an abelian category.
  • (e)
    The functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{Mod}}(\Lambda)\to{\rm{Mod}}(\Gamma)$ is fully faithful with essential image $\mathrm{Sub}^{2}(\mathcal{Q})$, so $\mathrm{Sub}^{2}(\mathcal{Q})\simeq{\rm{Mod}}(\Lambda)$ is an abelian category.
Proof.
The statement $(a)$ can be proved as in [ARS, chapter 5, Lemma 5.3]. For $(b)$, the fact that $M_{\Gamma}$ is projective follows from the duality $\mathrm{Hom}_{\Lambda}(-,M)$ between $\mathrm{add}(M)$ and $\mathrm{proj}(\Gamma^{op})$ and the fact that ${}_{\Lambda}\Lambda\in\mathrm{add}(M)$.
Other assertions deduce from $(a)$ and the fact that $\nu^{-1}$ establishes an equivalence between (finitely generated) projective modules and injective modules.
For $(c)$, the two functors $\mathrm{Hom}_{\Gamma}(I,-)$ and $D\mathrm{Hom}_{\Gamma}(-,Q)$ from ${\rm{{mod}}}(\Gamma)$ to ${\rm{{mod}}}(\Lambda)$ are right exact (in fact exact), so by Watts’ theorem, one need to verify that they coincident with $M\otimes_{\Gamma}-$ on the left regular module ${}_{\Lambda}\Lambda$, which is obvious by $Q=\mathrm{Hom}_{\Lambda}(M,D\Lambda)\simeq D(M)$ and $I\simeq\mathrm{Hom}_{\Gamma}(M,\Gamma)\simeq\nu^{-1}Q$. The existence of the adjoint triple follows from the usual tensor-Hom adjunction.
By $(a)$, the functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ restricts to an equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{add}(Q)$, with quasi-inverse $D\mathrm{Hom}_{\Gamma}(-,Q)$.
By the first lemma of [Au71, Chapter III, Section 4, Page 48]. this equivalence can be extended to an equivalence ${\rm{{mod}}}(\Lambda)\simeq\mathrm{sub}^{2}(\mathcal{Q})$, as $Q$ is injective. This proves $(d)$.
For $(d^{\prime})$, by Fact 1.11 $(a)(b)$, the equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{add}(Q)$ can be extended to ${\mathrm{Inj}}(\Lambda)\simeq\mathrm{Add}(Q)$. Again the latter can be extended to ${\rm{Mod}}(\Lambda)\simeq\mathrm{Sub}^{2}(\mathcal{Q})$. ∎
Theorem 3.5.
Let $\Lambda$ be an Artin algebra and $n$ be a positive integer. Then the following are equivalent:
  • $(a)$
    $\Lambda$ is $\tau_{n}$-selfinjective, i.e. it has $n$-precluster tilting modules;
  • $(b)$
    ${\rm{{mod}}}(\Lambda)$ is equivalent to ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$, where $\Gamma$ is an Artin algebra such that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$;
  • $(c)$
    ${\rm{Mod}}(\Lambda)$ is equivalent to ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$, where $\Gamma$ is an Artin algebra such that ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$ and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n}}\Gamma}$.
Proof.
$(a)\Rightarrow(b)$ follows from Lemma 3.4 $(d)$ and Proposition 2.1.
$(b)\Rightarrow(a)$. By Proposition 2.4, $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra and so by Iyama-Solberg correspondence [IS18, Theorem 4.5], there is an Artin algebra $\Lambda^{\prime}$ with $n$-precluster tilting $\Lambda^{\prime}$-module $M^{\prime}$ such that ${\rm{{mod}}}(\Lambda^{\prime})\simeq{\mathrm{Gproj}}^{\leq n-1}(\Gamma)\simeq{\rm{{mod}}}(\Lambda)$. Hence, $\Lambda$ and $\Lambda^{\prime}$ are also Morita equivalent.
The equivalence $(a)\Leftrightarrow(c)$ can be proved similarly with Lemma 3.4 (d) replaced by Lemma 3.4 (e) and Proposition 2.4 by Proposition 2.5. ∎
Remark 3.6.
Lemmas 1.6 and 3.4, allowed us to generalize our results from finitely generated modules to all modules, see Theorems 2.6 and 3.5. Of course, we are not limited to these lemmas and can follow other ways.
Let $\mathcal{Q}$ be the maximal injective summand of Artin algebra $\Gamma$. Then $\mathcal{Q}$ is product-complete. Recall that a $\Gamma$-module $\mathcal{Q}$ is product-complete, if ${\rm{Add}}(\mathcal{Q})={\rm Prod}(\mathcal{Q})$, where $\mathrm{Prod}(\mathcal{Q})$ is the smallest subcategory of ${\rm{Mod}}(\Gamma)$ containing $Q$ and closed under products and direct summands.
By [KS98, Proposition 3.5], there is a (locally Noetherian) Grothendieck category $\mathcal{A}$ and a functor $F:\mathcal{A}\longrightarrow{\rm{Mod}}(\Gamma)$ such that $F$ induces an equivalence between ${\mathrm{Inj}}(\mathcal{A})$ and ${\rm{Add}}(\mathcal{Q})$. By the proof of [KS98, Proposition 3.5], the functor $F$ is a composition of the right adjoint of a quotient functor with an exact functor and so $F$ is left exact. Hence, $F$ induces an equivalence between $\mathcal{A}$ and $\mathrm{Sub}^{2}(\mathcal{Q})$. On the other hand, by a similar argument like the proof of Proposition 2.1, ${\mathrm{GProj}}^{\leq n-1}(\Gamma)=\mathrm{Sub}^{2}(\mathcal{Q})$ and so ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category.
4. Cotorsion Pairs
As mentioned in the introduction, this section is devoted to study the relation between higher Auslander-Gorenstein and cotorsion pairs and we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS].
Let $n$ and $i$ be a non-negative integer. For an $n$-minimal Auslander-Gorenstein algebra $\Gamma$, some results by Martinez Villa show that the pair
$(\rm{dom}^{\geq i}(\Gamma),\mathrm{inj}^{\leq i}(\Gamma))$
is a hereditary cotorsion pair in ${\rm{{mod}}}(\Gamma)$, where $\rm{dom}^{\geq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules having the dominant dimension at least $i$ and $\mathrm{inj}^{\leq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules with injective dimension at most $i$, see [Ma92, Lemma 3, Proposition 5, and the corollary before it].
A result by Marczinzik shows that ${\mathrm{Gproj}}^{\leq n+1-i}(\Gamma)=\rm{dom}^{\geq i}(\Gamma)$ (see, [Ma18, Theorem 2.1]) and so, by [ET01, Theorem 10] or Salce’s lemma and [Ma92, Proposition 7], the pair
$({\mathrm{Gproj}}^{\leq n+1-i}(\Gamma),\mathrm{inj}^{\leq i}(\Gamma))$
is a hereditary complete cotorsion pair in ${\rm{{mod}}}(\Gamma)$ for every integer $i=0,1,\cdots,n+1$.
The following results describe the kernel of the above mentioned cotorsion pairs and relates them to the notion of tilting theory.
Proposition 4.1.
Let $n$ be a positive integer, $10\leq k\leq n+$ 1, $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra; and $Q$ be its maximal projective-injective summand. Then $T_{k}:=\Omega^{-k}(\Gamma)\oplus Q$ is a $k$-tilting module and ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Also for all $0\leq j\leq n+1-k$,
$\mathrm{sub}^{j}(T_{k})=\rm{dom}^{\geq j}(\Gamma)={\mathrm{Gproj}}^{\leq n+1-j}(\Gamma).$
Especially, for every $10\leq i\leq n+$ 1,
$\mathrm{sub}^{i}(T_{n+1-i})=\rm{dom}^{\geq i}(\Gamma)={\mathrm{Gproj}}^{\leq n+1-i}(\Gamma).$
Proof.
Let
$0\longrightarrow{}_{\Gamma}\Gamma\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\longrightarrow 0$
be the minimal injective resolution of $\Gamma$. Since $I^{0},I^{1},\cdots,I^{n}\in{\rm{add}}(\mathcal{Q})$, clearly, $\Gamma$ has an ${\rm{add}}(T_{k})$-copresentation of length $k$ and for every $10\leq k\leq n+$ 1, $\Omega^{-k}(\Gamma)\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Hence, for every $11\leq l\leq n+$ 1,
${\rm{Ext}}^{l}_{\Gamma}(T_{k},T_{k})\simeq{\rm{Ext}}^{l}_{\Gamma}(\Omega^{-k}(\Gamma),\Omega^{-k}(\Gamma))\simeq{\rm{Ext}}^{l}_{\Gamma}(\Omega^{k}(\Omega^{-k}(\Gamma)),\Gamma)={\rm{Ext}}^{l}_{\Gamma}(\Gamma,\Gamma)=0.$
Therefore, $T_{k}$ is a $k$-tilting module; see [HU96, Page 448] or [AC01, Section 2] for the definition.
We now show that for every $0\leq j\leq n+1-k$, $\mathrm{sub}^{j}(T_{k})=\rm{dom}^{\geq j}(\Gamma)={\mathrm{Gproj}}^{\leq n+1-j}(\Gamma).$
For every $0\leq j\leq n+1-k$, $T_{k}\in{\mathrm{Gproj}}^{\leq n+1-j}(\Gamma)$ and so by Lemma 2.3,
$\mathrm{sub}^{j}(T_{k})\subseteq\mathrm{sub}^{j}({\mathrm{Gproj}}^{\leq n+1-j}(\Gamma))\subseteq{\mathrm{Gproj}}^{\leq n+1-j}(\Gamma).$
Conversely, ${\mathrm{Gproj}}^{\leq n+1-j}(\Gamma)=\rm{dom}^{\geq j}(\Gamma)=\mathrm{sub}^{j}(\mathcal{Q})\subseteq\mathrm{sub}^{j}(T_{k}).$
To complete the proof, we must show that for every $10\leq k\leq n+$ 1, ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. Clearly, ${\rm{add}}(T_{k})\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so we only need to prove the converse.
Let $X\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$. By [Ho04b, Theorem 2.2], ${\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)=\mathrm{proj}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so ${\rm{pd}}(_{\Gamma}X)\leq k$. Hence, by [AC01, Lemma 2.2 $(a)$], ${}^{\perp}(T_{k}^{\perp})\subseteq\mathrm{proj}^{\leq k}(\Gamma)\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)$ and so
$X\in\mathrm{inj}^{\leq n+1-k}(\Gamma)=({\mathrm{Gproj}}^{\leq k}(\Gamma))^{\perp}\subseteq(^{\perp}(T_{k}^{\perp}))^{\perp}=T_{k}^{\perp}.$
On the other hand, by [AC01, Lemmas 2.3 $(a)$ and 2.4 $(a)$], for every $Y\in T_{k}^{\perp}$, there is a short exact sequence
$0\longrightarrow K_{0}\longrightarrow T^{\prime}_{0}\longrightarrow Y\longrightarrow o,$
where $T^{\prime}_{0}\in{\rm{add}}(T_{k})$ and $K_{0}\in T_{k}^{\perp}$. By applying this argument for finite steps, there is an exact sequence
$0\longrightarrow K_{k-1}\longrightarrow T^{\prime}_{k-1}\longrightarrow\cdots\longrightarrow T^{\prime}_{0}\longrightarrow Y\longrightarrow o,$
where $T^{\prime}_{0},\cdots,T^{\prime}_{k-1}\in{\rm{add}}(T_{k})$. Clearly $K_{k-1}\in X^{\perp_{[k+1,\infty)}}$ and $T^{\prime}_{0},\cdots,T^{\prime}_{k-1}\in{\rm{add}}(T_{k})\subseteq X^{\perp}$. Hence, a dimension shifting argument shows that $Y\in X^{\perp}$. Therefore, $X\in{{}^{\perp}(T_{k}^{\perp})}$. But by [AC01, Lemmas 2.3 $(a)$ and 2.4 $(a)$] there is a short exact sequence
$0\longrightarrow K\longrightarrow T\longrightarrow X\longrightarrow o,$
where $T\in{\rm{add}}(T_{k})$ and $K\in T_{k}^{\perp}$. Now, by applying functor ${\rm{Hom}}_{\Gamma}(X,-)$, we deduce that $X\in{\rm{add}}(T_{k})$. ∎
Among the above mentioned cotorsion pairs, one of them has a very nice characteristic and its cotorsion class is a torsion-free class of a hereditary torsion pair. This fact relates higher Auslander-Gorenstein Algebras to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS] and gives us a nice description of them.
We first recall some definitions and results in this direction and then by using them we will establish two equivalences for $n$-minimal Auslander-Gorenstein algebras.
Definition 4.2 (Torsion Cotorsion Triple).
A torsion cotorsion triple in an abelian category $\mathcal{A}$ is a triple of subcategories $(\mathcal{T},\mathcal{F},\mathcal{D})$ such that the pair $(\mathcal{T},\mathcal{F})$ is a torsion pair and the pair $(\mathcal{F},\mathcal{D})$ is a cotorsion pair, see [BBOS, Page 29 before Theorem 2.33].
Theorem 4.3 (Dual of [BBOS, Theorem 2.35]).
Let $\Gamma$ be an Artin algebra and $(\mathcal{T},\mathcal{F},\mathcal{D})$ be a torsion cotorsion triple in ${\rm{{mod}}}(\Gamma)$. Then the inverse of Auslander–Reiten translation defines the following equivalence.
$\mathcal{T}\simeq\frac{\mathcal{D}}{\mathrm{inj}(\Gamma)}.$
Proof.
This theorem is the dual version of [BBOS, Theorem 2.35] and for the convenience of the reader we give the proof.
By [ARS, Proposition IV.1.9], we know that the inverse of Auslander-Reiten translation $\tau^{-}$ defines an equivalence
[…]
Clearly, $\frac{\mathcal{D}}{\mathrm{inj}(\Gamma)}$ is a full subcategory of $\overline{{\rm{{mod}}}}(\Gamma)$ and one can easily see that $\mathcal{T}$ is a full subcategory of $\underline{{\rm{{mod}}}}(\Gamma)$. In fact, the class $\mathcal{F}$ contains projectives and so ${\rm{Hom}}_{\Gamma}(\mathcal{T},\mathrm{proj}(\Gamma))=0$. Thus there is no zero maps between objects in $\mathcal{T}$ factor through projective modules.
Now, to complete the proof, we show that a module $X$ is in $\mathcal{D}$ if and only if $\tau^{-}X$ is in $\mathcal{T}$. We observe that
$X\in\mathcal{D}\Longleftrightarrow{\rm{Ext}}^{1}_{\Gamma}(\mathcal{F},X)=0\Longleftrightarrow{\rm{Hom}}_{\Gamma}(\tau^{-}X,\mathcal{F})=0\Longleftrightarrow\tau^{-}X\in\mathcal{T},$
where the middle equivalence follows from Auslander-Reiten formula because the injective dimension of every element of $\mathcal{D}$ by Bauer-Botnan-Oppermann-Steen Correspondence is at most one, see [SY11, Corollary III.6.4] and [BBOS, Theorem 2.34]. ∎
As an immediate result of these results, we have the following equivalences.
Corollary 4.4.
Let $n$ be a positive integer and $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Then we have the following equivalences.
${{}^{\perp_{0}}\Gamma}\simeq\frac{\mathrm{inj}^{\leq 1}(\Gamma)}{\mathrm{inj}^{\leq 1}(\Gamma)\bigcap\mathrm{proj}^{\leq n}(\Gamma)}\simeq\frac{\mathrm{inj}^{\leq 1}(\Gamma)}{\mathrm{inj}(\Gamma)}.$
Proof.
Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, the triple $(^{\perp_{0}}\Gamma,{\mathrm{Gproj}}^{\leq n}(\Gamma),\mathrm{inj}^{\leq 1}(\Gamma))$ is a torsion cotorsion triple, see Corollary 2.7 and the arguments at the beginning of this section. Hence, by [BBOS, Theorem 2.33], we have the first equivalence. That is
${{}^{\perp_{0}}\Gamma}\simeq\frac{\mathrm{inj}^{\leq 1}(\Gamma)}{\mathrm{inj}^{\leq 1}(\Gamma)\bigcap\mathrm{proj}^{\leq n}(\Gamma)}.$
The second equivalence also follows from Theorem 4.3. ∎
5. Higher Auslander Algebras
As higher Auslander algebras are special cases of higher Auslander-Gorenstein algebras and are of independent interest, in this short section we restrict our attention to these algebras and rewrite some of results of previous sections for them.
Let $n$ be a positive integer. Recall that an Artin algebra $\Gamma$ is an $n$-Auslander algebra if its global dimension is at most $n+1$ and its dominant dimension is at least $n+1$, that is,
${\rm{gldim}}(\Gamma)\leq n+1\leq\mathrm{domdim}(\Gamma).$
$\Gamma$ is also called a higher Auslander algebra, if there is a positive integer $n$ such that $\Gamma$ is an $n$-Auslander algebra. Iyama introduced these algebras as a generalization of classical Auslander algebras [Iy11, Page 3].
As a result of Theorem 2.6, we have the following result for higher Auslander algebras.
Theorem 5.1 (Higher Auslander-Tachikawa Theorem).
Let $n$ be a positive integer and $\Gamma$ be an Artin algebra. Then the following statements are equivalent.
  • $(a)$
    $\Gamma$ is an $n$-Auslander algebra;
  • $(b)$
    $\mathrm{proj}^{\leq n-1}(\Gamma)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$, and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$;
  • $(c)$
    ${\mathrm{Proj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$, and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n}}\Gamma}$. ∎
Remark 5.2.
Let $\Gamma$ be an Artin algebra and $n$ be a positive integer. Then we have the following statements:
  • $(a)$
    If $\mathrm{proj}(\Gamma)$ is an abelian category, then
    $\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)\ \mathrm{and}\ {{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{1}}\Gamma};$
  • $(b)$
    If ${\mathrm{Proj}}(\Gamma)$ is an abelian category, then
    ${\mathrm{Proj}}^{\leq 1}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)\ \mathrm{and}\ {{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{1}}\Gamma}.$
Hence, Theorem 5.1 is precisely a generalization of Auslander-Tachikawa theorem (Theorem 1) and also provide a new proof for it.
Let us prove $(a)$, that of $(b)$ being similar. In fact, assume that $\mathrm{proj}(\Gamma)$ is an abelian category. By Step 1 of the proof of Proposition 2.4, we see that $\mathrm{proj}(\Gamma)$ has the same kernels as ${\rm{{mod}}}(\Gamma)$, so $\mathrm{gldim}(\Gamma)\leq 2$.
Let $X\in\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)$. Since $\mathrm{proj}(\Gamma)$ is an abelian category, by the proof of Proposition 3.1, there exists a projective $\Gamma$-module $Y$ such that $X$ is a submodule of $Y$. But $X$ is an injective $\Gamma$-module and so it is a direct summand of $Y$. Therefore, $X$ is a projective $\Gamma$-module. This shows that $\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$.
Let now $X\in{{}^{{\perp}_{0}}}\Gamma$. Then, there is an exact sequence
$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{0}\longrightarrow X\longrightarrow 0$
where $P_{0},P_{1}\in\mathrm{proj}(\Gamma)$. Since $X\in{{}^{{\perp}_{0}}}\Gamma$, ${\rm{Hom}}_{\Gamma}(X,P)=0$ for every finitely generated projective $\Gamma$-module $P$, $\beta$ is a surjective map in $\mathrm{proj}(\Gamma)$. Since $\mathrm{gldim}(\Gamma)\leq 2$, $K$ is projective. Thus the sequence $0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{0}\longrightarrow 0$ is exact in $\mathrm{proj}(\Gamma)$ and so by the exact sequence of abelian groups
$0\longrightarrow{\rm{Hom}}_{\Gamma}(P_{0},\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(P_{1},\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(K,\Gamma)$
one can easily see that ${\rm{Ext}}^{1}_{\Gamma}(X,\Gamma)=0$. Thus the condition ${{}^{{\perp}_{0}}}\Gamma\subseteq{{}^{{\perp}_{1}}}\Gamma$ holds.
Remark 5.3.
For $n\geq 2$, the conditions $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ in Theorem 5.1(b) are necessary. In fact, take an Artin algebra of global dimension $n-1$ which is not an $n$-Auslander algebra. Then $\mathrm{proj}^{\leq n-1}(\Gamma)={\rm{{mod}}}(\Gamma)$ is an abelian category, but by Theorem 5.1, the conditions that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ or ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ don’t hold.
Remark 5.4.
Assume that $\Gamma$ is an $n$-Auslander algebra such that ${\rm{gldim}}\ \Gamma=n+1={\rm{domdim}}\ \Gamma$. By Theorem 5.1, $\mathrm{proj}^{\leq n-1}(\Gamma)$ is an abelian category and $\mathrm{proj}^{\leq i}(\Gamma)$ is not an abelian category for each $i=0,\dots,n-2$. Hence, Theorem 5.1 can be used to study the abelianness of some subcategories of ${\rm{{mod}}}(\Gamma)$.
Now, like Theorem 3.5, by combining Auslander-Iyama correspondence and our results (Theorem 5.1), we have the following result that gives us a new characterization of the class of Artin algebras having $n$-cluster tilting modules, i.e. $n$-representation finite algebras. We first recall the definition of $n$-representation finite algebras.
Definition 5.5 ($n$-representation finite).
Let $\Lambda$ be an Artin algebra and $\mathcal{M}$ be a full subcategory of ${\rm{{mod}}}(\Lambda)$. Recall that $\mathcal{M}$ is an $n$-cluster-tilting subcategory of ${\rm{{mod}}}(\Lambda)$ if it is functorially finite in ${\rm{{mod}}}(\Lambda)$ and
[…]
Moreover, if $\mathcal{M}$ admits an additive generator $M$, i.e. $\mathcal{M}={\rm{add}}(M)$, we say that $\mathcal{M}$ is a finite $n$-cluster tilting subcategory and $M$ is an $n$-cluster tilting module.
[…]
Theorem 5.6.
Let $\Lambda$ be an Artin algebra and $n$ be a positive integer. Then the following are equivalent.
  • $(a)$
    $\Lambda$ is $n$-representation finite, i.e. it has an $n$-cluster tilting module;
  • $(b)$
    ${\rm{{mod}}}(\Lambda)$ is equivalent to $\mathrm{proj}^{\leq n-1}(\Gamma)$, where $\Gamma$ is an Artin algebra such that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$;
  • $(c)$
    ${\rm{Mod}}(\Lambda)$ is equivalent to ${\mathrm{Proj}}^{\leq n-1}(\Gamma)$, where $\Gamma$ is an Artin algebra such that ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$ and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n}}\Gamma}$. ∎
The following example gives a better perspective about some results of the paper. In the following, for every vertex $i$ of the given quiver $Q$, the corresponding simple, projective, and injective object will be deonted by $S(i)$, $P(i)$, and $I(i)$, respectively. For more details about quiver representations, see for example [SY11].
Example 5.7.
Let $\Gamma=KQ/I$ be an algebra with the quiver $Q$:
{tikzpicture}
and $I$ is the admissible ideal of $KQ$ generated by paths $\alpha_{i+1}\alpha_{i}$ for every $1\leq i\leq n$. Then:
  1. (1)
    By [SY11, Theorem I.10.5], the Nakayama algebra $\Gamma$ is of finite representation type with $32n+$ 3 indecomposable $\Gamma$-modules. The following table gives us more information about their homological dimensions.
    Ind. Modules    pd    id
    $P(1)$ 00 $n+1$
    $P(2)=I(3)$ 00 00
    $\vdots$ $\vdots$ $\vdots$
    $P(n)=I(n+1)$ 00 00
    $P(n+1)=I(2)$ 00 00
    $I(1)$ $n+1$ 00
    $S(1)$ $n+1$ $n+1$
    $S(2)$ $n$ $1$
    $\vdots$ $\vdots$ $\vdots$
    $S(j)\ (2\leq j\leq n+1)$ $n+2-j$ $j-1$
    $\vdots$ $\vdots$ $\vdots$
    $S(n+1)$ $1$ $n$
  2. (2)
    The global dimension of $\Gamma$ is $n+1$ and $P(i)=I(i+1)$ for $2\leq i\leq n$ and $P(n+1)=I(2)$.
  3. (3)
    The minimal injective resolution of $\Gamma$ is as follows:
    $0\rightarrow\Gamma\rightarrow(\bigoplus_{i=2}^{n+1}I(i))\oplus I(2)\rightarrow I(2)\rightarrow I(n+1)\rightarrow\cdots\rightarrow I(2)\rightarrow I(1)\rightarrow 0$.
    Hence $\Gamma$ is $(n+1)$-Gorenstein.
  4. (4)
    The dominant dimension of $\Gamma$ is $n+1$ and so $\Gamma$ is an $n$-Auslander algebra.
  5. (5)
    The full subcategory $\mathrm{proj}^{\leq n}(\Gamma)$ of ${\rm{{mod}}}(\Gamma)$ is generated by indecomposable $\Gamma$-modules $P(1),\cdots,$ $P(n+1),S(2),\cdots,S(n+1)$. Hence, by Corollary 2.7, the torsion class ${{}^{\perp_{0}}\Gamma}$ can have only $S(1)$ and $I(1)$ as the idecomposable objects. But ${\rm{Hom}}_{\Gamma}(I(1),P(n))\neq 0$ and so ${{}^{\perp_{0}}\Gamma}={\rm{add}}(S(1))$.
  6. (6)
    By Corollary 2.11 $(i)$, $P(1),\cdots,P(n+1)$ are all indecomposable ${\rm{Ext}}$-projective objects in $\mathrm{proj}^{\leq n}(\Gamma)$, i.e. $\mathcal{P}(\mathrm{proj}^{\leq n}(\Gamma))=\bigoplus_{i=1}^{n+1}P(i)$.
  7. (7)
    As mentioned in Section 4, the pair $(\mathrm{proj}^{\leq n}(\Gamma),\mathrm{inj}^{\leq 1}(\Gamma))$ is a cotorsion pair and so $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}^{\leq 1}(\Gamma)$ is the class of all ${\rm{Ext}}$-injective objects in $\mathrm{proj}^{\leq n}(\Gamma)$, i.e. $\mathcal{I}(\mathrm{proj}^{\leq n}(\Gamma))=(\bigoplus_{i=2}^{n+1}I(i))\oplus S(2).$
  8. (8)
    By Corollary 2.12 $(ii)$, the Auslander-Reiten sequence
    $0\longrightarrow P(1)\longrightarrow S(1)\oplus P(n+1)\longrightarrow I(1)\longrightarrow 0$
    in ${\rm{{mod}}}(\Gamma)$ induces the following Auslander-Reiten sequence in $\mathrm{proj}^{\leq n}(\Gamma)$:
    $0\longrightarrow P(1)\longrightarrow P(n+1)\longrightarrow S(n+1)\longrightarrow 0$
  9. (9)
    Since ${\rm{Ext}}^{1}_{\Gamma}(S(1),S(1))=0$, $S(1)$ is both ${\rm{Ext}}$-projective and ${\rm{Ext}}$-injective in ${}^{\perp_{0}}\Gamma$. Hence, $\mathcal{P}(^{\perp_{0}}\Gamma)=S(1)=\mathcal{I}(^{\perp_{0}}\Gamma)$.
Final Remark 5.8.
Let $R$ be a commutative artinian ring and $n$ be a positive integer. In 2017, Iyama and Jasso extended higher Auslander correspondence from Artin $R$-algebras of finite representation type to dualizing $R$-varieties. By their result, i.e. higher Auslander correspondence for dualizing $R$-varieties, we know that an Artin $R$-algebra $\Gamma$ is an $n$-Auslander algebra if and only if the category of all finitely generated projective $\Gamma$-modules is an $n$-abelian category [IJ17, Theorem 1.2]. But by some results due to Beligiannis [Be15], even for $n$-minimal Auslander-Gorenstein algebras, one can see that the category of all finitely generated Gorenstein projective $\Gamma$-modules is not an $n$-abelian category in general. Hence, based on our results, there is a difference between the abelianness of ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ and the $n$-abelianness of ${\mathrm{Gproj}}(\Gamma)$ in general.
Acknowledgements: The authors were supported by the National Natural Science Foundation of China (No. 12071137), by Key Laboratory of MEA (Ministry of Education), by Shanghai Key Laboratory of PMMP (No. 22DZ2229014), and by Fundamental Research Funds for the Central Universities.

### v2

@@
 Let $n$ be a positive integer.
-In this paper, by using torsion theoretic methods, we show that $n$-Auslander algebras can be characterized by the abelianness of the category of modules with projective dimension less than $n$ and a certain additional property, extending the classical Auslander-Tachikawa theorem.
+In this paper, by using torsion theoretic methods, we show that $n$-Auslander algebras can be characterized by the abelianness of the category of modules with projective dimension less than $n$ and two additional properties, extending the classical Auslander-Tachikawa theorem.
 By Auslander-Iyama correspondence a categorical characterization of the class of Artin algebras having $n$-cluster tilting modules is obtained.
 Since higher Auslander algebras are a special case of higher Auslander-Gorenstein algebras, the results are given in the general setting as extending previous results of Kong.
-Moreover, as an application of some results, we give categorical descriptions for the semisimplicity and selfinjectivity of an Artin algebra.
 Higher Auslander-Gorenstein Algebras are also studied from the viewpoint of cotorsion pairs and, as application, we show that they satisfy in two nice equivalences.
@@
 16G10; 16E65
-Introduction
+1. Introduction
 An important problem in the representation theory of algebras is to study algebras of finite representation type; i.e.
 Artin algebras that have only a finite number of isomorphism classes of finitely generated indecomposable modules.
-In 1971, Auslander [Au71] proved a remarkable classical result, called Auslander correspondence, about algebras of finite representation type which introduced a completely new insight to the representation theory of Artin algebras.
+In 1971, Auslander [3] proved a remarkable classical result, called Auslander correspondence, about algebras of finite representation type which introduced a completely new insight to the representation theory of Artin algebras.
 He showed that there is a bijective correspondence between the set of Morita equivalence classes of Artin algebras $\Lambda$ of finite representation type and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an Auslander algebra; i.e.
-i.e.
 Artin algebras $\Gamma$ whose global dimensions (denoted by ${\rm{gldim}}(\Gamma)$) are at most two and whose dominant dimensions (denoted by ${\rm{domdim}}(\Gamma)$) are at least two.
-See [Au71, Chapter III, Section 4, Page 52] or [ARS, Theorem VI.5.7].
+See [3, Page 52] or [5, Theorem VI.5.7].
 For an Artin algebra $\Gamma$, let ${\rm{Mod}}(\Gamma)$ be the category of all left $\Gamma$-modules and ${\rm{{mod}}}(\Gamma)$ be the category of all finitely generated left $\Gamma$-modules.
-If $\Gamma$ is an Auslander algebra, then the full subcategory $\mathrm{proj}(\Gamma)$ of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated projective $\Gamma$-modules is equivalent to a module category and so is itself an abelian category [Au71, §III, Section 4, Page 52].
+If $\Gamma$ is an Auslander algebra, then the full subcategory $\mathrm{proj}(\Gamma)$ of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated projective $\Gamma$-modules is equivalent to a module category and so is itself an abelian category [3, Page 52].
 This gave Tachikawa the motivation to study rings $\Gamma$ such that the full subcategory ${\mathrm{Proj}}(\Gamma)$ of ${\rm{Mod}}(\Gamma)$ consisting of all projective $\Gamma$-modules is abelian.
-He obtained some characterizations of such rings and proved that these algebras are precisely Auslander algebras [Ta74, Theorem 1].
+He obtained some characterizations of such rings and proved that these algebras are precisely Auslander algebras [44, Theorem 1].
 For Artin algebras, by combining Auslander’s result and Tachikawa’s result, we have the following characterization of Auslander algebras.
@@
     the full additive subcategory ${\mathrm{Proj}}(\Gamma)$ of ${\rm{Mod}}(\Gamma)$ is an abelian category.
-Hence, for Artin algebras, by Auslander correspondence and Auslander-Tachikawa theorem, there are bijections between the following four sets:
-  • (i)
-    Morita equivalence classes of Artin algebras $\Lambda$ such that $\Lambda$ is of finite representation type;
-  • (ii)
-    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an Auslander algebra;
-  • (iii)
-    Morita equivalence classes of Artin algebras $\Delta$ such that $\mathrm{proj}(\Delta)$ is an abelian category;
-  • (iv)
-    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{Proj}}(\Sigma)$ is an abelian category.
-Moreover, an Artin algebra $\Lambda$ is of finite representation type if and only if ${\rm{{mod}}}(\Lambda)\simeq\mathrm{proj}(\Gamma)$, where $\Gamma$ is an Auslander algebra; see also [Iy04, Theorem 3.3].
-In 2007, Iyama in a series of important papers [Iy07a, Iy07b] gave a generalization of Auslander correspondence, called higher Auslander correspondence or Auslander-Iyama correspondence.
-He showed that for every positive integer $n$, there is a bijective correspondence between the set of equivalence classes of $n$-cluster tilting subcategories with additive generators (Definition 5.5) of Artin algebras and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that ${\rm{gldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$, called $n$-Auslander algebras [Iy07b, Theorem 0.2].
-A natural question is whether there exists a characterization for higher Auslander algebras analogous to Auslander-Tachikawa Theorem.
-The first goal of this paper is to answer this question and we have the following characterization for higher Auslander algebras; see Theorem 5.1.
+In 2007, Iyama in a series of important papers [26, 27] gave a generalization of Auslander correspondence, called higher Auslander correspondence or Auslander-Iyama correspondence.
+He showed that for every positive integer $n$, there is a bijective correspondence between the set of equivalence classes of $n$-cluster tilting subcategories with additive generators (Definition 6.4) of Artin algebras and the set of Morita equivalence classes of Artin algebras $\Gamma$ such that ${\rm{gldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$, called $n$-Auslander algebras [27, Theorem 0.2].
+A natural question is whether there exists a characterization for higher Auslander algebras analogous to the Auslander-Tachikawa Theorem.
+The first goal of this paper is to answer this question and we have the following characterizations for higher Auslander algebras; see Theorem 6.1.
 Theorem 2 (Higher Auslander-Tachikawa Theorem).
@@
     ${\mathrm{Proj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$, and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n}}\Gamma}$.
-Here, for an Artin algebra $\Gamma$, we denote by $\Gamma^{{\rm{op}}}$ the opposite algebra of $\Gamma$, by ${\mathrm{Inj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all injective $\Gamma$-modules, by $\mathrm{inj}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated injective $\Gamma$-modules; for every non-negative integer $m$, by ${\mathrm{Proj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of projective dimension at most $m$, and by $\mathrm{proj}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of projective dimension at most $m$.
+Here, for an Artin algebra $\Gamma$, we denote by $\Gamma^{{\rm{op}}}$ the opposite algebra of $\Gamma$, by ${\mathrm{Inj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all injective $\Gamma$-modules, by $\mathrm{inj}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated injective $\Gamma$-modules; for every non-negative integer $m$, by ${\mathrm{Proj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all $\Gamma$-modules of projective dimension at most $m$, and by $\mathrm{proj}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules of projective dimension at most $m$.
 For $m\geq 1$, we write
-[…]
+$^{\perp_{m}}\Gamma:={}^{\perp_{[1,m]}}\Gamma=\{M\in{\rm{{mod}}}(\Gamma)\ |\ {\rm{Ext}}^{i}_{\Gamma}(M,\Gamma)=0,\ \forall\ 1\leq i\leq m\},$
 and
-[…]
+$^{\underline{\perp}_{m}}\Gamma:={}^{\underline{\perp}_{[1,m]}}\Gamma=\{M\in{\rm{Mod}}(\Gamma)\ |\ {\rm{Ext}}^{i}_{\Gamma}(M,\Gamma)=0,\ \forall\ 1\leq i\leq m\};$
 we also write
@@
 Note that only the abelianness property of the category of modules with projective dimension less than $n$ can not characterize $n$-Auslander algebras.
-For $n\geq 2$, the other conditions are necessary, although they hold automatically for $n=1$; see Remarks 5.2 and 5.3.
-In 1993, Auslander and Solberg [AS93] established a Gorenstein analogue of Auslander correspondence by using the technique of relative homology developed in [AS93a, AS93b, AS93c].
+For $n\geq 2$, the other conditions are necessary; see Remarks 6.2.
+In 1993, Auslander and Solberg [10] established a Gorenstein analogue of Auslander correspondence by using the technique of relative homology developed in [7, 8, 9].
 They introduced $\tau$-selfinjective algebras, where $\tau$ is the Auslander-Reiten translation, as a replacement of representation-finite algebras.
 The corresponding analogue of Auslander algebras are characterized as Artin algebras whose injective dimensions are at most $2$ and whose dominant dimensions are at least $2$, that is, Artin algebras $\Gamma$ such that ${\rm{id}}({}_{\Gamma}\Gamma)\leq 2\leq{\rm{domdim}}(\Gamma)$.
-Kong named this class of algebras as quasi-Auslander algebras [Ko14].
-In his work [Ko14] Kong also proved an extension of Auslander-Tachikawa’s theorem by characterizing quasi-Auslander algebras as Artin algebras over which the category of finitely generated Gorenstein projective modules is an abelian category.
-Hence, for Artin algebras, by Auslander-Solberg correspondence ([AS93, Proposition 5.11]) and Kong’s theorem [Ko14, Theorem 2.1], there are bijections between the following four sets:
-  • (i)
-    Equivalence classes of suitable modules of Artin algebras;
-  • (ii)
-    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is a quasi-Auslander algebra;
-  • (iii)
-    Morita equivalence classes of Artin algebras $\Delta$ such that ${\mathrm{Gproj}}(\Delta)$ is an abelian category;
-  • (iv)
-    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{GProj}}(\Sigma)$ is an abelian category.
-To obtain a higher analogue of Auslander-Solberg correspondence, Iyama and Solberg [IS18] introduced pre-cluster tilting subcategories in module categories of Artin algebras which are called $\tau_{n}$-selfinjective algebras as a higher version of $\tau$-selfjinjective algebras.
+Kong named this class of algebras as quasi-Auslander algebras [33].
+In his work [33] Kong also proved an extension of Auslander-Tachikawa theorem by characterizing quasi-Auslander algebras as Artin algebras over which the category of finitely generated Gorenstein projective modules is an abelian category.
+To obtain a higher analogue of Auslander-Solberg correspondence, Iyama and Solberg [31] introduced pre-cluster tilting subcategories in module categories of Artin algebras which are called $\tau_{n}$-selfinjective algebras as a higher version of $\tau$-selfjinjective algebras.
 They also gave the notion of $n$-minimal Auslander-Gorenstein algebras as a higher version of quasi-Auslander algebras, say Artin algebras $\Gamma$ with ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq{\rm{domdim}}(\Gamma)$.
 They showed that there exists a bijective correspondence between Morita equivalence classes of $n$-minimal Auslander-Gorenstein algebras and equivalence classes of finite $n$-precluster tilting subcategories of Artin algebras.
-In view of the above Iyama-Solberg correspondence, we could establish a Gorenstein version of Theorem 2 which are also higher versions of Kong’s result [Ko14, Theorem 2.1].
+In view of the above Iyama-Solberg correspondence, we could establish a Gorenstein version of Theorem 2 which are also higher versions of Kong’s result [33, Theorem 2.1].
 Theorem 3 (Theorem 3.6).
@@
     ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$ and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$.
-Here for every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$, by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
-Hence, for Artin algebras, by Iyama-Solberg correspondence [IS18, Theorem 4.5] and our results (Theorem 2.6), there are bijections between the following sets:
-  • (I)
-    Equivalence classes of finite $n$-precluster tilting subcategories of Artin algebras;
-  • (II)
-    Morita equivalence classes of Artin algebras $\Gamma$ such that $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra;
-  • (III)
-    Morita equivalence classes of Artin algebras $\Delta$ such that ${\mathrm{Gproj}}^{\leq n-1}(\Delta)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Delta)\cap\mathrm{inj}(\Delta)\subseteq\mathrm{proj}(\Delta)$, and ${{}^{\perp_{0}}\Delta}\subseteq{{}^{\perp_{n}}\Delta}$;
-  • (IV)
-    Morita equivalence classes of Artin algebras $\Sigma$ such that ${\mathrm{GProj}}^{\leq n-1}(\Sigma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Sigma)\cap{\mathrm{Inj}}(\Sigma)\subseteq{\mathrm{Proj}}(\Sigma)$, and ${{}^{\underline{\perp}_{0}}\Sigma}\subseteq{{}^{\underline{\perp}_{n}}\Sigma}$.
-While the results of Section 2 mainly make use of torsion theoretic methods, in Section 4, we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS].
-Especially, by Corollary 4.4, if $n$ is a positive integer and $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, then we have the following equivalences:
+Here for every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$, by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
+While the results of Section 3 mainly make use of torsion theoretic methods, in Section 5, we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [11].
+Especially, by Corollary 5.4, if $n$ is a positive integer and $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, then we have that
 ${{}^{\perp_{0}}\Gamma}\simeq\frac{\mathrm{inj}^{\leq 1}(\Gamma)}{\mathrm{inj}^{\leq 1}(\Gamma)\bigcap\mathrm{proj}^{\leq n}(\Gamma)}\simeq\frac{\mathrm{inj}^{\leq 1}(\Gamma)}{\mathrm{inj}(\Gamma)}.$
@@
 Let $\Gamma$ be an Artin algebra defined over $R$.
-We denote by $\mathbf{D}$ the canonical duality ${\rm{Hom}}_{R}(-,\mathbf{E}$ $(R/J_{R})):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$, where $\mathbf{E}(R/J_{R})$ is the injective envelop of $R/J_{R}$ as $R$-module.
+We denote by $\mathbf{D}$ the canonical duality ${\rm{Hom}}_{R}(-,\mathbf{E}$ $(R/J_{R})):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$, where $\mathbf{E}(R/J_{R})$ is the injective envelope of $R/J_{R}$ as an $R$-module.
 For a left $\Gamma$-module $X$, $\mathbf{E}(X)$ denotes the injective envelope of $X$; ${\rm{pd}}({}_{\Gamma}X)$ (resp.
@@
 For a module $M\in{\rm{{mod}}}(\Gamma)$ and $i\geq 0$, write $\mathrm{sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{add}(M))$, $\mathrm{fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{add}(M))$, $\mathrm{Sub}^{i}(M):=\mathrm{sub}^{i}(\mathrm{Add}(M))$, $\mathrm{Fac}_{i}(M):=\mathrm{fac}_{i}(\mathrm{Add}(M))$.
-In particular, $\mathrm{sub}(M):=\mathrm{sub}^{1}(M)$, $\mathrm{fac}(M)$ $:=\mathrm{fac}_{1}(M)$, $\mathrm{Sub}(M):=\mathrm{Sub}^{1}(M)$, $\mathrm{Fac}(M):=\mathrm{Fac}_{1}(M)$.
+In particular, $\mathrm{sub}(M):=\mathrm{sub}^{1}(M)$, $\mathrm{fac}(M)$ $\mathrm{fac}_{1}(M)$, $\mathrm{Sub}(M):=\mathrm{Sub}^{1}(M)$, $\mathrm{Fac}(M):=\mathrm{Fac}_{1}(M)$.
 2.2. Dominant Dimension
@@
 Note that the dominant dimension of a right module and the right dominant dimension of a ring are defined similarly.
-It is well-known that ${\rm{domdim}}({{}_{\Gamma}\Gamma})={\rm{domdim}}(\Gamma_{\Gamma})$ ; see [Mu68, Theorem 4].
+It is well-known that ${\rm{domdim}}({{}_{\Gamma}\Gamma})={\rm{domdim}}(\Gamma_{\Gamma})$ ; see [39, Theorem 4].
 So for the rest of the paper, we will denote both left and right dominant dimension of $\Gamma$, by ${\rm{domdim}}(\Gamma)$ and call it the dominant dimension of $\Gamma$.
@@
 So $M\in{{}^{\perp_{n}}\Gamma}$.
-The inclusion ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$ can be proves exactly in the same way. ∎
+The inclusion ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$ can be proved exactly in the same way. ∎
 2.3. Torsion Pairs
-Let $\mathcal{A}$ be an abelian category.
-If $X,Y\in\mathcal{A}$, then we denote the set of morphisms from $X$ to $Y$ in $\mathcal{A}$ by $\mathcal{A}(X,Y)$ or ${\rm{Hom}}_{\mathcal{A}}(X,Y)$.
-All subcategories considered are assumed to be full and closed under isomorphisms and finite direct sums.
-A torsion pair (or torsion theory) in the abelian category $\mathcal{A}$ is a pair $(\mathcal{X},\mathcal{Y})$ of subcategories such that $\mathcal{A}(X,Y)=0$ for all $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$, and these two classes are maximal for this property [Di66, Sections 1 and 2], that is,
+Let $\Gamma$ be an Artin algebra and $\mathcal{A}:={\rm{Mod}}(\Gamma)$.
+If $X,Y\in\mathcal{A}$, then we denote the set of morphisms from $X$ to $Y$ in $\mathcal{A}$ by $\mathcal{A}(X,Y)$.
+A torsion pair in $\mathcal{A}$ is a pair $(\mathcal{X},\mathcal{Y})$ of full subcategories of $\mathcal{A}$ such that $\mathcal{A}(X,Y)=0$ for all $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$, and these two classes are maximal for this property [16, Sections 1 and 2], that is,
 $\mathcal{X}={{}^{\perp_{0}}\mathcal{Y}}:=\{X\in\mathcal{A}\ |\ \mathcal{A}(X,Y)=0,\forall\ Y\in\mathcal{Y}\},$
@@
 the torsion-free class) of some torsion pair if and only if it is closed under quotients, direct sums and extensions (resp.
-subobjects, direct products and extensions, respectively) [Di66, Theorem 2.3].
-If $(\mathcal{X},\mathcal{Y})$ is a torsion pair in $\mathcal{A}$, to each object $A$ in $\mathcal{A}$ is associated a unique short exact sequence (called the canonical sequence for $A$)
-$0\longrightarrow xA\longrightarrow A\longrightarrow yA\longrightarrow 0$
-such that $xA\in\mathcal{X}$ and $yA\in\mathcal{Y}$.
-In fact, $xA$ is the maximal subobject of $A$ contained in $\mathcal{X}$ [Di66, Proposition 2.4].
-Also, the canonical sequence is functorial, i.e.
-it defines the functors $x:\mathcal{A}\longrightarrow\mathcal{X}$ given by $A\mapsto xA$ and $y:\mathcal{A}\longrightarrow\mathcal{Y}$ given by $A\mapsto yA$, see [BR07, Section 1.1].
-Note that the composition of the functor $x:\mathcal{A}\longrightarrow\mathcal{X}$ by the inclusion functor $\mathcal{X}\longrightarrow\mathcal{A}$ is an idempotent radical, see [St75, Proposition VI.2.3].
+subobjects, direct products and extensions) [16, Theorem 2.3].
 A torsion pair $(\mathcal{X},\mathcal{Y})$ in $\mathcal{A}$ is called hereditary if $\mathcal{X}$ is also hereditary, i.e.
-$\mathcal{X}$ is closed under subobjects (which is equivalent to $\mathcal{Y}$ being closed under taking injective envelopes [Di66, Theorem 2.9]).
-Let now $\Gamma$ be an Artin algebra.
-We would like to mention that in this case, a subcategory of ${\rm{{mod}}}(\Gamma)$ is the torsion class (resp.
+$\mathcal{X}$ is closed under subobjects (which is equivalent to $\mathcal{Y}$ being closed under taking injective envelopes [16, Theorem 2.9]).
+For more details about torsion theories in ${\rm{Mod}}(\Gamma)$; see [43, Chapter VI].
+We would like to mention that a subcategory of ${\rm{{mod}}}(\Gamma)$ is the torsion class (resp.
 the torsion-free class) of some torsion pair if and only if it is closed under quotients and extensions (resp.
-submodules and extensions, respectively).
-For more details about torsion theories in ${\rm{{mod}}}(\Gamma)$; see [AIR, Subsection 1.1] or [Iy04, Subsection 1.2].
-1.4. Functorially Finite Subcategories
-Let $\Gamma$ be an Artin algebra.
-A subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Gamma)$ is said to be covariantly finite in ${\rm{{mod}}}(\Gamma)$, if all representable covariant functors on ${\rm{{mod}}}(\Gamma)$ are finitely generated when considered as functors on $\mathcal{X}$ by restriction.
-In other words, for each $M\in{\rm{{mod}}}(\Gamma)$, there exist $X\in\mathcal{X}$ and a morphism $\varphi:M\longrightarrow X$ so that the sequence
-${\rm{Hom}}_{\Gamma}(X,-)\mid_{\mathcal{X}}\longrightarrow{\rm{Hom}}_{\Gamma}(M,-)\mid_{\mathcal{X}}\longrightarrow 0$
-is exact. Here, the notation $\mid_{\mathcal{X}}$ means restriction to $\mathcal{X}$.
-Dually, $\mathcal{X}$ is said to be contravariantly finite in ${\rm{{mod}}}(\Gamma)$ if the representable contravariant functors on ${\rm{{mod}}}(\Gamma)$ stay finitely generated when restricted to $\mathcal{X}$.
-A subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Gamma)$ is said to be functorially finite if it is both contravariantly and covariantly finite in ${\rm{{mod}}}(\Gamma)$.
-Note that if the pair $(\mathcal{X},\mathcal{Y})$ is a torsion pair in ${\rm{{mod}}}(\Gamma)$, then an easy investigation shows that $\mathcal{X}$ is a contravariantly finite and $\mathcal{Y}$ is covariantly finite in ${\rm{{mod}}}(\Gamma)$ [Sm84, Page 519].
+submodules and extensions).
+For more details about torsion theories in ${\rm{{mod}}}(\Gamma)$; see [1, Subsection 1.1] or [25, Subsection 1.2].
 2.4. (Co)tiliting Modules
-We recall some definitions and notations about (co)tilting modules and their relations to torsion theory; see [Ho82, As84, Sm84, AIR].
+We recall some definitions and notations about (co)tilting modules; see [22, 41].
 Definition 2.2 (Tilting and Cotilting Modules).
-Let $\Gamma$ be an Artin algebra. A $\Gamma$-module $T$ is said to be a tilting module, if the following three properties are satisfied.
+Let $\Gamma$ be an Artin algebra. A module $T\in{\rm{{mod}}}(\Gamma)$ is said to be a tilting module, if the following three properties are satisfied.
   • (i)
@@
   • (iii)
-    There is an exact sequence $0\longrightarrow\Gamma\longrightarrow T^{\prime}\longrightarrow T^{\prime\prime}\longrightarrow 0$ with $T^{\prime}$ and $T^{\prime\prime}$ direct summands of direct sums of copies of $T$.
-Dually, the $\Gamma$-module $T$ will be called a cotilting module if $\mathbf{D}(T)$ is a tilting $\Gamma^{\rm op}$-module.
+    There is an exact sequence $0\longrightarrow\Gamma\longrightarrow T^{\prime}\longrightarrow T^{\prime\prime}\longrightarrow 0$ with $T^{\prime}$ and $T^{\prime\prime}$ in ${\rm{add}}(T)$.
+Dually, a module $T\in{\rm{{mod}}}(\Gamma)$ will be called a cotilting module if $\mathbf{D}(T)$ is a tilting $\Gamma^{\rm op}$-module.
 Definition 2.3 (Ext-projectives and Ext-injectives).
@@
 $\mathrm{Ann}_{\Gamma}(\mathcal{X}):=\{\gamma\in\Gamma|\gamma X=0,\ \forall\ X\in\mathcal{X}\}$.
-For the convenience of the reader, we recall the following relevant result on torsion pairs that we will use of it in next sections.
-Theorem 1.4 ([Sm84, Theorem] and [AIR, Proposition 1.1]).
-Let $\Gamma$ be an Artin algebra and the pair $(\mathcal{X},\mathcal{Y})$ be a torsion pair in ${\rm{{mod}}}(\Gamma)$. Then the following conditions are equivalent.
-  • $(a)$
-    $\mathcal{X}$ is functorially finite;
-  • $(b)$
-    $\mathcal{Y}$ is functorially finite;
-  • $(c)$
-    $\mathcal{X}=\mathrm{fac}(X)$ for some $X\in{\rm{{mod}}}(\Gamma)$;
-  • $(d)$
-    $\mathcal{Y}=\mathrm{sub}(Y)$ for some $Y\in{\rm{{mod}}}(\Gamma)$;
-  • $(e)$
-    $\mathcal{P}(\mathcal{X})$ is a tilting $(\Gamma/\mathrm{Ann}_{\Gamma}(\mathcal{X}))$-module;
-  • $(f)$
-    $\mathcal{I}(\mathcal{Y})$ is a cotilting $(\Gamma/\mathrm{Ann}_{\Gamma}(\mathcal{Y}))$-module;
-  • $(g)$
-    $\mathcal{X}=\mathrm{fac}(\mathcal{P}(\mathcal{X}))$;
-  • $(h)$
-    $\mathcal{Y}=\mathrm{sub}(\mathcal{I}(\mathcal{Y}))$.
-1.6. Auslander-Reiten Theory
-In this subsection we recall several results in Auslander-Reiten theory.
-Let $\Gamma$ be an Artin algebra. The functor
-()*$(-)^{*}:={\rm{Hom}}_{\Gamma}(-,\Gamma):{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{{\rm{op}}})$
-defined by
-$M^{*}:={\rm{Hom}}_{\Gamma}(M,\Gamma)$
-induces a duality ()*$(-)^{*}:{\rm{add}}(\Gamma)\longrightarrow{\rm{add}}(\Gamma^{{\rm{op}}})$, see e.g. [ARS, Proposition II.4.3].
-For a $\Gamma$-module $X$, take a projective presentation $P_{1}\stackrel{{\scriptstyle g}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle f}}{{\longrightarrow}}X\longrightarrow 0.$ Define $\Gamma^{{\rm{op}}}$-module $\mathrm{Tr}X$ by an exact sequence $P_{0}^{*}\stackrel{{\scriptstyle g^{*}}}{{\longrightarrow}}P_{1}^{*}\longrightarrow\mathrm{Tr}X\longrightarrow 0.$ In fact $\mathrm{Tr}X$ is defined as the cokernel of $g^{*}$.
-On the other hand, the canonical duality $\mathbf{D}:{\rm{{mod}}}(\Gamma)\longrightarrow{\rm{{mod}}}(\Gamma^{\rm{\rm{op}}})$ induces a duality $\mathbf{D}:\underline{{\rm{{mod}}}}(\Gamma)\longrightarrow\overline{{\rm{{mod}}}}(\Gamma^{\rm{\rm{op}}})$ and we have mutually quasi-inverse equivalences
-$\tau:=\mathbf{D}\mathrm{Tr}:\underline{{\rm{{mod}}}}(\Gamma)\stackrel{{\scriptstyle\mathrm{Tr}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma^{{\rm{op}}})\stackrel{{\scriptstyle\mathbf{D}}}{{\longrightarrow}}\overline{{\rm{{mod}}}}(\Gamma)$
-$\tau^{-}:=\mathrm{Tr}\mathbf{D}:\overline{{\rm{{mod}}}}(\Gamma)\stackrel{{\scriptstyle\mathbf{D}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma^{{\rm{op}}})\stackrel{{\scriptstyle\mathrm{Tr}}}{{\longrightarrow}}\underline{{\rm{{mod}}}}(\Gamma)$
-called the Auslander-Reiten translations.
-Especially, $\tau$ gives a bijection from isoclasses of indecomposable non-projective $\Gamma$-modules to isoclasses of indecomposable non-injective $\Gamma$-modules, see e.g.
-[ARS, Propositions IV.1.9 and IV.1.10].
 2.5. Cotorsion Pairs
@@
 Let now $\mathcal{X}$ and $\mathcal{Y}$ be full subcategories of the abelian category $\mathcal{A}$.
-Recall that a pair $(\mathcal{X},\mathcal{Y})$ is called a cotorsion pair (or cotorsion theory) if $\mathcal{X}={}^{\perp_{1}}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp_{1}}$, see for instance, [EJ11, Definition 7.1.2].
+Recall that a pair $(\mathcal{X},\mathcal{Y})$ is called a cotorsion pair (or cotorsion theory) if $\mathcal{X}={}^{\perp_{1}}\mathcal{Y}$ and $\mathcal{Y}=\mathcal{X}^{\perp_{1}}$, see for instance, [18, Definition 7.1.2].
 In this case, the class $\mathcal{X}$ is called a cotorsion class and the class $\mathcal{Y}$ is called a cotorsion-free class.
@@
 $\mathcal{Y}$ is a special preenveloping class.
-Salce’s lemma (which goes back to [Sa79]) asserts that $(i)$ and $(ii)$ are equivalent in the case where $\mathcal{A}$ is a module category, see, e.g.
-[EJ11, Proposition 7.1.7].
-The proof of this lemma shows that if the abelian category $\mathcal{A}$ has enough injectives, then $(i)\Rightarrow(ii)$ and if $\mathcal{A}$ has enough projectives, then $(ii)\Rightarrow(i)$.
-In contrast to the case with torsion pairs, the constructions $x,\tilde{x}$ and $y,\tilde{y}$ are not functorial.
-But they define functors from $\mathcal{A}$ to $\frac{\mathcal{X}}{\mathcal{X}\cap\mathcal{Y}}$ and $\frac{\mathcal{Y}}{\mathcal{X}\cap\mathcal{Y}}$, respectively [BBOS, Lemma 2.11].
-Note that there are slightly different definitions of cotorsion pairs which have appeared in the literature, see [BR07, KS03].
-By [BR07, Remark V.3.2], there is no difference between the notion of a complete cotorsion pair that we recalled it in the above definition and the notion of a cotorsion pair in the sense of Beligiannis-Reiten [BR07, Definition V.3.1].
-Also, cotorsion pairs in the sense of Krause-Solberg [KS03, Definition 2.1] are hereditary complete cotorsion pairs.
 2.6. Projectivization and Injectivization
@@
 $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$
-restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$; see, for instance, [ARS, Proposition II.2.1 (c)].
+restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$; see, for instance, [5, Proposition II.2.1 (c)].
 Lemma 2.4.
@@
 Proof.
-The first statement is well known; see, for instance, [ARS, Proposition II.2.5].
+The first statement is well known; see, for instance, [5, Proposition II.2.5].
 For the second statement, since $P$ is finitely generated, $\mathrm{Hom}_{\Lambda}(P,-)$ commutes with direct sums, thus it establishes an equivalence between $\mathrm{Add}(P)$ and ${\mathrm{Proj}}(\Gamma)$.
-Then the proof proceeds as that of [ARS, Proposition II.2.5].
+Then the proof proceeds as that of [5, Proposition II.2.5].
 ∎
@@
 Proof.
-(a) This is the first lemma of [Au71, Section 4].
+(a) This is the first lemma of [3, Chapter III, Section 4, Page 48].
 (b) It is easy to see that $I$ is a finitely generated projective module, as $I\cong\nu^{-1}\mathcal{Q}$, where $\nu^{-1}$ is the inverse Nakayama functor.
-The two functors $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q})$ and $\mathrm{Hom}_{\Gamma}(I,-)$ are exact, in particular, right exact, so by Watts’ theorem (see, for example, [Ro09, Corollary 5.47]), they are isomorphic to a tensor functor.
+The two functors $\mathbf{D}\mathrm{Hom}_{\Gamma}(-,\mathcal{Q})$ and $\mathrm{Hom}_{\Gamma}(I,-)$ are exact, in particular, right exact, so by Watts’ theorem (see, for example, [40, Corollary 5.47]), they are isomorphic to a tensor functor.
 It suffices to check that $\mathbf{D}\mathrm{Hom}_{\Gamma}(\Gamma,\mathcal{Q})\cong\mathbf{D}\mathcal{Q}\cong\mathrm{Hom}_{\Gamma}(I,\Gamma)$.
@@
 commutes with direct sums, so the equivalence $\mathrm{add}(\mathcal{Q})\simeq\mathrm{inj}(\Lambda)$ extends to $\mathrm{Add}(\mathcal{Q})\simeq{\mathrm{Inj}}(\Lambda)$.
-As in the proof of the first lemma of [Au71, Section 4], the latter equivalence extends to another equivalence
+As in the proof of the first lemma of [3, Chapter III, Section 4], the latter equivalence extends to another equivalence
 $\mathrm{Hom}_{\Gamma}(I,-)\simeq\mathbf{D}\mathcal{Q}\otimes_{\Gamma}-:\mathrm{Sub}^{2}(\mathcal{Q})\to{\rm{Mod}}(\Lambda).$
 (d) follows from (a) and (c). ∎
-1.9. Injective Resolutions
-For the convenience of the reader, this subsection is devoted to state some important results about injective resolutions which we’ll make use of them in several proofs.
-Recall that an Iwanaga-Gorenstein algebra is a two-sided Noetherian ring such that its injective dimensions both as a left and as a right $\Gamma$-module are finite, which are equal by Zaks [Za69, Lemma A] ( see also [EJ11, Proposition 9.1.8]).
-Theorem 1.7 (Iwanaga [Iw79, Theorem 2]).
-Let $\Gamma$ be an Iwanaga-Gorenstein algebra with the minimal injective resolution
-$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow I^{1}\longrightarrow\cdots.$
-Then $\bigoplus_{i\geq 0}I^{i}$ is an injective cogenerator.
-Theorem 1.8 (Iwanaga [Iw80, Theorem 3 (1)]).
-Let $\Gamma$ be an Artinian Iwanaga-Gorenstein algebra with the minimal injective resolution
-$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow 0$
-Then the projective dimension of every direct summand $I$ of $I^{n}$ is $n$, i.e. ${\rm{pd}}({}_{\Gamma}I)=n$.
-Hence, for Gorenstein algebras of Gorenstein dimension $n+1$, we have the following proposition.
+2.7. Iwanaga-Gorenstein algebras
+Recall that an Artin algebra $\Gamma$ is called Iwanaga-Gorenstein if its injective dimensions both as a left and as a right $\Gamma$-module are finite, which are equal by Zaks [45, Lemma A] ( see also [18, Proposition 9.1.8]).
+For Iwanaga-Gorenstein algebras of injective dimension $n+1$, we have the following proposition.
 Proposition 2.6.
 Let $\Gamma$ be an Artinian Iwanaga-Gorenstein algebra of injective dimension at most $n+1$ with the minimal injective resolution
-$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\to 0$
+$0\longrightarrow{{}_{\Gamma}\Gamma}\longrightarrow I^{0}\longrightarrow\cdots\longrightarrow I^{n}\longrightarrow I^{n+1}\to 0.$
 Then $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq{\rm{add}}(\oplus_{i=0}^{n}I^{i}).$
@@
 If $M\in\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)$, then $M=\bigoplus_{l=1}^{r}M_{l}$, where $M_{l}$ is an injective indecomposable $\Gamma$-module and of projective dimension at most $n$.
-By Theorem 1.7, each $M_{l}$ is a direct summand of $I^{j}$ with $10\leq j\leq n+$ 1.
-But ${\rm{pd}}(M_{l})\leq n$, so by Theorem 1.8, $M_{l}$ is not a direct summand of $I^{n+1}$.
+By [23, Theorem 2], each $M_{l}$ is a direct summand of $I^{j}$ with $0\leq j\leq n+1$.
+But ${\rm{pd}}(M_{l})\leq n$, so by [24, Theorem 2], $M_{l}$ is not a direct summand of $I^{n+1}$.
 Thus $M_{l}$ is a direct summand of $I^{j}$ with $0\leq j\leq n$.
 This shows that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq{\rm{add}}(\oplus_{i=0}^{n}I^{i}).$
-Theorem 1.10 (Miyachi [Mi00, Coro llary 1.3]).
-Let
-$0\longrightarrow Y_{n}\longrightarrow\cdots\longrightarrow Y_{1}\longrightarrow Y_{0}\longrightarrow X\longrightarrow 0$
-be an exact sequence of $\Gamma$-modules. If
-$0\longrightarrow Y_{i}\longrightarrow I^{-i,0}\longrightarrow I^{-i,1}\longrightarrow\cdots$
-is an injective resolution of $Y_{i}$ ($0\leq i\leq n$), then $X$ has the injective resolution
-$0\longrightarrow X\longrightarrow T\longrightarrow\bigoplus_{r+s=1}I^{r,s}\longrightarrow\bigoplus_{r+s=2}I^{r,s}\longrightarrow\cdots\longrightarrow\bigoplus_{r+s=n}I^{r,s}\longrightarrow\cdots$
-where $T$ is a direct summand of $\bigoplus_{r+s=0}I^{r.s}$.
-1.10. Infinitely Generated Modules
-For later use, we record some facts about infinitely generated modules over an Artin algebra.
-Facts 1.11.
-Let $\Gamma$ is an Artin algebra.
-  • (a)
-    ([CE56, p. 17, Exercise 8]) Since $\Gamma$ is left Noetherian, arbitrary direct sums of injective (left) modules are injective modules, see also [Ch60, Proposition 4.1];
-  • (b)
-    ([KMT, Theorem 3.2]) Since $\Gamma$ is left Artinian, every injective module is a direct sum of indecomposable injective modules; see also Matlis [Ma58, Theorem 2.5];
-  • (c)
-    (Chase [Ch60, Theorem 3.3]) Since $\Gamma$ is left prefect and right coherent, direct products of projective modules are projective modules;
-  • (d)
-    (Müller [Mu70, Theorem 3]) Since $\Gamma$ is a semi-perfect ring, each projective module is a direct sum of finitely generated indecomposable projective modules.
 2.8. Gorenstein Projective Modules
@@
 ${\bf P}=\cdots\longrightarrow P_{1}\longrightarrow P_{0}\longrightarrow P^{0}\longrightarrow P^{1}\longrightarrow\cdots$
-such that $M\simeq\mathrm{Im}(P_{0}\longrightarrow P^{0})$ and such that ${\rm{Hom}}_{\Gamma}(-,Q)$ leaves the sequence ${\bf P}$ exact whenever $Q$ is a projective module.
-The class of all Gorenstein projective modules is denoted by ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{Gproj}}(\Gamma)$ denotes the subcategory of finitely generated Gorenstein projective modules; see [EJ11, Definition 10.2.1].
-Gorenstein injective modules are defined dually.
-Denote by ${\mathrm{GInj}}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all Gorenstein injective $\Gamma$-modules and by ${\mathrm{Ginj}}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated Gorenstein injective $\Gamma$-modules; see [EJ95, Definition 2.1] or [EJ11, Definition 10.1.1].
-The Gorenstein projective dimension of a $\Gamma$-module $M$ is at most $n\ (n\in\mathbb{N}_{0})$ if and only if $M$ has a Gorenstein projective resolution of length $n$, written as ${\rm{Gpd}}({}_{\Gamma}M)\leq n$.
-By [Ho04a, Theorem 2.10], ${\rm{Gpd}}({}_{\Gamma}M)\leq n$ if and only if there exists an exact sequence
-$0\longrightarrow P_{n}\longrightarrow\cdots\longrightarrow P_{1}\longrightarrow G\longrightarrow M\longrightarrow 0,$
-where $P_{n},\dots,P_{1}$ are projective modules and $G$ is a Gorenstein projective modules.
-Similarly, one can define the Gorenstein injective dimension ${\rm{Gid}}({}_{\Gamma}M)$ of $M$ [Ho04a, Definition 2.8].
-Note that if ${\rm{pd}}({}_{\Gamma}M)$ is finite, then ${\rm{Gpd}}({}_{\Gamma}M)={\rm{pd}}({}_{\Gamma}M)$ [Ho04a, Proposition 2.27].
-For every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$, consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$ and by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$, consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
+such that $M\simeq\mathrm{Im}(P_{0}\longrightarrow P^{0})$ and such that ${\rm{Hom}}_{\Gamma}(-,Q)$ leaves the sequence $\bf P$} exact whenever $Q$ is a projective module; see [18, Definition 10.2.1].
+The class of all Gorenstein projective modules is denoted by ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{Gproj}}(\Gamma)$ denotes the subcategory of finitely generated Gorenstein projective modules.
+The Gorenstein projective dimension of a $\Gamma$-module $M$ is at most $n\ (n\in\mathbb{N}_{0})$ if $M$ has a Gorenstein projective resolution of length $n$, written as ${\rm{Gpd}}({}_{\Gamma}M)\leq n$.
+For every non-negative integer $m$, we denote by ${\mathrm{GProj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{Mod}}(\Gamma)$ consisting of all $\Gamma$-modules of Gorenstein projective dimension at most $m$ and by ${\mathrm{Gproj}}^{\leq m}(\Gamma)$ the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules of Gorenstein projective dimension at most $m$.
 Like the classical case, the supremum of Gorenstein projective dimensions of all left $\Gamma$-modules is called left Gorenstein global dimension of $\Gamma$ and is denoted by ${\rm{Ggldim}}(\Gamma)$.
-By [BM10, Theorem 1.1], the left Gorenstein global dimension of $\Gamma$ is also the supremum of Gorenstein injective dimensions of all left $\Gamma$-modules.
-By [EJ11, Theorem 12.3.1], there is also no difference between left and right Gorenstein global dimensions of $\Gamma$ (as every Artin algebra is Noetherian).
-Also, the Gorenstein global dimension of $\Gamma$ is the supremum of Gorenstein projective dimensions of all finitely generated $\Gamma$-modules; see [Ch10, Theorem 3.2.8].
-Moreover, ${\rm{Ggldim}}(\Gamma)\leq n$ if and only if $\Gamma$ is an $n$-Iwanaga-Gorenstein algebra (i.e.
+Also, ${\rm{Ggldim}}(\Gamma)\leq n$ if and only if $\Gamma$ is an $n$-Iwanaga-Gorenstein algebra (i.e.
 ${\rm{id}}({}_{\Gamma}\Gamma)\leq n$ and ${\rm{id}}(\Gamma_{\Gamma})\leq n$).
-In fact, in this case, we have ${\rm{Ggldim}}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)={\rm{id}}(\Gamma_{\Gamma})$; see [EJ11, Corollary 12.3.2] or [Ch10, Corollary 3.2.6].
-For any left $\Gamma$-module $M$ with finite projective dimension, ${\rm{Gpd}}({}_{\Gamma}M)\leq n$ if and only if ${\rm{Ext}}^{i}_{\Gamma}(M,\Gamma$ )=0)=0) = 0 for all $i>n$; see [Ho04a, Theorem 2.20].
-Also, the Gorenstein injective dimension of $M$, ${\rm{Gid}}({}_{\Gamma}M)$, equals the usual injective dimension of $M$, i.e.
-if $M$ is a $\Gamma$-module with ${\rm{pd}}({}_{\Gamma}M)<\infty$, then ${\rm{Gid}}({}_{\Gamma}M)={\rm{id}}({}_{\Gamma}M)$; see [Ho04b, Theorem 2.1].
-This result has a dual counterpart and if $M$ is a $\Gamma$-module with ${\rm{id}}({}_{\Gamma}M)<\infty$, then ${\rm{Gpd}}({}_{\Gamma}M)={\rm{pd}}({}_{\Gamma}M)$; see [Ho04b, Theorem 2.2].
+In fact, in this case, we have ${\rm{Ggldim}}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)={\rm{id}}(\Gamma_{\Gamma})$; see [18, Corollary 12.3.2] or [14, Corollary 3.2.6].
 3. Higher Auslander-Gorenstein Algebras
-In this section, we will study higher Auslander-Gorenstein algebras introduced by Iyama and Solberg in [IS18] as a generalization of higher Auslander algebras and give a new characterization of them in terms of the abelianness property of the category of modules of Gorenstein projective dimension less than $n$.
+In this section, we will study higher Auslander-Gorenstein algebras introduced by Iyama and Solberg in [31] as a generalization of higher Auslander algebras and give a new characterization of them in terms of the abelianness property of the category of modules of Gorenstein projective dimension less than $n$.
 Of course, the results can be considered as a higher generalization of Auslander-Tachikawa theorem.
@@
 Let $n$ be a positive integer.
-Recall that an Artin algebra $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$ [IS18, Definition 1.1].
+Recall that an Artin algebra $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if ${\rm{id}}({}_{\Gamma}\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$ [31, Definition 1.1].
 $\Gamma$ is also called a higher Auslander-Gorenstein algebra, if $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra for a certain $n$.
-By [AR94, Corollary 5.5], these algebras are Iwanaga-Gorenstein algebras.
-It is also easily checked that any $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is either selfinjective or satisfies ${\rm{id}}({}_{\Gamma}\Gamma)=n+1={\rm{domdim}}(\Gamma)$; see [IS18, Proposition 4.1].
-Notice that, as mentioned in Subsection 1.11, for Gorenstein algebras, there is no difference between Gorenstein global dimention and injective dimension of $\Gamma$, i.e.
+By [4, Corollary 5.5], these algebras are Iwanaga-Gorenstein algebras.
+It is also easily checked that any $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is either selfinjective or satisfies ${\rm{id}}({}_{\Gamma}\Gamma)=n+1={\rm{domdim}}(\Gamma)$; see [31, Proposition 4.1].
+Notice that, as mentioned in Subsection 2.8, for Iwanaga-Gorenstein algebras, there is no difference between Gorenstein global dimension and injective dimension of $\Gamma$, i.e.
 $\mathrm{Ggldim}(\Gamma)={\rm{id}}({}_{\Gamma}\Gamma)$, and so $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if and only if ${\rm{Ggldim}}(\Gamma)\leq n+1\leq\rm{domdim}(\Gamma)$.
@@
 Let $X^{\prime}$ be the cokernel of the morphism $E^{0}\longrightarrow E^{1}$.
-Since ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}X^{\prime})\leq n+1$, so by [Ho04a, Theorem 2.20] $,{\rm{Gpd}}({}_{\Gamma}X)\leq n-1$, i.e.
+Since ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}X^{\prime})\leq n+1$, so by [20, Theorem 2.20] ${\rm{Gpd}}({}_{\Gamma}X)\leq n-1,$ i.e.
 $X\in{\mathrm{Gproj}}^{\leq n-1}(\Gamma)$.
@@
 $0\longrightarrow X\longrightarrow\mathbf{E}(X)\longrightarrow K\longrightarrow 0,$
-we have ${\rm{Gpd}}({}_{\Gamma}K)\leq 1$ and by [Ho04a, Theorem 2.10], there exists an exact sequence
+we have ${\rm{Gpd}}({}_{\Gamma}K)\leq 1$ and by [20, Theorem 2.10], there exists an exact sequence
 $0\longrightarrow P^{\prime}\longrightarrow G\longrightarrow K\longrightarrow 0$
 such that $G$ is a finitely generated Gorenstein projective module and $P^{\prime}$ is a finitely generated projective $\Gamma$-module.
-Since ${\rm{domdim}}(\Gamma)\geq n+1\geq 2$, both $\mathbf{E}(G)$ and $\mathbf{E}(\mathbf{E}(P^{\prime})/P^{\prime})$ are projective, so by Theorem 1.10, it follows that $\mathbf{E}(K)$ is a projective $\Gamma$-module as well.
+Since ${\rm{domdim}}(\Gamma)\geq n+1\geq 2$, both $\mathbf{E}(G)$ and $\mathbf{E}(\mathbf{E}(P^{\prime})/P^{\prime})$ are projective, so by [38, Corollary 1.3], it follows that $\mathbf{E}(K)$ is a projective $\Gamma$-module as well.
 Hence, $X$ fits into
@@
 with both $\mathbf{E}(X)$ and $\mathbf{E}(K)$ being projective-injective, hence $X$ falls into $\mathrm{sub}^{2}(\mathcal{Q})$.
-Now assume that $X$ is a $\Gamma$-module such that ${\rm{Gpd}}({}_{\Gamma}X)\leq n-1$. By [Ho04a, Theorem 2.10], there exists an exact sequence
+Now assume that $X$ is a $\Gamma$-module such that ${\rm{Gpd}}({}_{\Gamma}X)\leq n-1$. By [20, Theorem 2.10], there exists an exact sequence
 (3.2) $0\longrightarrow P_{n-1}\longrightarrow\cdots\longrightarrow P_{1}\longrightarrow G\longrightarrow X\longrightarrow 0$
-where $P_{i}$ is a finitely generated projective module for every $11\leq i\leq n-$ 1 and $G$ is a finitely generated Gorenstein projective module.
+where $P_{i}$ is a finitely generated projective module for every $1\leq i\leq n-1$ and $G$ is a finitely generated Gorenstein projective module.
 Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, ${\rm{Ggldim}}(\Gamma)\leq n+1\leq{\rm{domdim}}(\Gamma)$ and so every projective $\Gamma$-module $P_{i}$ has the minimal injective resolution
@@
 such that $J^{-i,k}\in\mathrm{proj}(\Gamma)\ \cap\ \mathrm{inj}(\Gamma)$, for every $k=0,\dots,n$.
-Applying Theorem 1.10 to the exact sequence (2.2) together with (2.1) and (2.3) then immediately gives the required result.
+Applying [38, Corollary 1.3] to the exact sequence (3.2) together with (3.1) and (3.3) then immediately gives the required result.
 We have shown that ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)=\mathrm{sub}^{2}(\mathcal{Q})$. ∎
@@
 Let $\Gamma$ be an Artin algebra.
-If a full additive category $\mathcal{C}$ of ${\rm{{mod}}}(\Gamma)$ has kernels and contains $\Gamma$, then the kernels in $\mathcal{C}$ coincident with the kernels in ${\rm{{mod}}}(\Gamma)$ and the inclusion functor from $\mathcal{C}$ to ${\rm{{mod}}}(\Gamma)$ is left exact.
+If a full additive category $\mathcal{C}$ of ${\rm{{mod}}}(\Gamma)$ has kernels and contains $\Gamma$, then the kernels in $\mathcal{C}$ coincides with the kernels in ${\rm{{mod}}}(\Gamma)$ and the inclusion functor from $\mathcal{C}$ to ${\rm{{mod}}}(\Gamma)$ is left exact.
 Proof.
@@
 We consider the following two complexes and the natural isomorphism between them.
-[…]
+${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\rm{Hom}}_{\Gamma}(\Gamma,K^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${{\rm{Hom}}_{\Gamma}(\Gamma,X)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${{\rm{Hom}}_{\Gamma}(\Gamma,Y)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{k^{\prime}}$${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Y}$
 Since $\Gamma\in\mathcal{C}$ the upper complex is an exact sequence of abelian groups, so the lower complex is also exact in the category of abelian groups.
@@
 Lemma 3.3.
-Let $\Gamma$ be an Artin algebra and $n\geq 1,0\leq k\leq n+1$. The following statements are equivalents:
+Let $\Gamma$ be an Artin algebra and $n\geq 1,0\leq k\leq n+1$. The following statements are equivalent:
   • $(a)$
@@
 Proof.
-$(a)\Rightarrow(b)$: We prove by induction on $k$.
+$(a)\Rightarrow(b)$.
+We prove by induction on $k$.
 For $k=0$, this is trivial.
@@
 Form the short exact sequence $0\to Y\to X\to Z\to 0$.
-As ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$, and combining with ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$ by [Ho04a, Theorem 2.20].
+As ${\rm{Ggldim}}(\Gamma)\leq n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$, and combining with ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$ by [20, Theorem 2.20].
 Suppose $(b)$ holds for $k$ ($1\leq k\leq n$). Then for $Y\in\mathrm{sub}^{k+1}({\mathrm{Gproj}}^{\leq n-k}(\Gamma))$, we have an exact sequence
@@
 $K^{1}\in\mathrm{sub}^{k}({\mathrm{Gproj}}^{\leq n-k}(\Gamma))\subseteq\mathrm{sub}^{k}({\mathrm{Gproj}}^{\leq n-k+1}(\Gamma))\subseteq{\mathrm{Gproj}}^{\leq n-k+1}(\Gamma)$
-where the last inclusion holds by inductive assumption. Combining with ${\rm{Gpd}}({}_{\Gamma}X^{0})\leq n-k$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n-k$ by[Ho04a, Theorem 2.20].
-$(b)\Rightarrow(a)$: Let $X\in{\rm{{mod}}}(\Gamma)$ and $k\geq 1$. Given a presentation of $X$ by finitely generated projective modules:
+where the last inclusion holds by inductive assumption. Combining with ${\rm{Gpd}}({}_{\Gamma}X^{0})\leq n-k$, we get ${\rm{Gpd}}({}_{\Gamma}Y)\leq n-k$ by [20, Theorem 2.20].
+$(b)\Rightarrow(a)$. Let $X\in{\rm{{mod}}}(\Gamma)$ and $k\geq 1$. Given a resolution of $X$ by finitely generated projective modules:
 $P_{k-1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{k-2}\longrightarrow\cdots\longrightarrow X\longrightarrow 0,$
@@
 Then $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra if ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is an abelian category, $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$, ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$.
-Before giving the proof, we would like to mention that torsion pairs play an important rule in the proof of Tachikawa’s theorem ([Ta74, Theorem 1]), Iyama’s theorem ([Iy04, Theorem 3.3]), and Kong’s theorem ([Ko14, Theorem 2.1]).
-We also follow this procedure and prove the result in four steps.
+Before giving the proof, we would like to mention that torsion pairs play an important rule in the proof of Tachikawa theorem ([44, Theorem 1]), Iyama’s theorem ([25, Theorem 3.3]), and Kong’s theorem ([33, Theorem 2.1]).
+We also follow this procedure and prove the result in three steps.
 Proof.
 Step I. We show that ${\rm{Ggldim}}(\Gamma)\leq n+1$.
-Since ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is abelian, it has kernels which, by Lemma 2.2, coincident with the kernels in ${\rm{{mod}}}(\Gamma)$.
+Since ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ is abelian, it has kernels which, by Lemma 3.2, coincides with the kernels in ${\rm{{mod}}}(\Gamma)$.
 Hence, ${\rm{Ggldim}}(\Gamma)\leq n+1$ follows from Lemma 3.3.
@@
 $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ $\Rightarrow\mathbf{E}(X)\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
-The idea is to show that ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is the torsion-free class of some hereditary torsion pair, and so is closed under taking injective envelopes [Di66, Theorem 2.9].
+The idea is to show that ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is the torsion-free class of some hereditary torsion pair, and so is closed under taking injective envelopes [16, Theorem 2.9].
 Obviously the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under extensions.
 By Lemma 3.3 ($b^{\prime}$), it is also closed under taking submodules.
-Hence, the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under taking extensions and submodules and so is the torsion-free class of some torsion pair $(\mathcal{T},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ in ${\rm{{mod}}}(\Gamma)$ [Di66, Theorem 2.3].
+Hence, the subcategory ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is closed under taking extensions and submodules and so is the torsion-free class of some torsion pair $(\mathcal{T},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ in ${\rm{{mod}}}(\Gamma)$ [16, Theorem 2.3].
 We now show that ${{}^{\perp_{0}}\Gamma}={{}^{{\perp}_{0}}{\mathrm{Gproj}}^{\leq n}(\Gamma)}=\mathcal{T}$.
 Let $X\in\mathcal{T}$.
-Then ${\rm{Hom}}_{\Gamma}(X,M)=0$ for $M=\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, we get $X\in{{}^{\perp_{0}}\Gamma}$.
-This shows that ${{}^{\perp_{0}}\Gamma}\supseteq\mathcal{T}$.
-On the other hand, suppose that $X\in{{}^{\perp_{0}}}\Gamma$.
+Then ${\rm{Hom}}_{\Gamma}(X,\Gamma)=0$ as $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, and so $X\in{{}^{\perp_{0}}\Gamma}$.
+This shows ${{}^{\perp_{0}}\Gamma}\supseteq\mathcal{T}$.
+On the other hand, suppose $X\in{{}^{\perp_{0}}}\Gamma$.
 By $^{\perp_{0}}\Gamma\subseteq{}^{\perp_{n}}\Gamma$, ${\rm{Ext}}^{i}_{\Gamma}(X,\Gamma)=0$ for every $0\leq i\leq n$.
 Hence, if $Y$ is a projective $\Gamma$-module, then ${\rm{Ext}}^{i}_{\Gamma}(X,Y)=0$ for every $0\leq i\leq n$.
-Thus, by a dimension shifting argument, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0={\rm{Ext}}^{1}_{\Gamma}(X,Y)$ for each $Y\in{\mathrm{Proj}}^{\leq n-1}(\Gamma)$.
+If $Y$ is a $\Gamma$-module with ${\rm{pd}}({}_{\Gamma}Y)=1$, then $Y$ has a projective resolution as the following.
+$0\longrightarrow P_{1}\longrightarrow P_{0}\longrightarrow Y\longrightarrow 0$
+Applying ${\rm{Hom}}_{\Gamma}(X,-)$ to this exact sequence gives the long exact sequence
+[…]
+and so ${\rm{Ext}}^{i}_{\Gamma}(X,Y)=0$ for every $0\leq i\leq n-1$.
+By applying this argument for finite steps, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0={\rm{Ext}}^{1}_{\Gamma}(X,Y)$ for each $Y\in\mathrm{proj}^{\leq n-1}(\Gamma)$.
 We also deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in\mathrm{proj}^{\leq n}(\Gamma)$.
-Also, if $Y$ is a Gorenstein projective $\Gamma$-module, then $Y$ is a submodule of a projective $\Gamma$-module and so ${\rm{Hom}}_{\Gamma}(X,Y)=0$.
-Let now $Y$ be a $\Gamma$-module with ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$. Then, by [Ho04a, Theorem 2.10], there exists an exact sequence
+Let now $Y$ be a $\Gamma$-module with ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$. Then, by [20, Theorem 2.10], there exists an exact sequence
 $0\longrightarrow K\longrightarrow G\longrightarrow Y\longrightarrow 0$
 such that $G$ is a Gorenstein projective module and $K$ is a $\Gamma$-module of projective dimension less than $n$.
-By applying ${\rm{Hom}}_{\Gamma}(X,-)$ to this exact sequence, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
+By applying ${\rm{Hom}}_{\Gamma}(X,-)$ to this exact sequence, we deduce that ${\rm{Hom}}_{\Gamma}(X,Y)=0$ for each $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ as $G$ is a submodule of a projective $\Gamma$-module and the projective dimension of $K$ is less than $n$.
 We have shown that ${{}^{\perp_{0}}\Gamma}=\mathcal{T}$.
@@
 Hence, ${{}^{\perp_{0}}}\Gamma$ is also closed under taking submodules and so the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a hereditary torsion pair in ${\rm{{mod}}}(\Gamma)$.
-This finishes the proof of Step 2.
+This finishes the proof of Step II.
 Step III. We show that ${\rm{domdim}}(\Gamma)\geq n+1$.
 By Step II, for every $X\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, $\mathbf{E}(X)\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
-On the other hand, as mentioned in Subsection 1.11, by [Ho04b, Theorem 2.2], the projective dimension of $\mathbf{E}(X)$ equals to its Gorenstein projective dimension and so $\mathbf{E}(X)\in\mathrm{proj}^{\leq n}(\Gamma)$.
+On the other hand, by [21, Theorem 2.2], the projective dimension of $\mathbf{E}(X)$ equals to its Gorenstein projective dimension and so $\mathbf{E}(X)\in\mathrm{proj}^{\leq n}(\Gamma)$.
 Hence, by the assumption, $\mathbf{E}(X)$ is a projective $\Gamma$-module.
@@
 with $I^{i}\in\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$ for each $0\leq i\leq n$ and hence ${\rm{domdim}}\ \Gamma\geq n+1$. ∎
-Propositions 2.1 and 2.4 could be enhanced to infinitely generated modules, that is, we could show the following result.
-Since its proof is analogous to that of Propositions 2.1 and 2.4, we only give a sketch of proof.
+Propositions 3.1 and 3.4 could be enhanced to infinitely generated modules, that is, we could show the following result.
+Since its proof is analogous to that of Propositions 3.1 and 3.4, we only give a sketch of proof.
 Proposition 3.5.
@@
 Let $X$ be an arbitrary injective $\Gamma$-module.
-By Facts 1.11 $(b)$, $X$ is a direct sum of indecomposable injective $\Gamma$-modules, the latter being finitely generated over an Artin algebra.
+By [32, Theorem 3.2], $X$ is a direct sum of indecomposable injective $\Gamma$-modules, the latter being finitely generated over an Artin algebra.
 From this fact, we deduce that each injective $\Gamma$-module is a direct sum of finitely generated injective $\Gamma$-modules.
@@
 $\Leftarrow$”.
-For this direction, the proof is nearly the same as the proof of the implication $\Leftarrow$” in Proposition 2.4 except that in Step II, one needs to observe that the subcategory ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is also closed under taking direct products and to use [Di66, Theorem 2.3] to conclude.
+For this direction, the proof is nearly the same as the proof of the implication $\Leftarrow$” in Proposition 3.4 except that in Step II, one needs to observe that the subcategory ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is also closed under taking direct products and to use [16, Theorem 2.3] to conclude.
 ∎
-Now by combining Propositions 2.1, 2.4, and 2.5 we have the following theorem that can be considered as a higher generalization of Auslander-Tachikawa theorem.
+Now by combining Propositions 3.1, 3.4, and 3.5 we have the following theorem that can be considered as a higher generalization of Auslander-Tachikawa theorem.
 Theorem 3.6.
@@
     ${\mathrm{GProj}}^{\leq n-1}(\Gamma)$ is an abelian category, ${\mathrm{Proj}}^{\leq n}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)$ and ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{}^{\underline{\perp}_{n}}\Gamma$;
-Based on the proof of Theorem 2.6 (in fact, Propositions 2.1, 2.4, and 2.5) we have the following results.
+Based on the proof of Theorem 3.6 (in fact, Propositions 3.1, 3.4, and 3.5) we have the following results.
 Corollary 3.7.
@@
 Proof.
-In Step 2 of the proof of Proposition 2.4, we showed that the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a hereditary torsion pair.
+In Step II of the proof of Proposition 3.4, we showed that the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a hereditary torsion pair.
 A similar argument shows that the pair $({{}^{\underline{\perp}_{0}}\Gamma},{\mathrm{GProj}}^{\leq n}(\Gamma))$ is also a hereditary torsion pair.
@@
 Similarly, for every non-negative integer $i$, ${\mathrm{Gproj}}^{\leq i}(\Gamma)$ is also closed under taking injective envelopes. ∎
-Kong used of the notion of submodule categories and characterized Artin algebras whose Gorenstein projective modules form an abelian category; see [Ko14, Corollary 2.3].
+Kong used of the notion of submodule categories and characterized Artin algebras whose Gorenstein projective modules form an abelian category; see [33, Corollary 2.3].
 In the following, we will prove a similar property for higher Auslander-Gorenstein algebras.
@@
 ∎
-A combination of our results and some classical results about torsion pairs gives us a better perspective about higher Auslander-Gorenstein algebras, see Corollaries 2.10 and 2.11.
+A combination of our results and some classical results about torsion pairs gives us a better perspective about higher Auslander-Gorenstein algebras, see Corollaries 3.10 and 3.11.
 Corollary 3.10.
@@
 By Corollary 3.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair.
-By [AR94, Proposition 5.8], ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is a functorially finite subcategory of ${\rm{{mod}}}(\Gamma)$ and so by Theorem 1.4 that we recalled it in the Subsection 1.5, all the statements hold and also are equivalent with each other.
+By [4, Proposition 5.8], ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is a functorially finite subcategory of ${\rm{{mod}}}(\Gamma)$ and so by [41, Theorem] and [1, Proposition 1.1], all the statements hold and also are equivalent with each other.
 Note that $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ and so $\mathrm{Ann}_{\Gamma}({\mathrm{Gproj}}^{\leq n}(\Gamma))=0$.
-But, by [Sm84, Lemma 0.2], $\mathrm{Ann}_{\Gamma}({{}^{\perp_{0}}\Gamma})\neq 0$.
+But, by [41, Lemma 0.2], $\mathrm{Ann}_{\Gamma}({{}^{\perp_{0}}\Gamma})\neq 0$.
 ∎
@@
 Proof.
-By Corollary 2.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair and so, by [Ho82, Lemmas 1,2, and 3], all the statements hold; see also [AS81, Corollaries 3.4 and 3.7].
+By Corollary 3.7, the pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair and so, by [22, Lemmas 1,2, and 3], all the statements hold; see also [6, Corollaries 3.4 and 3.7].
 ∎
@@
 Let $n$ be a positive integer, $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra, and $t$ be the idempotent radical corresponding to the torsion pair $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$.
-Moreover, suppose $X\in{{}^{\perp_{0}}\Gamma}$ and $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ are indecomposable $\Gamma$-modules.Then the following statements hold.
+Moreover, suppose $X\in{{}^{\perp_{0}}\Gamma}$ and $Y\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$ are indecomposable $\Gamma$-modules.
+Then the following statements hold.
   • (i)
@@
 Proof.
-By [Ho82, Lemmas 2 and 3], all the statements hold; see also [AS81, Corollaries 3.4 and 3.7]. ∎
-3. Selfinjective and $\tau_{n}$-selfinjective Algebras
-As an application of our results in Section 2, in this section we study selfinjective and $\tau_{n}$-selfinjective algebras and characterize them in a categorical sense.
-We first have the following result that gives a categorical description for when an $n$-minimal Auslander-Gorenstein algebra $\Gamma$ is self-injective.
-Recall that an Artin algebra $\Gamma$ is said to be selfinjective if it is injective as a left $\Gamma$-module.
-Proposition 3.1.
-Let $n$ be a positive integer and an Artin algebra $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra. Then the following statements are equivalent:
-  • $(a)$
-    $\Gamma$ is selfinjective;
-  • $(b)$
-    ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category;
-  • $(c)$
-    ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$;
-  • $(d)$
-    ${\mathrm{GProj}}^{\leq n}(\Gamma)$ is an abelian category;
-  • $(e)$
-    ${{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{n+1}}\Gamma}$.
-Proof.
-$(a)\Leftrightarrow(b)$.
-Let ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ be an abelian category and $\Gamma$ be non-selfinjective.
-Then ${\rm{Ggldim}}(\Gamma)$ $=n+1$, so there exists a finitely generated $\Gamma$-module $X$ such that ${\rm{Gpd}}({}_{\Gamma}X)=n+1$.
-Let
-$0\to K\stackrel{{\scriptstyle\kappa}}{{\to}}P_{0}\stackrel{{\scriptstyle\varepsilon}}{{\longrightarrow}}X\longrightarrow 0$
-be a short exact sequence with $P_{0}\in\mathrm{proj}(\Gamma)$.
-Then $K\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
-If $\pi:P_{0}\longrightarrow Y$ is the cokernel of the natural monomorphism $\kappa$ in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$, then the sequence
-$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}Y$
-is an exact sequence in ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ as ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category. Since $\Gamma\in{\mathrm{Gproj}}^{\leq n}(\Gamma)$, the sequence
-$0\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,K)\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,P_{0})\longrightarrow{\rm{Hom}}_{\Gamma}(\Gamma,Y)$
-of abelian groups is also exact and so the sequence
-$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{0}\stackrel{{\scriptstyle\pi}}{{\longrightarrow}}Y$
-is also exact in ${\rm{{mod}}}(\Gamma)$.
-Thus the induce map $\overline{\pi}:X\simeq P_{0}/K\longrightarrow Y$ is a monomorphism.
-Hence, $X$ is a submodule of $Y$ and so there exists a short exact sequence
-$0\longrightarrow X\longrightarrow Y\longrightarrow Z\longrightarrow 0$
-in ${\rm{{mod}}}(\Gamma)$ such that ${\rm{Gpd}}({}_{\Gamma}Y)\leq n$.
-Since ${\rm{Ggldim}}(\Gamma)=n+1$, ${\rm{Gpd}}({}_{\Gamma}Z)\leq n+1$ and so ${\rm{Gpd}}({}_{\Gamma}X)\leq n$, which is a contradiction.
-We have shown that if ${\mathrm{Gproj}}^{\leq n}(\Gamma)$ is an abelian category, then $\Gamma$ is selfinjective. The other direction is trivial.
-$(a)\Leftrightarrow(c)$.
-Assume that ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$.
-Suppose that $\Gamma$ has Gorenstein global dimension $n+1$.
-We first show that ${{}^{\perp_{n+1}}\Gamma}\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
-Let $X\in{}^{\perp_{n+1}}\Gamma$.
-Then ${\rm{Ext}}_{\Gamma}^{n+1}(X,\Gamma)=0$.
-Since ${\rm{Ggldim}}(\Gamma)=n+1$, ${\rm{Gpd}}({}_{\Gamma}X)\leq n$ and hence ${}^{\perp_{n}}\Gamma\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
-Then ${{}^{\perp_{0}}\Gamma}\subseteq{}^{\perp_{n+1}}\Gamma\subseteq{\mathrm{Gproj}}^{\leq n}(\Gamma)$.
-By Corollary 2.7, $({{}^{\perp_{0}}\Gamma},{\mathrm{Gproj}}^{\leq n}(\Gamma))$ is a torsion pair in ${\rm{{mod}}}(\Gamma)$ and so ${{}^{\perp_{0}}\Gamma}=0$.
-Therefore, ${\mathrm{Gproj}}^{\leq n}(\Gamma)={\rm{{mod}}}(\Gamma)$, which contradicts the assumption ${\rm{Ggldim}}(\Gamma)=n+1$.
-So ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n+1}}\Gamma}$ implies that $\Gamma$ is selfinjective.
-The other direction is trivial.
-$(a)\Leftrightarrow(d)$. The proof is similar to that of $(a)\Leftrightarrow(b)$.
-$(a)\Leftrightarrow(e)$. The proof is similar to that of $(a)\Leftrightarrow(c)$. ∎
-Now we can give a categorical description for the selfinjectivity of an Artin algebra.
-Corollary 3.2.
-Let $\Gamma$ be an Artin algebra.
-Then $\Gamma$ is selfinjective if and only if the categories ${\mathrm{Gproj}}(\Gamma)$ and ${\mathrm{Gproj}}^{\leq 1}(\Gamma)$ are abelian categories if and only if the categories ${\mathrm{GProj}}(\Gamma)$ and ${\mathrm{GProj}}^{\leq 1}$ $(\Gamma)$ are abelian categories.
-Proof.
-Let the categories ${\mathrm{Gproj}}(\Gamma)$ and ${\mathrm{Gproj}}^{\leq 1}(\Gamma)$ be abelian categories.
-By [Ko14, Theorem 2.1], $\Gamma$ is an Auslander algebra and hence by Proposition 3.1, $\Gamma$ is selfinjective.
-For the second equivalence, the proof is similar.
-∎
-We now turn our attention to $\tau_{n}$-selfinjective algebras. Let us recall their definition.
+By [22, Lemmas 2 and 3], all the statements hold; see also [6, Corollaries 3.4 and 3.7]. ∎
+4. $\tau_{n}$-selfinjective Algebras
+As an application of our results in Section 3, in this section we study $\tau_{n}$-selfinjective algebras and characterize them in a categorical sense. Let us recall their definition.
 Definition 4.1 ($\tau_{n}$-selfinjective algebras).
@@
 The concept of generating subcategories is defined dually.
-We also call $\mathcal{X}$ $n$-rigid if ${\rm{Ext}}^{i}_{\Lambda}(\mathcal{X},\mathcal{X})=0$ for each $0<i<n$.
-Let now $\tau_{n}$ and $\tau_{n}^{-}$ be the $n$-Auslander-Reiten translations; see [Iy07a, Subsection 1.4] or [Iy11, Definition 1.1].
-A generating-cogenerating $n$-rigid subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Lambda)$ is called $n$-precluster tilting subcategory if it is functorially finite, $\tau_{n}(\mathcal{X})\subseteq\mathcal{X}$, and $\tau_{n}^{-}(\mathcal{X})\subseteq\mathcal{X}$ [IS18, Definition 3.2].
+Also $\mathcal{X}$ is called $n$-rigid if ${\rm{Ext}}^{i}_{\Lambda}(\mathcal{X},\mathcal{X})=0$ for each $0<i<n$.
+Let now $\tau_{n}$ and $\tau_{n}^{-}$ be the $n$-Auslander-Reiten translations; see [26, Subsection 1.4] or [28, Definition 1.1].
+A generating-cogenerating $n$-rigid subcategory $\mathcal{X}$ of ${\rm{{mod}}}(\Lambda)$ is called $n$-precluster tilting subcategory if it is functorially finite, $\tau_{n}(\mathcal{X})\subseteq\mathcal{X}$, and $\tau_{n}^{-}(\mathcal{X})\subseteq\mathcal{X}$ [31, Definition 3.2].
 Moreover, if $\mathcal{X}$ admits an additive generator $M$, we say that $\mathcal{X}$ is a finite $n$-precluster tilting subcategory and $M$ is an $n$-precluster tilting module.
-By following Iyama-Solberg [IS18, Definition 3.4], Artin algebras which have $n$-precluster tilting modules are also called $\tau_{n}$-selfinjective.
+By following Iyama-Solberg [31, Definition 3.4], Artin algebras which have $n$-precluster tilting modules are also called $\tau_{n}$-selfinjective.
 […]
@@
 $\tau_{n}$-selfinjective algebras.
-We need the following lemma which is of independent interest and follows from Lemmas 1.5 and 1.6.
+We need the following lemma which is of independent interest and follows from Lemmas 2.4 and 2.5.
 Lemma 4.2.
@@
   • (a)
-    The functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$, which further restricts to another equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$, so $\mathrm{add}(\mathcal{Q})$ $=\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$.
+    The functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ restricts to an equivalence $\mathrm{add}(M)\simeq\mathrm{proj}(\Gamma)$, which further restricts to another equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$, so $\mathrm{add}(\mathcal{Q})$ $\mathrm{proj}(\Gamma)\cap\mathrm{inj}(\Gamma)$.
   • (b)
@@
 Proof.
-The statement $(a)$ can be proved as in [ARS, chapter 5, Lemma 5.3].
+The statement $(a)$ can be proved as in [5, Lemma V.5.3].
 For $(b)$, the fact that $M_{\Gamma}$ is projective follows from the duality $\mathrm{Hom}_{\Lambda}(-,M)$ between $\mathrm{add}(M)$ and $\mathrm{proj}(\Gamma^{op})$ and the fact that $_{\Lambda}\Lambda\in\mathrm{add}(M)$.
 Other assertions deduce from $(a)$ and the fact that $\nu^{-1}$ establishes an equivalence between (finitely generated) projective modules and injective modules.
-For $(c)$, the two functors $\mathrm{Hom}_{\Gamma}(I,-)$ and $D\mathrm{Hom}_{\Gamma}(-,Q)$ from ${\rm{{mod}}}(\Gamma)$ to ${\rm{{mod}}}(\Lambda)$ are right exact (in fact exact), so by Watts’ theorem, one need to verify that they coincident with $M\otimes_{\Gamma}-$ on the left regular module ${}_{\Lambda}\Lambda$, which is obvious by $Q=\mathrm{Hom}_{\Lambda}(M,D\Lambda)\simeq D(M)$ and $I\simeq\mathrm{Hom}_{\Gamma}(M,\Gamma)\simeq\nu^{-1}Q$.
+For $(c)$, the two functors $\mathrm{Hom}_{\Gamma}(I,-)$ and $D\mathrm{Hom}_{\Gamma}(-,Q)$ from ${\rm{{mod}}}(\Gamma)$ to ${\rm{{mod}}}(\Lambda)$ are right exact (in fact exact), so by Watts’ theorem, one need to verify that they coincides with $M\otimes_{\Gamma}$- on the left regular module $_{\Lambda}\Lambda$, which is obvious by $Q=\mathrm{Hom}_{\Lambda}(M,D\Lambda)\simeq D(M)$ and $I\simeq\mathrm{Hom}_{\Gamma}(M,\Gamma)\simeq\nu^{-1}Q$.
 The existence of the adjoint triple follows from the usual tensor-Hom adjunction.
 By $(a)$, the functor $\mathrm{Hom}_{\Lambda}(M,-):{\rm{{mod}}}(\Lambda)\to{\rm{{mod}}}(\Gamma)$ restricts to an equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{add}(Q)$, with quasi-inverse $D\mathrm{Hom}_{\Gamma}(-,Q)$.
-By the first lemma of [Au71, Chapter III, Section 4, Page 48].
+By the first lemma of [3, Chapter III, Section 4, Page 48].
 this equivalence can be extended to an equivalence ${\rm{{mod}}}(\Lambda)\simeq\mathrm{sub}^{2}(\mathcal{Q})$, as $Q$ is injective.
 This proves $(d)$.
-For $(d^{\prime})$, by Fact 1.11 $(a)(b)$, the equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{add}(Q)$ can be extended to ${\mathrm{Inj}}(\Lambda)\simeq\mathrm{Add}(Q)$.
+For $(e)$, by [13, p.
+17, Exercise 8] and [32, Theorem 3.2], the equivalence $\mathrm{inj}(\Lambda)\simeq\mathrm{add}(Q)$ can be extended to ${\mathrm{Inj}}(\Lambda)\simeq\mathrm{Add}(Q)$.
 Again the latter can be extended to ${\rm{Mod}}(\Lambda)\simeq\mathrm{Sub}^{2}(\mathcal{Q})$.
@@
 $(b)\Rightarrow(a)$.
-By Proposition 2.4, $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra and so by Iyama-Solberg correspondence [IS18, Theorem 4.5], there is an Artin algebra $\Lambda^{\prime}$ with $n$-precluster tilting $\Lambda^{\prime}$-module $M^{\prime}$ such that ${\rm{{mod}}}(\Lambda^{\prime})\simeq{\mathrm{Gproj}}^{\leq n-1}(\Gamma)\simeq{\rm{{mod}}}(\Lambda)$.
+By Proposition 3.4, $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra and so by Iyama-Solberg correspondence [31, Theorem 4.5], there is an Artin algebra $\Lambda^{\prime}$ with $n$-precluster tilting $\Lambda^{\prime}$-module $M^{\prime}$ such that ${\rm{{mod}}}(\Lambda^{\prime})\simeq{\mathrm{Gproj}}^{\leq n-1}(\Gamma)\simeq{\rm{{mod}}}(\Lambda)$.
 Hence, $\Lambda$ and $\Lambda^{\prime}$ are also Morita equivalent.
@@
 Remark 4.4.
-Lemmas 1.6 and 3.4, allowed us to generalize our results from finitely generated modules to all modules, see Theorems 2.6 and 3.5.
+Lemmas 2.5 and 4.2, allowed us to generalize our results from finitely generated modules to all modules, see Theorems 3.6 and 4.3.
 Of course, we are not limited to these lemmas and can follow other ways.
@@
 Recall that a $\Gamma$-module $\mathcal{Q}$ is product-complete, if ${\rm{Add}}(\mathcal{Q})={\rm Prod}(\mathcal{Q})$, where $\mathrm{Prod}(\mathcal{Q})$ is the smallest subcategory of ${\rm{Mod}}(\Gamma)$ containing $Q$ and closed under products and direct summands.
-By [KS98, Proposition 3.5], there is a (locally Noetherian) Grothendieck category $\mathcal{A}$ and a functor $F:\mathcal{A}\longrightarrow{\rm{Mod}}(\Gamma)$ such that $F$ induces an equivalence between ${\mathrm{Inj}}(\mathcal{A})$ and ${\rm{Add}}(\mathcal{Q})$.
-By the proof of [KS98, Proposition 3.5], the functor $F$ is a composition of the right adjoint of a quotient functor with an exact functor and so $F$ is left exact.
+By [34, Proposition 3.5], there is a (locally Noetherian) Grothendieck category $\mathcal{A}$ and a functor $F:\mathcal{A}\longrightarrow{\rm{Mod}}(\Gamma)$ such that $F$ induces an equivalence between ${\mathrm{Inj}}(\mathcal{A})$ and ${\rm{Add}}(\mathcal{Q})$.
+By the proof of [34, Proposition 3.5], the functor $F$ is a composition of the right adjoint of a quotient functor with an exact functor and so $F$ is left exact.
 Hence, $F$ induces an equivalence between $\mathcal{A}$ and $\mathrm{Sub}^{2}(\mathcal{Q})$.
@@
 5. Cotorsion Pairs
-As mentioned in the introduction, this section is devoted to study the relation between higher Auslander-Gorenstein and cotorsion pairs and we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS].
+As mentioned in the introduction, this section is devoted to study the relation between higher Auslander-Gorenstein and cotorsion pairs and we present some results on higher Auslander-Gorenstein algebras from the viewpoint of cotorsion pairs and relate them to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [11].
 Let $n$ and $i$ be a non-negative integer. For an $n$-minimal Auslander-Gorenstein algebra $\Gamma$, some results by Martinez Villa show that the pair
 $({\rm{dom}}^{\geq i}(\Gamma),\mathrm{inj}^{\leq i}(\Gamma))$
-is a hereditary cotorsion pair in ${\rm{{mod}}}(\Gamma)$, where $\rm{dom}^{\geq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules having the dominant dimension at least $i$ and $\mathrm{inj}^{\leq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules with injective dimension at most $i$, see [Ma92, Lemma 3, Proposition 5, and the corollary before it].
-A result by Marczinzik shows that ${\mathrm{Gproj}}^{\leq n+1-i}(\Gamma)=\rm{dom}^{\geq i}(\Gamma)$ (see, [Ma18, Theorem 2.1]) and so, by [ET01, Theorem 10] or Salce’s lemma and [Ma92, Proposition 7], the pair
+is a hereditary cotorsion pair in ${\rm{{mod}}}(\Gamma)$, where ${\rm{dom}}^{\geq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules having the dominant dimension at least $i$ and $\mathrm{inj}^{\leq i}(\Gamma)$ denotes the full subcategory of ${\rm{{mod}}}(\Gamma)$ consisting of all finitely generated $\Gamma$-modules with injective dimension at most $i$, see [36, Lemma 3, Proposition 5, and the corollary before it].
+A result by Marczinzik shows that ${\mathrm{Gproj}}^{\leq n+1-i}(\Gamma)={\rm{dom}}^{\geq i}(\Gamma)$ (see, [35, Theorem 2.1]) and so, by [17, Theorem 10] or Salce’s lemma and [36, Proposition 7], the pair
 $({\mathrm{Gproj}}^{\leq n+1-i}(\Gamma),\mathrm{inj}^{\leq i}(\Gamma))$
@@
 Proposition 5.1.
-Let $n$ be a positive integer, $10\leq k\leq n+$ 1, $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra; and $Q$ be its maximal projective-injective summand.
+Let $n$ be a positive integer, $0\leq k\leq n+1$, $\Gamma$ be an $n$-minimal Auslander-Gorenstein algebra; and $Q$ be its maximal projective-injective summand.
 Then $T_{k}:=\Omega^{-k}(\Gamma)\oplus Q$ is a $k$-tilting module and ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$.
@@
 $\mathrm{sub}^{j}(T_{k})={\rm{dom}}^{\geq j}(\Gamma)={\mathrm{Gproj}}^{\leq n+1-j}(\Gamma).$
-Especially, for every $10\leq i\leq n+$ 1,
+Especially, for every $0\leq i\leq n+1$,
 $\mathrm{sub}^{i}(T_{n+1-i})={\rm{dom}}^{\geq i}(\Gamma)={\mathrm{Gproj}}^{\leq n+1-i}(\Gamma).$
@@
 be the minimal injective resolution of $\Gamma$.
-Since $I^{0},I^{1},\cdots,I^{n}\in{\rm{add}}(\mathcal{Q})$, clearly, $\Gamma$ has an ${\rm{add}}(T_{k})$-copresentation of length $k$ and for every $10\leq k\leq n+$ 1, $\Omega^{-k}(\Gamma)\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$.
-Hence, for every $11\leq l\leq n+$ 1,
+Since $I^{0},I^{1},\cdots,I^{n}\in{\rm{add}}(\mathcal{Q})$, clearly, $\Gamma$ has an ${\rm{add}}(T_{k})$-copresentation of length $k$ and for every $0\leq k\leq n+1$, $\Omega^{-k}(\Gamma)\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$.
+Hence, for every $1\leq l\leq n+1$,
 ${\rm{Ext}}^{l}_{\Gamma}(T_{k},T_{k})\simeq{\rm{Ext}}^{l}_{\Gamma}(\Omega^{-k}(\Gamma),\Omega^{-k}(\Gamma))\simeq{\rm{Ext}}^{l}_{\Gamma}(\Omega^{k}(\Omega^{-k}(\Gamma)),\Gamma)={\rm{Ext}}^{l}_{\Gamma}(\Gamma,\Gamma)=0.$
-Therefore, $T_{k}$ is a $k$-tilting module; see [HU96, Page 448] or [AC01, Section 2] for the definition.
+Therefore, $T_{k}$ is a $k$-tilting module; see [19, Page 448] or [2, Section 2] for the definition.
 We now show that for every $0\leq j\leq n+1-k$, $\mathrm{sub}^{j}(T_{k})={\rm{dom}}^{\geq j}(\Gamma)={\mathrm{Gproj}}^{\leq n+1-j}(\Gamma).$
@@
 Conversely, ${\mathrm{Gproj}}^{\leq n+1-j}(\Gamma)={\rm{dom}}^{\geq j}(\Gamma)=\mathrm{sub}^{j}(\mathcal{Q})\subseteq\mathrm{sub}^{j}(T_{k}).$
-To complete the proof, we must show that for every $10\leq k\leq n+$ 1, ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$.
+To complete the proof, we must show that for every $0\leq k\leq n+1$, ${\rm{add}}(T_{k})={\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$.
 Clearly, ${\rm{add}}(T_{k})\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so we only need to prove the converse.
 Let $X\in{\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$.
-By [Ho04b, Theorem 2.2], ${\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)=\mathrm{proj}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so ${\rm{pd}}(_{\Gamma}X)\leq k$.
-Hence, by [AC01, Lemma 2.2 $(a)$], ${}^{\perp}(T_{k}^{\perp})\subseteq\mathrm{proj}^{\leq k}(\Gamma)\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)$ and so
+By [21, Theorem 2.2], ${\mathrm{Gproj}}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)=\mathrm{proj}^{\leq k}(\Gamma)\cap\mathrm{inj}^{\leq n+1-k}(\Gamma)$ and so ${\rm{pd}}(_{\Gamma}X)\leq k$.
+Hence, by [2, Lemma 2.2 $(a)$], $^{\perp}(T_{k}^{\perp})\subseteq\mathrm{proj}^{\leq k}(\Gamma)\subseteq{\mathrm{Gproj}}^{\leq k}(\Gamma)$ and so
 $X\in\mathrm{inj}^{\leq n+1-k}(\Gamma)=({\mathrm{Gproj}}^{\leq k}(\Gamma))^{\perp}\subseteq(^{\perp}(T_{k}^{\perp}))^{\perp}=T_{k}^{\perp}.$
-On the other hand, by [AC01, Lemmas 2.3 $(a)$ and 2.4 $(a)$], for every $Y\in T_{k}^{\perp}$, there is a short exact sequence
+On the other hand, by [2, Lemmas 2.3 $(a)$ and 2.4 $(a)$], for every $Y\in T_{k}^{\perp}$, there is a short exact sequence
 $0\longrightarrow K_{0}\longrightarrow T^{\prime}_{0}\longrightarrow Y\longrightarrow o,$
@@
 Therefore, $X\in{{}^{\perp}(T_{k}^{\perp})}$.
-But by [AC01, Lemmas 2.3 $(a)$ and 2.4 $(a)$] there is a short exact sequence
+But by [2, Lemmas 2.3 $(a)$ and 2.4 $(a)$] there is a short exact sequence
 $0\longrightarrow K\longrightarrow T\longrightarrow X\longrightarrow o,$
@@
 Among the above mentioned cotorsion pairs, one of them has a very nice characteristic and its cotorsion class is a torsion-free class of a hereditary torsion pair.
-This fact relates higher Auslander-Gorenstein Algebras to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [BBOS] and gives us a nice description of them.
+This fact relates higher Auslander-Gorenstein Algebras to the notion of torsion-cotorsion triples introduced by Bauer, Botnan, Oppermann and Steen [11] and gives us a nice description of them.
 We first recall some definitions and results in this direction and then by using them we will establish two equivalences for $n$-minimal Auslander-Gorenstein algebras.
 Definition 5.2 (Torsion Cotorsion Triple).
-A torsion cotorsion triple in an abelian category $\mathcal{A}$ is a triple of subcategories $(\mathcal{T},\mathcal{F},\mathcal{D})$ such that the pair $(\mathcal{T},\mathcal{F})$ is a torsion pair and the pair $(\mathcal{F},\mathcal{D})$ is a cotorsion pair, see [BBOS, Page 29 before Theorem 2.33].
-Theorem 4.3 (Dual of [BBOS, Theorem 2.35]).
+A torsion cotorsion triple in an abelian category $\mathcal{A}$ is a triple of subcategories $(\mathcal{T},\mathcal{F},\mathcal{D})$ such that the pair $(\mathcal{T},\mathcal{F})$ is a torsion pair and the pair $(\mathcal{F},\mathcal{D})$ is a cotorsion pair, see [11, Page 29 before Theorem 2.33].
+Theorem 5.3 (Dual of [11, Theorem 2.35]).
 Let $\Gamma$ be an Artin algebra and $(\mathcal{T},\mathcal{F},\mathcal{D})$ be a torsion cotorsion triple in ${\rm{{mod}}}(\Gamma)$.
@@
 Proof.
-This theorem is the dual version of [BBOS, Theorem 2.35] and for the convenience of the reader we give the proof.
-By [ARS, Proposition IV.1.9], we know that the inverse of Auslander-Reiten translation $\tau^{-}$ defines an equivalence
-[…]
+This theorem is the dual version of [11, Theorem 2.35] and for the convenience of the reader we give the proof.
+By [5, Proposition IV.1.9], we know that the inverse of Auslander-Reiten translation $\tau^{-}$ defines an equivalence
+$\overline{{\rm{{mod}}}}(\Gamma)$$\underline{{\rm{{mod}}}}(\Gamma).$$\scriptstyle{\tau^{-}=\mathrm{Tr}\mathbf{D}}$
 Clearly, $\frac{\mathcal{D}}{\mathrm{inj}(\Gamma)}$ is a full subcategory of $\overline{{\rm{{mod}}}}(\Gamma)$ and one can easily see that $\mathcal{T}$ is a full subcategory of $\underline{{\rm{{mod}}}}(\Gamma)$.
@@
 $X\in\mathcal{D}\Longleftrightarrow{\rm{Ext}}^{1}_{\Gamma}(\mathcal{F},X)=0\Longleftrightarrow{\rm{Hom}}_{\Gamma}(\tau^{-}X,\mathcal{F})=0\Longleftrightarrow\tau^{-}X\in\mathcal{T},$
-where the middle equivalence follows from Auslander-Reiten formula because the injective dimension of every element of $\mathcal{D}$ by Bauer-Botnan-Oppermann-Steen Correspondence is at most one, see [SY11, Corollary III.6.4] and [BBOS, Theorem 2.34].
+where the middle equivalence follows from Auslander-Reiten formula because the injective dimension of every element of $\mathcal{D}$ by Bauer-Botnan-Oppermann-Steen Correspondence is at most one, see [42, Corollary III.6.4] and [11, Theorem 2.34].
 ∎
@@
 Since $\Gamma$ is an $n$-minimal Auslander-Gorenstein algebra, the triple $(^{\perp_{0}}\Gamma,{\mathrm{Gproj}}^{\leq n}(\Gamma),\mathrm{inj}^{\leq 1}(\Gamma))$ is a torsion cotorsion triple, see Corollary 3.7 and the arguments at the beginning of this section.
-Hence, by [BBOS, Theorem 2.33], we have the first equivalence.
+Hence, by [11, Theorem 2.33], we have the first equivalence.
 That is
@@
 $\Gamma$ is also called a higher Auslander algebra, if there is a positive integer $n$ such that $\Gamma$ is an $n$-Auslander algebra.
-Iyama introduced these algebras as a generalization of classical Auslander algebras [Iy11, Page 3].
+Iyama introduced these algebras as a generalization of classical Auslander algebras [28, Page 3].
 As a result of Theorem 3.6, we have the following result for higher Auslander algebras.
@@
 Remark 6.2.
-Let $\Gamma$ be an Artin algebra and $n$ be a positive integer. Then we have the following statements:
-  • $(a)$
-    If $\mathrm{proj}(\Gamma)$ is an abelian category, then
-    $\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)\ \mathrm{and}\ {{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{1}}\Gamma};$
-  • $(b)$
-    If ${\mathrm{Proj}}(\Gamma)$ is an abelian category, then
-    ${\mathrm{Proj}}^{\leq 1}(\Gamma)\cap{\mathrm{Inj}}(\Gamma)\subseteq{\mathrm{Proj}}(\Gamma)\ \mathrm{and}\ {{}^{\underline{\perp}_{0}}\Gamma}\subseteq{{}^{\underline{\perp}_{1}}\Gamma}.$
-Hence, Theorem 5.1 is precisely a generalization of Auslander-Tachikawa theorem (Theorem 1) and also provide a new proof for it.
-Let us prove $(a)$, that of $(b)$ being similar.
-In fact, assume that $\mathrm{proj}(\Gamma)$ is an abelian category.
-By Step 1 of the proof of Proposition 2.4, we see that $\mathrm{proj}(\Gamma)$ has the same kernels as ${\rm{{mod}}}(\Gamma)$, so $\mathrm{gldim}(\Gamma)\leq 2$.
-Let $X\in\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)$.
-Since $\mathrm{proj}(\Gamma)$ is an abelian category, by the proof of Proposition 3.1, there exists a projective $\Gamma$-module $Y$ such that $X$ is a submodule of $Y$.
-But $X$ is an injective $\Gamma$-module and so it is a direct summand of $Y$.
-Therefore, $X$ is a projective $\Gamma$-module.
-This shows that $\mathrm{proj}^{\leq 1}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$.
-Let now $X\in{{}^{{\perp}_{0}}}\Gamma$. Then, there is an exact sequence
-$0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{0}\longrightarrow X\longrightarrow 0$
-where $P_{0},P_{1}\in\mathrm{proj}(\Gamma)$.
-Since $X\in{{}^{{\perp}_{0}}}\Gamma$, ${\rm{Hom}}_{\Gamma}(X,P)=0$ for every finitely generated projective $\Gamma$-module $P$, $\beta$ is a surjective map in $\mathrm{proj}(\Gamma)$.
-Since $\mathrm{gldim}(\Gamma)\leq 2$, $K$ is projective.
-Thus the sequence $0\longrightarrow K\stackrel{{\scriptstyle\kappa}}{{\longrightarrow}}P_{1}\stackrel{{\scriptstyle\beta}}{{\longrightarrow}}P_{0}\longrightarrow 0$ is exact in $\mathrm{proj}(\Gamma)$ and so by the exact sequence of abelian groups
-$0\longrightarrow{\rm{Hom}}_{\Gamma}(P_{0},\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(P_{1},\Gamma)\longrightarrow{\rm{Hom}}_{\Gamma}(K,\Gamma)$
-one can easily see that ${\rm{Ext}}^{1}_{\Gamma}(X,\Gamma)=0$. Thus the condition ${{}^{{\perp}_{0}}}\Gamma\subseteq{{}^{{\perp}_{1}}}\Gamma$ holds.
-Remark 5.3.
 For $n\geq 2$, the conditions $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ and ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ in Theorem 6.1(b) are necessary.
-In fact, take an Artin algebra of global dimension $n-1$ which is not an $n$-Auslander algebra.
+In fact, take an Artin algebra of global dimension n1n-1 which is not an $n$-Auslander algebra.
 Then $\mathrm{proj}^{\leq n-1}(\Gamma)={\rm{{mod}}}(\Gamma)$ is an abelian category, but by Theorem 6.1, the conditions that $\mathrm{proj}^{\leq n}(\Gamma)\cap\mathrm{inj}(\Gamma)\subseteq\mathrm{proj}(\Gamma)$ or ${{}^{\perp_{0}}\Gamma}\subseteq{{}^{\perp_{n}}\Gamma}$ don’t hold.
@@
 Recall that $\mathcal{M}$ is an $n$-cluster-tilting subcategory of ${\rm{{mod}}}(\Lambda)$ if it is functorially finite in ${\rm{{mod}}}(\Lambda)$ and
-[…]
+$\begin{array}[]{rcl}\mathcal{M}&=&\{X\in{\rm{{mod}}}(\Lambda)\ \ |\ \ {\rm{Ext}}^{i}_{\Lambda}(X,\mathcal{M})=0,\ \forall\ 0<i<n\}\\ &=&\{X\in{\rm{{mod}}}(\Lambda)\ \ |\ \ {\rm{Ext}}^{i}_{\Lambda}(\mathcal{M},X)=0,\ \forall\ 0<i<n\}.\end{array}$
 Moreover, if $\mathcal{M}$ admits an additive generator $M$, i.e.
@@
 The following example gives a better perspective about some results of the paper.
-In the following, for every vertex $i$ of the given quiver $Q$, the corresponding simple, projective, and injective object will be deonted by $S(i)$, $P(i)$, and $I(i)$, respectively.
-For more details about quiver representations, see for example [SY11].
+In the following, for every vertex $i$ of the given quiver $Q$, the corresponding simple, projective, and injective object will be denoted by $S(i)$, $P(i)$, and $I(i)$, respectively.
+For more details about quiver representations, see for example [42].
 Example 6.6.
 Let $\Gamma=KQ/I$ be an algebra with the quiver $Q$:
-{tikzpicture}
+[…]
 and $I$ is the admissible ideal of $KQ$ generated by paths $\alpha_{i+1}\alpha_{i}$ for every $1\leq i\leq n$. Then:
   1. (1)
-    By [SY11, Theorem I.10.5], the Nakayama algebra $\Gamma$ is of finite representation type with $32n+$ 3 indecomposable $\Gamma$-modules.
+    By [42, Theorem I.10.5], the Nakayama algebra $\Gamma$ is of finite representation type with $2n+3$ indecomposable $\Gamma$-modules.
 The following table gives us more information about their homological dimensions.
     Ind. Modules pd id
-    $P(1)$ 00 $n+1$
-    $P(2)=I(3)$ 00 00
+    $P(1)$ 0 $n+1$
+    $P(2)=I(3)$ 0 0
     $\vdots$ $\vdots$ $\vdots$
-    $P(n)=I(n+1)$ 00 00
-    $P(n+1)=I(2)$ 00 00
-    $I(1)$ $n+1$ 00
+    $P(n)=I(n+1)$ 0 0
+    $P(n+1)=I(2)$ 0 0
+    $I(1)$ $n+1$ 0
     $S(1)$ $n+1$ $n+1$
-    $S(2)$ $n$ $1$
+    $S(2)$ $n$ 11
     $\vdots$ $\vdots$ $\vdots$
-    $S(j)\ (2\leq j\leq n+1)$ $n+2-j$ $j-1$
+    $S(j)\ (2\leq j\leq n+1)$ n+2jn+2-j j1j-1
     $\vdots$ $\vdots$ $\vdots$
-    $S(n+1)$ $1$ $n$
+    $S(n+1)$ 11 $n$
   2. (2)
@@
     $0\rightarrow\Gamma\rightarrow(\bigoplus_{i=2}^{n+1}I(i))\oplus I(2)\rightarrow I(2)\rightarrow I(n+1)\rightarrow\cdots\rightarrow I(2)\rightarrow I(1)\rightarrow 0$.
-    Hence $\Gamma$ is $(n+1)$-Gorenstein.
   4. (4)
@@
     in ${\rm{{mod}}}(\Gamma)$ induces the following Auslander-Reiten sequence in $\mathrm{proj}^{\leq n}(\Gamma)$:
-    $0\longrightarrow P(1)\longrightarrow P(n+1)\longrightarrow S(n+1)\longrightarrow 0$
+    $0\longrightarrow P(1)\longrightarrow P(n+1)\longrightarrow S(n+1)\longrightarrow 0.$
   9. (9)
@@
 By their result, i.e.
-higher Auslander correspondence for dualizing $R$-varieties, we know that an Artin $R$-algebra $\Gamma$ is an $n$-Auslander algebra if and only if the category of all finitely generated projective $\Gamma$-modules is an $n$-abelian category [IJ17, Theorem 1.2].
-But by some results due to Beligiannis [Be15], even for $n$-minimal Auslander-Gorenstein algebras, one can see that the category of all finitely generated Gorenstein projective $\Gamma$-modules is not an $n$-abelian category in general.
+higher Auslander correspondence for dualizing $R$-varieties, we know that an Artin $R$-algebra $\Gamma$ is an $n$-Auslander algebra if and only if the category of all finitely generated projective $\Gamma$-modules is an $n$-abelian category [29, Theorem 1.2].
+But by some results due to Beligiannis [12], even for $n$-minimal Auslander-Gorenstein algebras, one can see that the category of all finitely generated Gorenstein projective $\Gamma$-modules is not an $n$-abelian category in general.
 Hence, based on our results, there is a difference between the abelianness of ${\mathrm{Gproj}}^{\leq n-1}(\Gamma)$ and the $n$-abelianness of ${\mathrm{Gproj}}(\Gamma)$ in general.
-Acknowledgements: The authors were supported by the National Natural Science Foundation of China (No.
-12071137), by Key Laboratory of MEA (Ministry of Education), by Shanghai Key Laboratory of PMMP (No.
-22DZ2229014), and by Fundamental Research Funds for the Central Universities.
+Acknowledgments: It is our great pleasure to thank many reseachers for their insightful suggestions and remarks, which include, but not restricted to, Zhaoyong Huang, Xiaojin Zhang, Bin Zhou, and Shijie Zhu.